### v5

# A Global Minimum Tax for Large Firms Only:

Abstract
The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms. We analyze tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT. Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals. However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals. Our calibration exercise shows that the implementation of a 15% GMT rate results in a regime where only the havens adopt split tax rates. Upon GMT introduction, welfare and tax revenues fall in the tax havens but rise in the non-haven, yielding a positive net gain worldwide.
Keywords: multinational firms; tax avoidance; global minimum tax; profit shifting
JEL classification: F23, H25, H87
1 Introduction
Profit shifting by multinational enterprises (MNEs) has long been a major problem for corporate taxation worldwide. Using aggregate data, Tørsløv et al., (2023) estimate that more than one-third of all foreign-earned corporate profits of multinationals are shifted to tax havens. The important role of tax havens for profit shifting is confirmed in other studies using microdata of multinational affiliates/headquarters in France (Davies et al.,, 2018), the UK (Bilicka,, 2019), and the US (Guvenen et al.,, 2022). In response to the large revenue losses caused by profit shifting, the OECD has launched an action plan to fight base erosion in OECD countries, and in particular the profit shifting to tax havens (OECD,, 2013). A core development in this endeavor is the introduction of a global minimum tax (GMT), which is applied to large MNEs (OECD, 2020a, ; OECD, 2020b, ). The GMT with a tax rate of 15% has been agreed upon by a group of more than 130 countries in 2021 including many tax havens, and most of these countries are expected to have the GMT enacted by 2026.111The list of countries joining the agreement on the GMT, also known as Pillar 2, is found in OECD, (2021). The United States has already enacted, in its 2017 Tax Cuts and Jobs Act, a tax on Global Intangible Low-Tax Income (GILTI) that works in many respects like a GMT (see Chodorow-Reich et al.,, 2024 for details). The global revenue gains from a 15% GMT are estimated to be in the range of 155 to 192 billion USD, or 6.5 to 8.1% of global corporate tax revenues (Hugger et al.,, 2024, Section 8.2). [^2] Baraké et al., (2022, Table 1) break down the global revenue changes of the GMT by country and arrive at a revenue gain of around 55 to 67 billion EUR for the US, and a gain of similar magnitude for the EU as a whole. Estimates of both Hugger et al., (2024) and Baraké et al., (2022) are based on models where governments and multinational firms mechanically respond to the GMT. Wang, (2020) sets up a quantitative corporate tax framework that endogenizes the behavior of both firms and governments. Using similar frameworks, Ferrari et al., (2023) and Shen, (2024) quantify the effects of the GMT. While no econometric estimates exist for the global minimum tax yet, there are experiences from similar policies enacted at the national level. See Buettner and Poehnlein, (2024) for an analysis of the effects of a minimum tax rate for German municipalities.
An important limitation of the GMT is, however, that it applies only to “large” multinationals, defined as multinational groups that have had no less than 750 million EUR in total annual revenues in at least two out of the last four years. [^3] Since 2016, multinational groups above this threshold size are also required to file individual reports on their activities, profits, and taxes paid in each of the countries in which they have a presence (country-by-country reporting). Fig. 1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of all MNEs, which are responsible for about 90% of all MNE profits, have consolidated accounts above this threshold and are therefore covered by the GMT. [^4] This corresponds to the estimate of the OECD, which arrives at a coverage rate of 90% of profits, using the Orbis database and other data sources (OECD, 2020a, , # 505, p. 233). Nevertheless, a substantial amount of MNE profits exceeding 300 billion EUR remain outside the scope of the GMT. Moreover, the MNE sample in Orbis is not comprehensive and the database is known for oversampling large firms (Bajgar et al.,, 2020). Therefore, a coverage rate of the GMT of 90% of all MNE profits can be considered as an upper bound of the true profit share covered by the GMT.
One implication of this incomplete coverage is that low-tax countries might respond to the introduction of the GMT by using a split corporate tax system, where a lower tax rate applies to firms below the GMT threshold. In fact, such a split is inherent in the regulation that tax havens are allowed to implement the GMT by raising their (low) general tax rate only for MNEs above the GMT threshold through a specific top-up tax. [^5] This is known as the Qualified Domestic Marginal Top-Up Tax (QDMTT). See Devereux, (2023) for a discussion. Tax havens are fully aware of this option. Ireland and Liechtenstein, for example, have already decided to keep their general tax rates at 12.5%, but top up the tax rate to 15% for affiliates of foreign MNEs above the threshold (Government of Ireland,, 2023; PwC,, 2024 for Liechtenstein). Hungary and Bulgaria have adopted a similarly split tax rate by joining the GMT agreement while maintaining their regular tax rates of 9% and 10%, respectively, for firms not covered by the GMT (Ernst & Young,, 2023 for Hungary; PwC,, 2023 for Bulgaria). [^6] The split corporate tax system is also expected to be introduced in Bermuda and Jersey (Government of Bermuda,, 2024; Government of Jersey,, 2023). These responses indicate that the effects of the GMT cannot be analyzed assuming a single, uniform corporate tax rate across all countries.
Figure 1: GMT coverage
Source: Orbis database, own calculations.
Notes: The left bar in each year shows the share of MNEs covered by the GMT out of all MNEs. The right bar in each year shows the share of pre-tax profits earned by MNEs covered by GMT out of pre-tax profits by all MNEs. The MNEs covered by GMT are those whose annual revenues exceed $750$ million EUR in at least two of the last four years out of all MNEs. See Appendix B for details.
Against this background, this paper analyzes under which conditions tax competition leads countries to differentiate their tax rates for large and small MNEs in response to the introduction of a GMT. We also study the welfare implications of such split tax system. Finally, we explore whether gradual reforms to the GMT—either through an increase in the GMT rate or an expansion of its coverage—enhance welfare for each country and for the world as a whole.
To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits from there to a set of symmetric tax-haven countries. Countries maximize their welfare, a weighted sum of tax revenues and (for the non-haven) the post-tax income of MNEs’ owners, and they compete for the MNEs’ profits under the GMT with incomplete coverage. Both the non-haven and the havens can decide to commit, in the first stage, to set a uniform tax rate for all their multinationals, or not. In the second stage, all countries choose their tax rates, either uniform or split, where the tax rate applicable for large MNEs must be at or above the GMT level. In the third stage, multinationals make their profit-shifting choices.
Our analysis shows that the introduction of a partial GMT leads to a sequence of tax competition regimes between the non-haven and the havens. If the GMT rate is binding for the tax havens, but still low enough, each of the tax havens will choose to commit to the GMT rate for all MNEs, in order to increase the (uniform) tax rate in the non-haven country. As the GMT rate is further increased, however, the non-cooperative tax equilibrium regime transitions to one where the havens split their tax rates by setting a GMT for large MNEs and a rate lower than the GMT for small MNEs. For sufficiently high GMT rates, a tax regime eventually emerges where the non-haven splits its tax rate as well. This is because a higher GMT rate limits competition for profits from large MNEs, leading the countries to focus instead on profits from small MNEs.
The implications for welfare and tax revenues are as follows. At the regime-switching GMT rates, if one country starts to split its tax rate, the other country discretely loses welfare and tax revenues. In general, however, the non-haven always gains from gradual increases in the GMT rate, whereas the havens weakly gain only for sufficiently low levels of the GMT, and lose otherwise. Gradual increases in the coverage rate of the GMT have similar effects, with the important difference that they never reduce tax revenues in the non-haven country or globally. These results suggest that, although both countries benefit from the introduction of a moderate GMT rate, conflicts of interest will arise once the GMT rate or coverage is increased.
In the final step, we calibrate our model using estimates of profit shifting from a global sample of countries (Tørsløv et al.,, 2023). In our benchmark calibration, the introduction of a $15\%$ GMT rate causes a regime shift: tax haven countries adopt split tax rates, while the non-haven maintains a uniform tax rate above the GMT. This result aligns with the actual responses observed in several tax haven countries, as described above. Regarding welfare and tax revenues, the introduction of the GMT benefits the non-haven while harming havens, leading to net welfare gains worldwide. This conflict of interest between the two groups of countries persists in the two gradual reforms likely to follow: increasing the GMT rate from $15\%$ to $18\%$ and expanding GMT coverage to all MNEs. However, the conflict is less pronounced for the latter reform.
Our model contributes to the recent theoretical literature on the effects of tax competition in the presence of a GMT. In a model with many non-haven and haven countries, Johannesen, (2022) shows that introducing a GMT will always benefit the tax havens, because the GMT reduces the competition between them. In contrast, the welfare effects of the GMT are generally ambiguous for the non-haven country, because of the offsetting changes in tax revenues and firms’ post-tax profits. Hebous and Keen, (2023) also analyze the effects on welfare in a model where two countries differ in terms of both their size and their valuation of public spending. Their quantitative results indicate that GMT levels up to a range of 17 to 20% can constitute strict Pareto improvements, increasing welfare in both high-tax and low-tax countries. Hindriks and Nishimura, (2022) consider governments who choose not only tax rates but also tax enforcement levels. They show that the introduction of the GMT hinders tax enforcement cooperation between the high-tax and the low-tax country, making the high-tax country worse off. Finally, Janeba and Schjelderup, (2023) study a three-country model where two non-haven countries compete for real foreign direct investment (FDI) while simultaneously competing for profit shifting with a third, tax-haven country. In this setting the introduction of the GMT intensifies tax competition for real FDI. While tax revenues in the non-haven countries will still increase if countries compete via corporate tax rates, net revenue gains are zero if competition is via lump-sum subsidies. [^7] Using an international trade model of firms without profit-shifting motives, Nocco et al., (2024) examine a rich set of corporate taxes and subsidies on sales, production, and profits. Specifically, they identify the globally welfare-maximizing combination of non-discriminatory taxes and subsidies including a global minimum tax on firms’ profits.
All these models have in common that the GMT must be levied uniformly on all firms, thus ignoring the existence of a threshold below which the GMT does not apply. In contrast, our model highlights the critical role of this threshold, as it allows countries to use different tax rates on firms with revenues above and below the threshold. This feature links our model to the literature on discriminatory tax competition, which examines whether tax revenues are higher in the non-cooperative tax equilibrium when countries may, or may not, set different tax rates on tax bases with different degrees of international mobility (Janeba and Peters,, 1999; Keen,, 2001; Janeba and Smart,, 2003). [^8] This literature has been extended to cover differences in country size (Bucovetsky and Haufler,, 2007), and imperfect competition in product markets (Gaigné and Wooton,, 2011). A related type of discrimination arises when profit shifting is monitored by governments in a deliberately loose way (Peralta et al.,, 2006). In our setting, tax discrimination will not be profitable for either country in the absence of the GMT, but it will arise in equilibrium if a sufficiently high level of the GMT is imposed. A distinct, but closely related literature has addressed the issue of whether the existence of tax havens—by allowing non-haven countries to tax-discriminate in favor of mobile, multinational firms—raises or reduces tax revenues and welfare in the non-haven countries (Slemrod and Wilson,, 2009; Hong and Smart,, 2010; Johannesen,, 2010; Elsayyad and Konrad,, 2012).
Our analysis proceeds as follows. In Section 2 we set up our model of tax competition and characterize the sub-game perfect Nash equilibria that result in different tax regimes. Section 3 then turns to the welfare effects of introducing a GMT, and of increasing its tax rate and coverage. Section 4 calibrates our model to real-world data and quantifies its effects. Section 5 concludes.
2 The model
2.1 Setup
We consider a tax competition model between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$. Tax competition arises between each of the tax havens and the non-haven. There are a large number of heterogeneous, multinational enterprises (MNEs) owned by individuals in the non-haven, each having a headquarters in the non-haven and an affiliate in each of the tax havens. The headquarters in the non-haven makes exogenous profits in this country, $\pi$, which are distributed in the range $[\underline{\pi},\infty)$, with a cumulative distribution function $F(\pi)$. Hence, all real activity occurs in the non-haven, and this is where the “true” profits accrue in its entirety. MNEs can shift a share $\theta\in[0,1]$ of their non-haven headquarters’ profits to the set of tax haven affiliates, in order to maximize their post-tax profits. [^9] This is the simplest possible setting for profit shifting. See e.g., Johannesen, (2022) or Krautheim and Schmidt-Eisenlohr, (2011) for a similar assumption. In particular, this setting avoids the possible non-existence of Nash equilibria that arise in models where the location of heterogeneous firms is endogenous (Baldwin and Okubo,, 2009; Davies and Eckel,, 2010; Haufler and Stähler,, 2013).
We assume that the non-haven and all tax havens have signed the GMT agreement, and are therefore bound to set tax rates of at least the GMT rate, $t_{M}$. [^10] Devereux, (2023) discusses how the specific institutional setting of the GMT, in particular the Under-Taxed Payments Rule (UTPR), gives incentives to both non-haven countries and haven countries to join the GMT agreement, once this has been initiated by a critical number of large non-havens. Note also that we assume the tax base of the GMT to be the same as that of national corporate taxes, thus ignoring extra tax deductions (labelled “substance-based income exclusion”) under the GMT. See Schjelderup and Stähler, (2024) for a study focusing on the latter issue. However, the GMT applies only to MNEs that make exogenous profits no smaller than $\pi_{M}$, called “large MNEs” in our analysis. Therefore, countries remain free to set a lower tax rate than the GMT for all MNEs that make profits less than $\pi_{M}$, called “small MNEs.”
The non-haven country maximizes a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of the heterogeneous multinationals, which are all owned by residents of the non-haven country. Since tax havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
We consider a three-stage game. In the first stage, the non-haven and haven countries simultaneously choose whether to commit to a single tax rate for all MNEs, or not. [^11] This is related to the literature on endogenous timing, or leadership, in tax competition, where countries commit in a pre-play stage to set taxes early or late. See Kempf and Rota-Graziosi, (2010). To comply with the GMT agreement, this single tax rate has to be equal to, or higher than, the GMT rate. In our main model, we assume that the set of tax havens collectively decides on whether to set a single GMT rate, or split their tax rates. This assumption simplifies the algebra, but it does not change the qualitative nature of our results. [^12] We discuss the alternative of a decentralized first stage decision of the tax havens in Appendix A.2. In the second stage, all countries choose their tax rate(s) in a non-cooperative way. Countries that have committed to a single tax rate in the first stage set a single tax rate, whereas countries that have not committed choose different tax rates for large and small MNEs (where the tax rate on large MNEs has to be at least $t_{M}$). In the final stage, MNEs engage in profit shifting, given the tax rates in the non-haven and the symmetric tax havens. We solve this three-stage game by backward induction. The equilibrium concept we rely on is the sub-game perfect Nash equilibrium, which we call the tax-competition equilibrium.
2.2 Profit shifting by multinationals
Each MNE shifts a share $\theta_{h}$ of its exogenous pre-tax profits $\pi$ from their non-haven affiliate to each of the tax haven affiliates $h$. Firms incur transaction costs when shifting profits. These costs can either be thought of as concealment costs, which have to be incurred to hide the profit shifting from the non-haven country’s tax authorities, or as the expected fine that is to be paid when profit-shifting is detected and sanctioned. We assume that the costs of shifting $\theta_{h}\pi$ of profits to each haven $h$, given by $C_{h}(\cdot)$, are convex and, for analytical simplicity, quadratic in the level of profit-shifting, $\theta_{h}\pi$. Higher total (true) profits make it easier, however, to hide a given amount of profit-shifting. [^13] This specification is widely used in the literature on profit shifting (e.g., Hines and Rice,, 1994; Huizinga and Laeven,, 2008; Suárez Serrato,, 2018). Therefore, the shifting costs for a firm with profits $\pi$ to a haven $h$ are $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$. The exogenous parameter $\delta$ captures the ease with which profits can be shifted across countries. It incorporates any efforts that countries take to prevent profit-shifting. [^14] In Hindriks and Nishimura, (2022), this cost parameter is endogenously chosen by governments, either cooperatively or non-cooperatively.
Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the tax haven countries, a MNE maximizes its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\Pi}$, by choosing the share of profit shifting to each haven:
$\max_{\{\theta_{h}\}_{h=1}^{H}}\ \widetilde{\Pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)=(1-t_{n})\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi+\sum_{h=1}^{H}(1-t_{h})\theta_{h}\pi-\sum_{h=1}^{H}\frac{\delta(\theta_{h}\pi)^{2}}{2\pi}.$ (1)
The MNEs’ optimal level profit shifting to each tax haven, $\theta_{h}$ is then:
$\theta_{h}=\frac{t_{n}-t_{h}}{\delta},$ (2)
which is independent of $\pi$ and therefore holds for all MNEs simultaneously. [^15] In principle, our analysis can incorporate $\theta<0$, which occurs if $t_{n}<t_{h}$. However, this case will never arise in the tax-competition equilibrium we will see shortly.
The optimal $\theta$ defined by (2) allows us to incorporate the MNEs’ adjustment to tax differentials in a very simple and compact form. It implies that MNEs of different size will respond to international tax differentials with the same tax base elasticity. Substituting (2) into (1) and using the symmetry of tax havens, the tax base of a MNE with profits $\pi$ in the non-haven country is $TB_{n}=(1-\sum_{h=1}^{H}\theta_{h})\pi=[1-H(t_{n}-t_{h})/\delta]\pi$. This yields a tax base elasticity of an individual MNE in the non-haven country such that $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, which is independent of $\pi$.
The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive. Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity. However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs). [^16] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms. Applying a similar empirical strategy to US firms, Coles et al., (2022) also find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size. See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different size. Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark. As a consequence, there is no reason per-se in our model that countries differentiate tax rates between MNEs of different size. However, as we will see below, a split tax regime can nevertheless occur in equilibrium as a result of introducing a GMT for large firms only.
2.3 Governments’ tax setting choices
Before examining the effects of GMT, we first solve the unconstrained optimization problem in the absence of a GMT. The non-haven country $n$ maximizes a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, where all MNE owners reside exclusively in the non-haven country. The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal valuation of public funds.
The welfare premium on tax revenues, relative to private income, is a standard assumption in international taxation, which motivates positive tax rates in equilibrium. It can be given two alternative interpretations. The first is a redistributive (or political-economy) motive, when the government uses the corporate tax to redistribute income from capital owners to other households in the economy, which possess no earned income. The other interpretation is that corporate tax revenue is used to reduce other distortive taxes, and the (exogenous) deadweight loss associated with them. [^17] The latter argument assumes that there is a fixed excess burden of taxation in a country’s tax system (see e.g., Keen and Lahiri,, 1998). This is a plausible assumption in our setting, as we focus on the corporation tax, which accounts for less than 10% of total tax revenue in all developed countries. Either setting leads to a welfare gain for the non-haven country from levying a positive corporation tax on its MNEs. Finally, for $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues. This special case underlies, for example, the analyses in Hindriks and Nishimura, (2022) or Janeba and Schjelderup, (2023).
The government of country $n$ therefore maximizes
[…]
$=(1-t_{n})\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi+\sum_{h=1}^{H}(1-t_{h})\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi-\sum_{h=1}^{H}\frac{\delta}{2}\left(\frac{t_{n}-t_{h}}{\delta}\right)^{2}\Pi$
$\qquad\qquad+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
[…]
Each of the tax haven countries $h$ only cares about its tax revenues (multiplied by the same marginal value of public funds, $\lambda$, as for country $n$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
[…]
To determine world welfare, we take a utilitarian approach and define $G_{W}\equiv G_{n}+\sum_{h=1}^{H}G_{h}$, using (2.3) and (4).
Solving the first-order conditions of the two countries yields best responses: [^18] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
$t_{n}=\frac{(\lambda-1)\left(\sum_{h=1}^{H}t_{h}+\delta\right)}{H(2\lambda-1)},\qquad t_{h}=\frac{t_{n}}{2}.$ (5)
Solving for the optimal tax rates in the unconstrained benchmark (“Regime 0”) then gives
$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)},\qquad t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}.$ (6)
[…]
A few comments on (6) are in order. First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$. Second, they are rising in the marginal value of public funds, $\lambda$. A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the valuation of tax revenues vis-à-vis the post-tax profits of MNEs. The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well. Finally, each of the havens without an independent tax base will choose one-half of the tax rate of the non-haven. [^19] Note the important difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT. This contrast arises from different specifications of profit-shifting costs. Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country. This winner-takes-all situation among the havens drives their equilibrium tax rates to zero. By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs. Hence each MNEs has an incentive to diversify its allocation of shifted profits across all havens. As we discuss in our calibration in Section 4, a setting with positive tax rates in haven countries in the unconstrained tax equilibrium is more in line with the empirical evidence.
We can also derive the share of shifted profits in this unconstrained tax equilibrium. This is given by:
$\sum_{h=1}^{H}\theta_{h}^{0}=H\cdot\left(\frac{t_{n}^{0}-t_{h}^{0}}{\delta}\right)=\frac{H}{\delta}\cdot\frac{\delta(\lambda-1)}{H(3\lambda-1)}=\frac{\lambda-1}{3\lambda-1}<\lim_{\lambda\to\infty}\sum_{h=1}^{H}\theta_{h}^{0}=\frac{1}{3}.$ (7)
Hence, when the non-haven country maximizes tax revenues ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium. For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
2.4 Tax regimes
With a GMT in place, there is a first stage of the game in which each country chooses to commit or not to a single tax rate, which can either be the GMT rate or a tax rate above the GMT. If a country does not commit to a single tax rate, then it will choose to split tax rates in the second stage, and set different tax rates for large and small MNEs. MNEs with profits equal to or greater than the threshold $\pi_{M}$, i.e., $\pi\in[\pi_{M},\infty)$, are subject to the GMT requirement (large MNEs), whereas those with $\pi\in[\underline{\pi},\pi_{M})$ are not (small MNEs). We denote by $\phi$ the share of total profits earned by large MNEs, or the coverage rate of the GMT, which is exogenously given:
[…]
Non-haven Haven
single non-GMT rate single GMT rate split tax rate
single non-GMT rate Regime 0 Regime 1 Regime 2
single GMT rate Regime 3
split tax rate Regime 4
Table 1: Possible regimes of the non-cooperative tax game
Table 1 shows that this sequence yields three potential choices for each of the two sets of countries. Of the nine potential tax regimes, four regimes cannot occur in equilibrium. In these potential regimes, if the non-haven sets a single GMT rate, the havens would choose a single tax that is at the same level or even higher. Hence tax revenues in each haven country would be zero. Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven country for small MNEs. Alternatively, if the haven sets a single non-GMT or a single GMT tax rate, the non-haven will never want to split its tax rate, because the tax base elasticity of its MNEs is independent of profits. Hence only five possible tax regimes remain.
In Regime 0, the GMT rate is not binding for any country; this is the case of unconstrained tax competition analyzed in the previous section. As the GMT rate is continuously increased, the tax competition equilibrium passes through four other regimes. In the following, we first state the full characterization of the sub-game perfect Nash equilibrium of the three-stage tax competition game, before commenting on each of these regimes.
Proposition 1
(Equilibrium regimes with binding GMT)
Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal valuation of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$. As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterized by the following four regimes, with all symmetric haven countries choosing the same tax schedule.
  • (i)
    Regime 1: $t_{M}\in\left[t_{M}^{0},t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\right]$.  The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\pi_{M},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\pi_{M},\infty).$
  • (ii)
    Regime 2: $t_{M}\in\left(t_{M}^{1},t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}\right]$.  The non-haven sets a single non-GMT rate, and the haven splits its tax rate and chooses the GMT rate for large MNEs, but a lower rate than the GMT for small MNEs:
    $t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}\ \ \text{for}\ \pi\in[\pi_{M},\infty),\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}\ \ &\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}\ \ &\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$
  • (iii)
    Regime 3: $t_{M}\in\left(t_{M}^{2},t_{M}^{3}\equiv\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}\right]$.  The non-haven chooses a single GMT rate, and the haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs:
    $\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
  • (iv)
    Regime 4: $t_{M}\in(t_{M}^{3},1]$.  Both the non-haven and the haven split their tax rates and choose the GMT rate for large MNEs, but a lower tax rate for small MNEs:
    $\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
Proof: See Appendix A.1.
In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates. Our discussion below will focus on tax revenue effects. This is strictly correct only for the tax havens. It is also possible for the non-haven, however, because the marginal value of public funds, $\lambda$, is constant in our analysis, and it always exceeds the valuation of private income. All formal proofs are relegated to the Appendix.
In Regime 1, where $t_{M}$ rises above the haven’s unconstrained tax rate $t_{h}^{0}$, the haven is forced to set $t_{M}$ for large MNEs. By contrast, the non-haven’s unconstrained rate is still above the GMT. As the non-haven is not bound by the GMT, it will continue to choose a single tax rate above the GMT for all MNEs. For the tax havens, the choice is therefore whether to apply the GMT also to small MNEs, or use a split tax schedule. By committing to a single GMT rate for all MNEs, the havens can induce the non-haven to set a higher uniform tax rate than in the unconstrained equilibrium. This argument is strongest when tax havens collectively decide on whether to commit to a single tax rate, as we assume in our main model. [^20] This is seen from the non-haven’s best response function (5), which is increasing in $t_{h}$. However, as we show in Appendix A.2, the same qualitative argument also holds for decentralised decision-making of tax havens. Havens therefore trade off the gains from inducing a higher uniform tax rate in the non-haven against the revenue losses from not undercutting the non-haven’s tax rate even more for small MNEs. In Regime 1, the first effect dominates and the havens find it profitable to commit to the GMT for all MNEs. As a result, all countries choose higher uniform tax rates in Regime 1, as compared to the unconstrained equilibrium without a GMT.
In Regime 2, where $t_{M}$ rises above the regime-switching rate $t_{M}^{1}$, the non-haven is still not bound by the GMT, and will continue to set a single tax rate above the GMT for all MNEs. For the havens, however, the trade-off described in Regime 1 above changes. As the GMT is now higher, there is less to gain from setting a uniform GMT rate for all MNEs, because the tax differential and thus shifted profits are reduced. [^21] This is seen from Regime 2 of Proposition 1, where the tax differential, $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases in $t_{M}$. At the same time, given the higher GMT rate, there is more to gain for the havens from undercutting the non-haven’s tax rate for small MNEs. In Regime 2, therefore, the havens will find it optimal to split their tax rate and set $t_{M}$ only for the large MNEs, but choose a lower tax rate for small MNEs. [^22] In practice, this split is done by choosing the lower tax rate for small MNEs as the general corporate tax rate, but topping up this tax rate for large MNEs. See footnote 5 in the introduction. Johannesen, (2022) has shown formally that it is always rational for tax havens to use the QDMTT up to the total tax level of the GMT, rather than leaving the top-up taxation to the home countries of the MNEs. Note, finally, that the upper boundary of Regime 2, $t_{M}^{2}$, is rising in the GMT coverage rate, $\phi$. This is because the non-haven country faces less competition from the havens (i.e., higher tax rates $t_{h}$) for large firms. If large MNEs make up a large part of the corporate tax base, the non-haven country will therefore be able to set a uniform tax rate above $t_{M}$ for a wider range of GMT rates.
In Regime 3, where $t_{M}$ rises above the regime-switching rate $t_{M}^{2}$, the non-haven will no longer charge a tax rate above the GMT level, but it is bound by the GMT tax rate $t_{M}$ for large firms. By committing to set $t_{M}$ for all MNEs, it induces the haven countries to charge a higher tax rate for small MNEs. This argument is analogous to the one that applied to the tax havens in Regime 1. In equilibrium, therefore, the non-haven sets $t_{M}$ for all firms, rather than splitting its tax rate, and the havens respond by charging $t_{M}/2$ for small firms.
Finally, in Regime 4, where $t_{M}>t_{M}^{3}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set the GMT rate for all firms. [^23] This is seen from Regime 3 of Proposition 1, where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases in $t_{M}$. To avoid large tax base losses from small MNEs, the non-haven therefore abandons its commitment and also undercuts the GMT for small MNEs. Again, this argument is analogous to the one for the havens to end their commitment in Regime 2. In the Regime 4 equilibrium, both countries therefore split their tax rates and set their unconstrained tax rates for small MNEs, while adhering to the GMT for large MNEs.
Figure 2: Equilibrium tax rates for different GMT rates
Notes: The figure shows the equilibrium tax rates in the non-haven (a) and the haven (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes (c). Parameter values given in Table 2 affect the precise levels of tax rates and regime-switching points, but do not affect qualitatively the sequence and pattern of tax-competition equilibria.
Figure 2 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of $t_{M}$. The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b). The equilibrium regimes are summarized in panel (c). The solid curves give the tax rate(s) in the tax-competition equilibrium under the GMT; the dashed horizontal line is the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0). Tax rates in both countries generally rise as the GMT increases, but there are distinct patterns in the different regimes. At the switch from Regime 1 to 2, the haven countries start splitting their tax rates and their tax rate on small MNEs drops discontinuously. At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0. The haven countries respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
3 Welfare effects of the GMT
Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of the introduction of the GMT and of two gradual reforms: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$. As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case. Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
3.1 Introduction of the GMT
We first compare welfare levels in the tax-competition equilibrium to those in the unconstrained equilibrium, Regime 0. This is summarized in:
Proposition 2
(GMT introduction)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1, and assume the GMT coverage rate is not too low so that $\phi\in(3/4,1)$. Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
  • (i)
    a rise in welfare in the non-haven country for all Regimes 1 to 4.
  • (ii)
    a rise in welfare in the haven country for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
  • (iii)
    a rise in world welfare for all Regimes 1 to 4.
Proof: See Appendix A.3.
Introducing a small GMT slightly higher than the havens’ unconstrained rate $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the haven countries. At the same time, profit shifting is reduced in comparison to the no-GMT benchmark due to a smaller tax differential. The sum of these effects must certainly benefit the non-haven. It also benefits the tax havens, as long as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits. However, once the GMT rate exceeds a certain threshold, $t_{M}^{+}$, which still lies in Regime 1, the reduced tax differential and thus the fall in shifted profits will dominate from the perspective of the haven countries. Therefore, for all levels of the GMT above $t_{M}^{+}$, tax havens lose from the introduction of the GMT while the non-haven gains. [^24] This conflict of interest does not arise in (Johannesen,, 2022, Proposition 6), where havens gain from the introduction of a GMT at any level, as long as profit shifting is not completely eliminated. In his setting where MNE shift profits only to the lowest-tax country, all the tax havens end up setting their tax rate to zero. Therefore, revenues in the unconstrained tax competition equilibrium are zero for each haven, and any positive GMT rate generates a revenue increase for the representative haven. See also footnote (19).
In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0. As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs. In this case, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.252525If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate will instead benefit the havens in Regime 3. See Appendix A.3 for details. Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes. These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
3.2 Gradual reforms of the GMT
We now turn to two likely reform options after the GMT has been introduced. We first study the effects of a gradually increasing GMT rate. The welfare effects of this reform are summarized in:
Proposition 3
(Gradual reform: increasing GMT rate $t_{M}$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1. A marginal increase in the GMT rate has the following effects:
  • (i)
    welfare in the non-haven country increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it discretely falls.
  • (ii)
    welfare in the haven country increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases for the remainder of Regime 1 and in Regime 2. It increases in Regime 3 and discretely falls at the last regime-switching point, $t_{M}^{3}$. It does not change in Regime 4.
  • (iii)
    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it discretely falls.
Proof: See Appendices A.3 and A.4.
As is true for the introduction of the GMT (Proposition 2), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium. An exception is the regime-switching point $t_{M}^{1}$ where the havens start splitting their tax rate and set a discretely lower tax rate for small MNEs. Small MNEs then shift more profits to the havens, hurting the non-haven.
For the havens, the welfare (tax revenue) effects of an increase in the GMT depend on the negative effect of reduced profit-shifting opportunities and the positive effect of a higher tax rate in the non-haven. This leads to non-monotonic changes in the havens’ welfare. In Regime 1, the induced higher tax rate in the non-haven dominates initially, but as the GMT continues to rise, this is overcompensated by the reduced levels of profit shifting. The latter is also true throughout Regime 2. In Regime 3, the effect of a higher tax rate dominates again, as the non-haven commits to the uniform GMT rate in order to raise the havens’ tax rates. This regime, where commitment induces effective tax coordination, is similar to (the first part of) Regime 1. It ends at the regime-switching point $t_{M}^{3}$ where the non-haven starts splitting its tax rate. This leads to a discrete fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens.
The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive. At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, where one of the countries starts splitting its tax rate and discretely lowers the tax rate on small MNEs, tax competition is tightened and world welfare falls at the margin.
Next we examine the welfare effects of another likely reform option: a marginal increase in the GMT coverage rate $\phi$, while keeping the GMT rate $t_{M}$ fixed. As shown in (8), this is achieved by reducing the threshold profit level $\pi_{M}$ above which the GMT rate must be charged. The results are summarized in:
Proposition 4
(Gradual reform: increasing GMT coverage rate $\phi$)
Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1. Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
  • (i)
    welfare in the non-haven is unaffected in Regime 1 and increases in Regimes 2 to 4.
  • (ii)
    welfare in the haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
  • (iii)
    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
Proof: See Appendix A.5.
In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates. Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates. Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies fewer small MNEs, reduces aggregate profit shifting. This unambiguously hurts the haven and benefits the non-haven. From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
There are two exceptional cases where increasing $\phi$ causes a regime switch. The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rate. This regime switch benefits the non-haven and the world, while it has no effect on the havens. The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT. This latter regime switch, however, leaves welfare in all countries and the world unaffected.
In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3). A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1. One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
4 Quantitative implications
We now examine which regime can be expected under the currently imposed GMT rate. For this purpose, we use a calibrated version of our model and explore its quantitative effects. We calibrate our unconstrained model without GMT (Regime 0) to match basic data on international profit shifting. The results are summarized in Table 2. The coverage rate of the GMT is set at $\phi=0.9$, in line with OECD, 2020a and our own calculations from the ORBIS database (see Figure 1 and Appendix B). The marginal valuation of public funds (MVPF) is set at $\lambda=1.5$ in our calibration benchmark, corresponding to the benchmark value used in Hebous and Keen, (2023). The figures for total pre-tax profits of MNEs, $\Pi$, and the number of tax havens, $H$, are taken from Tørsløv et al., (2023) and the Country-by-Country Report Statistics (CbCR); see the notes to Table 2 for details. We then calculate the cost parameter of profit shifting, $\delta$, to exactly match the GDP-weighted average of effective tax rates in non-haven countries of $18.6\%$, as reported in Tørsløv et al., (2023). Among the non-targeted moments, our model somewhat underestimates the havens’ tax rate in the data, as well as the total profits shifted by MNEs and –accordingly– the non-havens’ loss of tax revenues. Overall, however, our simple model explains the data reasonably well.
Parameter Source
$\Pi=4,623$bUSD Total pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
$\phi=0.9$ Profit share of MNEs covered by GMT OECD, 2020a , Orbis
$H=40$ Number of tax havens Tørsløv et al., (2023)
$\lambda=1.5$ Marginal valuation of public funds Hebous and Keen, (2023)
$\delta=27.4$ Cost of profit shifting Calibrated
Targeted moment Model Data Source
Corp. tax rate of non-haven (Regime 0): $t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ $18.6\%$ $18.6\%$ Tørsløv et al., (2023)
Non-targeted moment Model Data Source
Corp. tax rate of haven (Regime 0): $t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
MNE’s shifted profits: $\sum_{h}\theta_{h}^{0}\Pi$ $660$bUSD $969$bUSD Tørsløv et al., (2023)
Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ $129$bUSD $247$bUSD Tørsløv et al., (2023)
Table 2: Calibration of the model
Notes: Data are from 2019. The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ”Table1” and ”Table U1” in the excel file ”1975-2019 updated estimates: Tables”). The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world. There are $40$ haven countries across the world. The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates. Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7). Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs. With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
Figure 3: Equilibrium welfare and shifted profits
Notes: The figure shows the equilibrium levels of welfare normalized by the MVPF, $\lambda$, in the non-haven (a) and the sum of the $H=40$ haven countries (b), and the aggregate shifted profits (c), for different rates of GMT. Parameter values are given in Table 2. The solid curves in each panel are the equilibrium values under the partial GMT ($\phi=0.9$); the dashed horizontal lines are those in the unconstrained equilibrium (Regime 0).
For this calibrated version of our model, the effects of introducing a GMT at varying tax rates $t_{M}$ on equilibrium tax schedules are shown in Figures 2, which we already discussed in Section 2.4. Figure 3 illustrates the effects on welfare in both non-haven and haven countries, and those on aggregate shifted profits, as compared to their counterparts in the unconstrained equilibrium (Regime 0). To make numbers comparable with tax revenues, the levels of welfare are normalized by the MVPF, $\lambda$.
The most important result is that the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 2. In this regime, the havens find it profitable to split their tax rate and to set a lower tax rate for small MNEs as compared to large MNEs (Proposition 1(ii)). [^26] This result also holds if we assume that the first stage decision of whether to split tax rates or not is made by each tax haven individually (see Appendix A.2). In this case the range of GMT rates that support Regime 1 will be smaller than in our benchmark model, while the range of $t_{M}$ that supports an equilibrium in Regime 2 rises accordingly. This corresponds to the observations mentioned in the Introduction that several tax havens have announced to maintain their low statutory tax rate while topping up the taxation of large MNEs to comply with the GMT, known as the QDMTT (see footnote 5). The GMT introduction reduces the total amount of profits shifted, thereby increasing the non-haven’s welfare while reducing the havens’ welfare (tax revenues).
Panel (a) of Table 3 further quantifies the welfare and revenue effects of a 15% GMT rate for our calibrated model, with the benchmark MVPF of $\lambda=1.5$. Welfare and tax revenue increases in the non-haven country clearly outweigh the revenue losses in the havens, leading to an increase in global welfare and tax revenues. Our calibrated gain in tax revenue in the non-haven (107 billion USD) is close to the summed revenue gains for Europe and the U.S. in Baraké et al., (2022) and worldwide tax revenue gains are in the same range as in Hugger et al., (2024) (see footnote 2). [^27] The percentage increases in corporate tax revenue in our model are not comparable to those in Baraké et al., (2022) and Hugger et al., (2024), as we do not model purely domestic firms. The welfare gain for the non-haven is smaller than its revenue gain, because the private income in the non-haven (the post-tax profits of MNEs net of profit-shifting costs) falls after the introduction of GMT.
(a) MVPF: $\lambda=1.5$.  Regime 0 to 2 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
Tax rate after GMT introduction $19.9\%$ for all $\begin{cases}9.9\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
Welfare change in bUSD (change) $53$ ($1.6\%$) $-7$ ($-12.1\%$) $46$ ($1.4\%$)
Revenue change in bUSD (change) $107$ ($14.5\%$) $-7$ ($-12.1\%$) $99$ ($12.4\%$)
Shifted-profits change in bUSD (change) $267$ $(41.9\%)$ $-267$ $(-41.9\%)$ 00 $(0\%)$
(b) MVPF: $\lambda=3$.  Regime 0 to 1 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
Tax rate after GMT introduction $20.9\%$ for all $15.0\%$ for all —–
Welfare change in bUSD (change) $122$ ($6.4\%$) $2$ ($2.0\%$) $124$ ($6.1\%$)
Revenue change in bUSD (change) $168$ ($26.0\%$) $2$ ($2.0\%$) $170$ ($22.6\%$)
Shifting-profits change in bUSD (change) $425$ $(36.8\%)$ $-425$ $(-36.8\%)$ 00 $(0\%)$
(c) MVPF: $\lambda=10$.  Regime 0 to 1 Non-haven Haven World
Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
Tax rate after GMT introduction $21.3\%$ for all $15.0\%$ for all —–
Welfare change in bUSD (change) $169$ ($17.3\%$) $12.4$ ($9.3\%$) $181$ ($16.3\%$)
Revenue change in bUSD (change) $185$ ($31.1\%$) $12.4$ ($9.3\%$) $197$ ($27.1\%$)
Shifting-profits change in bUSD (change) $463$ $(32.2\%)$ $-463$ $(-32.2\%)$ 00 $(0\%)$
Table 3: Quantitative effects of introducing a $15\%$ GMT
Notes: Each of panels (a) to (c) show the effects of a $15\%$ GMT introduction for different values of the MVPF, $\lambda\in\{1.5,3,10\}$. Results are based on the calibrated parameter values described in Table 2; the GMT coverage rate is $90\%$ ($\phi=0.9$) in particular. The welfare levels are normalized by the MVPF, $\lambda$. With $\lambda=1.5$, the $15\%$ GMT introduction leads the economy from Regime 0 to 2; with $\lambda\in\{3,10\}$, it leads instead to Regime 1. The haven refers to the sum of the $H=40$ haven-countries. Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
One may argue that the MVPF is indeed higher than our benchmark value of $\lambda=1.5$, as evidenced by the very substantial efforts, discussed in the introduction, that non-haven countries take to increase their corporate tax revenues. Panels (b) and (c) of Table 3 therefore present the results of two robustness checks with high values of the MVPF: $\lambda=3$ and $\lambda=10$. The latter value, given in Panel (c) approaches a Leviathan government objective of tax revenue maximization. With these increased values of $\lambda$, the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 1, in which both countries employ a uniform tax rate on all MNEs. With the higher valuation of tax revenues in the non-haven country, the tax havens will find it profitable to commit to a uniform GMT rate for all MNEs, which in turn leads the non-haven to choose a higher uniform tax rate above the GMT, as compared to the benchmark in Panel (a). The GMT rate of $15\%$ therefore acts as a coordination device between the non-haven and the havens, benefitting both countries. [^28] These results depend on the assumption that tax havens make their first stage choice collectively. With decentralized first stage decisions of tax havens (cf. Appendix A.2), the upper boundary for a Regime 1 equilibrium, $t_{M}=t_{M}^{a}$, is calculated to be $9.37\%$, $9.42\%$, and $9.45\%$ for respective values of the marginal valuation of public funds, $\lambda=1.5$, $3$, and $10$. This implies that the introduction of a $15\%$ GMT rate leads to a Regime 2 equilibrium regardless of the level of $\lambda$. For $\lambda=1.5$, the quantitative effects under decentralized decision making of havens are the same as those in Panel (a) of Table 3. For $\lambda=3$ and $\lambda=10$, the effects differ from those in Panels (b) and (c) of Table 3, however, and yield a conflict of interests between the non-haven and the havens that is analogous to the one in Panel (a). Although observations at the time of writing somewhat favor the GMT resulting in split tax rates in the havens, corresponding to Regime 2, we leave open the question of which regime will eventually emerge.
(a) Gains from $t_{M}=0.15$ to $t_{M}=0.18$ Non-haven Haven World
Welfare change in bUSD (change) $34$ ($1.0\%$) $-17$ ($-31.0\%$) $18$ ($0.5\%$)
Revenue change in bUSD (change) $59$ ($7.0\%$) $-17$ ($-31.0\%$) $42$ ($4.7\%$)
Shifting-profits change in bUSD (change) $146$ $(37.9\%)$ $-146$ $(-37.9\%)$ 00 $(0\%)$
(b) Gains from $\phi=0.9$ to $\phi=1.0$ Non-haven Haven World
Welfare change in bUSD (change) $5.4$ ($0.2\%$) $-0.5$ ($-0.9\%$) $4.9$ ($0.1\%$)
Revenue change in bUSD (change) $10.7$ ($1.3\%$) $-0.5$ ($-0.9\%$) $10.3$ ($1.1\%$)
Shifting-profits change in bUSD (change) $27$ $(7.0\%)$ $-27$ $(-7.0\%)$ 00 $(0\%)$
Table 4: Quantitative effects of gradual reforms
Notes: Panel (a) shows gains/losses from an increase in the GMT rate from $t_{M}=0.15$ to $0.2$; panel (b) shows gains/losses from an increase in the GMT coverage rate from $\phi=0.9$ to $1.0$. Results are based on the calibrated parameter values described in Table 2; the MVPF is $\lambda=1.5$ in particular. The welfare levels are normalized by the MVPF, $\lambda$. The haven refers to the sum of the $H=40$ haven-countries. Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
Finally, Table 4 provides a quantitative assessment of two gradual reforms that are likely to follow the newly introduced GMT rate of $15\%$ with incomplete coverage. The first reform is to increase the GMT rate from $15\%$ to $18\%$ while maintaining the coverage rate at $\phi=90\%$ (Panel (a)). The second is to broaden the GMT coverage from $90\%$ to $100\%$ while keeping the GMT rate at $15\%$ (Panel (b)). In both scenarios, we assume our benchmark MVPF of $\lambda=1.5$ and thus, the equilibrium remains in Regime 2 after the reforms. The results indicate that the non-haven gains more from the $18\%$ GMT rate ($+1.0\%$ in welfare; $+7.0\%$ in revenues) than from the full GMT coverage ($+0.2\%$ in welfare; $+1.3\%$ in revenues). The greater gains for the non-haven from the $18\%$ GMT rate are driven by the stronger reduction in profit shifting to the havens ($-37.9\%$), as compared to a full GMT coverage ($-7.0\%$). Hence the increase in the GMT rate creates a much sharper conflict of interest with the tax havens, as compared to the extension of GMT coverage. In this sense, covering more MNEs under the GMT may be a politically more feasible option than increasing the GMT rate. However, even a full coverage of the GMT will likely bring only moderate gains to the non-haven and to the world, while entailing higher administrative and compliance costs.
5 Conclusion
This paper has analyzed the effects of a global minimum tax (GMT) that is confined to large multinational enterprises (MNEs), thus leaving at least 10% of the global multinational tax base outside its scope. Using a simple model with profit shifting by heterogeneous MNEs, we have shown that this partial coverage of the GMT gives rise to a sequence of tax competition equilibria between a non-haven and a set of symmetric tax haven countries. In particular, introducing a low GMT rate that still binds the tax havens induces them to commit to a single GMT rate for all MNEs and thus results in higher welfare and tax revenues in both sets of countries. However, a further increase in the GMT rate leads first the tax havens, and then the non-haven, to split their tax rate for large and small MNEs, creating a conflict of interest between the two groups of countries. A similar conflict also arises from broadening the coverage of MNEs subject to the GMT.
The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the haven countries split their tax rate and undercut the GMT for small MNEs (Regime 2). This finding corresponds to the observation that tax havens such as Ireland, Liechtenstein, Bermuda, or Bulgaria maintain their regular corporate tax rates below the level of 15%, but top up the taxation of large multinationals to the level of the GMT. In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the tax havens lose, resulting in a positive net gain globally.
Our main objective in this paper has been to understand the implications of an incomplete coverage of the GMT. We have done so in a highly stylized model, in order to develop some sharp intuitions. With the insights gained, it would be fruitful to extend our basic model to incorporate further important aspects of the global minimum tax and the responses of multinational firms. A first extension would allow MNEs to adjust their real pre-tax profits in response to changes in tax rates, for example, by modifying their investment decision. A second extension would incorporate a splitting response of multinationals seeking to benefit from potentially lower tax rates on small MNEs. A third extension would consider the substance-based income exclusion, i.e., additional tax deductions for MNEs that increase investment and employment. We leave these extensions for further research.
Appendix A: Proof of Propositions
A.1. Proof of Proposition 1
Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$. A new equilibrium regime emerges when $t_{M}\geq t_{M}^{0}$. Since haven countries are all symmetric, we are concerned with the representative haven country $h$ in what follows.
Regime 1: $t_{M}\in[t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
When the non-haven sets a single non-GMT rate and the haven sets a single GMT rate, the equilibrium tax rates and payoffs are
$t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)},\qquad t_{h}^{1}=t_{M}.$ (A.1)
$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\qquad G_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
When the non-haven sets a single non-GMT rate and the haven splits its tax rates, the equilibrium tax rates are
$t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]},\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$ (A.2)
$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\qquad G_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
Given the non-haven’s single non-GMT rate, the haven prefers to set a single GMT rate, if
$G_{h}^{1}-G_{h}^{2}=$ (A.3)
$\frac{\lambda\overbrace{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]}^{\geq 0}\overbrace{\left[\delta(\lambda-1)(2\lambda-1)-Ht_{M}\left\{\lambda(3\lambda-1)-\phi(\lambda-1)^{2}\right\}\right]}^{\geq 0}}{\delta H^{2}(2\lambda-1)[\lambda(3+\phi)-(1+\phi)]^{2}}(1-\phi)\Pi\geq 0,$
where the strict equality holds at $t_{M}\in\{t_{M}^{0},t_{M}^{1}\}$. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate, and the haven sets a single GMT rate, given by (A.1). Table A1 shows the relevant payoffs in Regimes 1 and 2, where the GMT rate is not binding for the non-haven.
Non-haven Haven
single GMT rate split tax rate
single non-GMT rate $G_{n}^{1},\quad G_{h}^{1}$ $G_{n}^{2},\quad G_{h}^{2}$
single GMT rate
split tax rate
Table A1: Payoff matrix for Regimes 1 ($t_{M}\in[t_{M}^{1},t_{M}^{2}]$) and 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$)
Notes: In Regimes 1 and 2, the GMT is binding only for the haven. The potential regimes where the non-haven does choose a single GMT rate or split tax rates are irrelevant because, whenever possible, it always chooses a single non-GMT rate that solves the unconstrained maximization problem.
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rates in response to a non-haven’s single non-GMT rate. The non-haven is still not bound by the GMT and continues to chooses its optimal unconstrained tax rate. Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rates, with tax rates given by (A.2).
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximizing rate in response to the haven’s splitting rates. The non-haven thus chooses either a single GMT rate or a splitting tax schedule. In response, the haven must always split its tax rate; otherwise it cannot obtain positive tax revenues.
When the non-haven sets a single GMT rate and the haven splits its tax rates, the equilibrium tax rates and payoffs are
$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\qquad G_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
When both the non-haven and the haven split their tax rates, the equilibrium tax rates are
$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\qquad G_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
Given the haven’s splitting rates, the non-haven prefers to set a single GMT rate, if
$\begin{split}&G_{n}^{3}-G_{n}^{4}=\frac{\overbrace{\left[Ht_{M}(3\lambda-1)-2\delta(\lambda-1)\right]}^{>0}\overbrace{\left[2\delta(\lambda-1)(8\lambda-3)-Ht_{M}(3\lambda-1)(4\lambda-1)\right]}^{\geq 0}}{8\delta H(3\lambda-2)^{2}}(1-\phi)\Pi\geq 0,\end{split}$ (A.6)
which is always fulfilled in Regime 3. The strict equality holds at $t_{M}=t_{M}^{3}$ and it holds that $t_{M}>t_{M}^{2}>\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ for any $\lambda>1/3$. Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rates, as given in (A.4).
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rates in response to the havens’ split rates. Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates, as given in (A.5).
Non-haven Haven
single GMT rate split tax rate
single non-GMT rate
single GMT rate $G_{n}^{3},\quad G_{h}^{3}$
split tax rate $G_{n}^{4},\quad G_{h}^{4}$
Table A2: Payoff matrix for Regimes 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$) and 4 ($t_{M}\in(t_{M}^{3},1]$)
Notes: In Regimes 3 and 4, the GMT is binding for both the non-haven and the haven. Therefore, the regimes where the non-haven chooses a single-non GMT rate are irrelevant. The regimes where the haven sets a single GMT are also irrelevant because, by doing so, the haven would obtain non-positive welfare/tax revenues.
A.2. Decentralized decision of tax havens
In our main model, we have assumed that tax havens collectively decide on whether to set a single GMT rate, or to split their tax rate. We here examine an alternative, decentralized decision process in the first stage of our game where each single tax haven decides whether to split or not, given the other havens’ decisions. To simplify this analysis, we assume that countries can split their tax rate only if they are bound by the GMT. This reflects the fact that preferential tax regimes for specific tax bases are not allowed under the OECD’s action plan to fight base erosion.
Suppose then that $t_{M}\in[t_{M}^{0},t_{M}^{2}]$. Out of the total of $H$ tax havens, there are $H_{M}$ havens that commit to a single GMT rate while the remaining $H_{S}=H-H_{M}$ havens split their tax rate. When the non-haven sets a single non-GMT rate and a single haven $h$ sets a uniform GMT rate, the equilibrium tax rates are
$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{1}=t_{M},$ (A.7)
which is the counterpart of (A.1). Let $G_{h}^{1}(H_{M})$ be the haven’s equilibrium payoff with commitment.
If the single haven $h$ splits its tax rate instead, whereas the non-haven sets a uniform non-GMT rate, the equilibrium tax rates are
$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{2}=\begin{cases}t_{n}(H_{M})/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},$ (A.8)
which is the counterpart of (A.2). Let $G_{h}^{2}(H_{M})$ be the haven’s equilibrium payoff under splitting.
We then ask if a single haven has an incentive to choose a uniform GMT rate given that all the other $H-1$ havens split their tax rates. A single haven $h$ has a unilateral incentive to do so, if the following holds:
$\Delta G_{h}^{12}(H_{M}=1)$ $\equiv G_{h}^{1}(H_{M}=1)-G_{h}^{2}(H_{M}=0)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta}{\delta H^{2}[\lambda(3+\phi)-(1+\phi)]^{2}[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]}\geq 0,$ (A.9)
where $\Theta\equiv A-Bt_{M},\quad A\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
$B\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
$\text{and}\qquad Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0\ \ \text{or}\ \ t_{M}\geq\frac{\delta(\lambda-1)}{H(3\lambda-1)}\equiv t_{M}^{0}.$
It is straightforward to show that $\Theta>0$ holds at $t_{M}^{0}$ and hence, by continuity, also for $t_{M}$ marginally above $t_{M}^{0}$. Therefore, there must be a positive range of $t_{M}$, with $t_{M}^{0}<t_{M}\leq t_{M}^{a}\equiv A/B$ for which the inequality (A.9) holds for a haven $h$. By symmetry, this argument applies to all havens.
For a Regime 1 equilibrium with commitment to occur, a further condition is that no single haven $h$ must have an incentive to split, given that all the other $H_{M}-1$ havens commit to a single GMT rate. A single haven $h$ has no unilateral incentive to split its tax rate, if the following holds:
$\Delta G_{h}^{12}(H_{M}=H)$ $\equiv G_{h}^{1}(H_{M}=H)-G_{h}^{2}(H_{M}=H-1)$
$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta^{\prime}}{\delta H(2\lambda-1)[\phi(\lambda-1)+\lambda(4H-1)-(2H-1)]^{2}}\geq 0,$ (A.10)
where $\Theta^{\prime}\equiv C-Dt_{M},\quad C\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
$D\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
and $t_{M}\geq t_{M}^{0}$ holds again. The equivalent condition is $t_{M}\leq t_{M}^{b}\equiv C/D$. We can immediately see $\Delta G_{h}^{12}(H_{M}=H)<0$ for $t_{M}\in(t_{M}^{b},t_{M}^{2}]$ and check $t_{M}^{a}<t_{M}^{b}<t_{M}^{1}<t_{M}^{2}$:
$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{\Theta^{\prime\prime}[(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}]}>0,$
$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}}{\Theta^{\prime\prime}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0,$
$\Theta^{\prime\prime}\equiv 3(3H-1)\lambda^{2}-2(3H-2)\lambda+H-1+\phi(\lambda-1)[\lambda(3H-5)-(H-3)]>0.$
We can conclude that for $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, all the havens set a single GMT rate and the non-haven sets a single non-GMT rate; for $t_{M}\in(t_{M}^{a},t_{M}^{2}]$, all the havens split their tax rate and the non-haven sets a single non-GMT rate. The former regime corresponds to Regime 1 and the latter one to Regime 2 in the text. The only difference is that the range of Regime 1 is smaller when havens make their first stage decision in a decentralized way, implying that the havens’ commitment is less likely. This is intuitive because the commitment of a single haven does not induce the non-haven to increase its tax rate by as much as the simultaneous commitment of many havens. Finally, in the remaining range of the GMT, $t_{M}\in(t_{M}^{2},1]$, the difference between collective vs. decentralized decisions of havens plays no role any more. In either setting, the tax havens always have an incentive to split their tax rate, whether the non-haven sets a single GMT rate (Regime 3), or splits its tax rate as well (Regime 4).
A.3. Proofs of Propositions 3 and 4
Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
Comparing the welfare levels of countries in Regime 1 with those in Regime 0 gives
$G_{n}^{1}-G_{n}^{0}=\frac{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[Ht_{M}\lambda^{2}(3\lambda-1)+\delta(\lambda-1)(7\lambda^{2}-8\lambda+2)\right]}{2\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi>0,$
$G_{h}^{1}-G_{h}^{0}=\frac{\lambda\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(2\lambda-1)-Ht_{M}\lambda(3\lambda-1)\right]}{\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi\equiv\Delta$
$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\end{cases},$
$G_{W}^{1}-G_{W}^{0}=\frac{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)-Ht_{M}\lambda^{2}(3\lambda-1)\right]}{2\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi>0,$
where the last inequality holds from
$t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)}{H\lambda^{2}(3\lambda-1)}.$
The marginal effect of the GMT is
$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,\qquad\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)}{2H\lambda}\end{cases},$
$\frac{\partial G_{W}^{1}}{\partial t_{M}}=\frac{\delta(\lambda-1)(2\lambda-1)-H\lambda^{2}t_{M}}{\delta(2\lambda-1)}\Pi>0,$
where the last inequality holds because of
$t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(2\lambda-1)}{2H\lambda}.$
In Regime 1, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven rises if $t_{M}<\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$ after the GMT introduction and it falls if $t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$.
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
For the non-haven country, we see
$\frac{\partial G_{n}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{1}}=\frac{(\lambda-1)[8\lambda^{3}-5\lambda^{2}-2\lambda+1+\phi(\lambda-1)^{2}]}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi>0,$
$\frac{\partial G_{n}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)[8\lambda^{2}-5\lambda+1+\phi(\lambda-1)]}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi>0.$
From these results and the fact that $G_{n}^{2}$ is a quadratic function of $t_{M}$, it follows that $G_{n}^{2}$ increases with $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{n}^{2}-G_{n}^{0}>0$, it is sufficient to show
$G_{n}^{2}-G_{n}^{0}\big{|}_{t_{M}=t_{M}^{1}}=\frac{\delta(\lambda-1)^{3}[(3\lambda-1)(16\lambda^{3}-13\lambda^{2}+1)-\phi(\lambda-1)^{3}(8\lambda-3)]}{2H(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi>0,$
which holds for any $\lambda>1$ and $\phi\in(0,1)$.
For the haven country, we get
$\frac{\partial G_{h}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{1}}=-\frac{2\lambda(2\lambda-1)(\lambda-1)^{2}}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi<0,$
$\frac{\partial G_{h}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{2}}=-\frac{4\lambda^{2}(\lambda-1)}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi<0.$
From these results and the fact that $G_{h}^{2}$ is a quadratic function of $t_{M}$, $G_{h}^{2}$ must be decreasing in $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{h}^{2}-G_{h}^{0}>0$, it is sufficient to show
$G_{h}^{2}-G_{h}^{0}\big{|}_{t_{M}=t_{M}^{1}}=-\frac{\delta\lambda(\lambda-1)^{5}[3\lambda-1+\phi(\lambda-1)]}{H(3\lambda-1)^{2}(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi<0,$
which holds for all $\lambda>1$ and $\phi\in(0,1)$.
For the world, we see
$\frac{\partial G_{W}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{1}}=\frac{(\lambda-1)[4\lambda^{3}+\lambda^{2}-4\lambda+1+\phi(\lambda-1)^{2}]}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi>0,$
$\frac{\partial G_{W}^{2}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)^{2}(4\lambda-1+\phi)}{[3\lambda-1+\phi(\lambda-1)][3\lambda-1-\phi(\lambda-1)]}\phi\Pi>0.$
Hence $G_{W}^{2}$ must be increasing in $t_{M}$ for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$. To prove $G_{W}^{2}-G_{W}^{0}>0$, we must show
$G_{W}^{2}-G_{W}^{0}\big{|}_{t_{M}=t_{M}^{1}}=\frac{\delta(\lambda-1)^{3}[(3\lambda-1)(14\lambda^{3}-9\lambda^{2}-2\lambda+1)-\phi(\lambda-1)^{3}(10\lambda-3)]}{2H(3\lambda-1)^{2}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]^{2}}\phi\Pi>0,$
which holds for all $\lambda>1$ and $\phi\in(0,1)$.
In Regime 2, the welfare levels of the non-haven and the world therefore rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven country, we have
$\frac{\partial G_{n}^{3}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)[2\lambda-1+\phi(2\lambda+1)]}{2[3\lambda-1-\phi(\lambda-1)]}\Pi>0,$
$\frac{\partial G_{n}^{3}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{3}}=\frac{(\lambda-1)[\phi(8\lambda-3)-(2\lambda-1)]}{2(3\lambda-1)}\Pi>0,$
so that $G_{n}^{3}$ must increase with $t_{M}$ for $t_{M}\in(t_{M}^{2},t_{M}^{3}]$. To prove $G_{n}^{3}-G_{n}^{0}>0$, we show that the following must hold for all levels of $\lambda>1$ and $\phi\in(0,1)$:
$G_{n}^{3}-G_{n}^{0}\big{|}_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(16\lambda^{3}-13\lambda^{2}+3)-\phi(\lambda-1)^{2}(8\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
For the haven country, we see
$\frac{\partial G_{h}^{3}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{2}}=\frac{\lambda(\lambda-1)}{3\lambda-1-\phi(\lambda-1)}(1-\phi)\Pi>0,$
$\frac{\partial G_{h}^{3}}{\partial t_{M}}\bigg{|}_{t_{M}=t_{M}^{3}}=\frac{\lambda(\lambda-1)(8\lambda-3)}{(3\lambda-1)(4\lambda-1)}(1-\phi)\Pi>0,$
which implies that $G_{h}^{3}$ decreases with $t_{M}$ for $t_{M}\in(t_{M}^{2},t_{M}^{3}]$. To prove $G_{h}^{3}-G_{h}^{0}<0$, we must show that
$G_{h}^{3}-G_{h}^{0}\big{|}_{t_{M}=t_{M}^{3}}=\frac{\delta\lambda(\lambda-1)^{2}[8(2\lambda-1)(3\lambda-1)-\phi(8\lambda-3)^{2}]}{H(3\lambda-1)^{2}(4\lambda-1)^{2}}\Pi<0.$
This inequality holds under our assumptions of $\lambda>1$ and $\phi\in(3/4,1)$, noting that $\phi>\frac{3}{4}>\frac{8(2\lambda-1)(3\lambda-1)}{(8\lambda-3)^{2}}$.
Considering that both the non-haven and the haven increases their welfare along with a gradual increase in $t_{M}$, world welfare must also increase as $t_{M}$ rises. To prove $G_{W}^{3}-G_{W}^{0}>0$, we show that, for all $\lambda>1$ and $\phi\in(0,1)$, it holds that
$G_{W}^{3}-G_{W}^{0}\big{|}_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(14\lambda^{2}-15\lambda+3)-\phi(\lambda-1)^{2}(10\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
In Regime 3, the welfare levels of the non-haven, the haven, and the world all rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
For the non-haven country, we see
$G_{n}^{4}-G_{n}^{0}\big{|}_{t_{M}=t_{M}^{3}}=\frac{\delta(\lambda-1)^{2}(8\lambda-3)^{2}}{2H(3\lambda-1)^{2}(4\lambda-1)}\phi\Pi>0,$
for $\lambda>1$ and $\phi\in(0,1)$. Moreover, the marginal effect of the GMT is always positive: $\partial G_{n}^{4}/\partial t_{M}=(\lambda-1)\phi\Pi>0$. With these observations and the fact that $G_{n}^{4}-G_{n}^{0}$ is linear in $t_{M}$, we conclude that $G_{n}^{4}-G_{n}^{0}>0$ for $t_{M}\in(t_{M}^{3},1]$.
For the haven country, we have
$G_{h}^{4}-G_{h}^{0}=-\frac{\delta\lambda(\lambda-1)^{2}}{H(3\lambda-1)^{2}}\phi\Pi>0,$
which holds for $\lambda>1$ and $\phi\in(0,1)$. The marginal effect of the GMT is zero: $\partial G_{h}^{4}/\partial t_{M}=0$ for $t_{M}\in(t_{M}^{3},1]$. Combining these results leads to $G_{W}^{4}-G_{W}^{0}>0$ and $\partial G_{W}^{4}/\partial t_{M}>0$.
In Regime 4, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate. On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
This proves Proposition 3 and a part of Proposition 4. The proof of the results on the regime-switching points is given in Appendix 3.
A.4. Regime-switching points
At the three regime switching GMT rates, $t_{M}^{1}$, $t_{M}^{2}$ and $t_{M}^{3}$, welfare in the non-haven, the haven, and the world changes as follows.
Change from Regime 1 to 2 at $t_{M}^{1}$
$G_{n}^{2}-G_{n}^{1}\big{|}_{t_{M}=t_{M}^{1}}=G_{W}^{2}-G_{W}^{1}\big{|}_{t_{M}=t_{M}^{1}}$
$=-\frac{\delta(\lambda-1)^{3}[\lambda(3\lambda-1)(3\lambda-2)-\phi(\lambda-1)^{2}]}{2H(\lambda-1)[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}(1-\phi)\Pi<0,$
$G_{h}^{2}-G_{h}^{1}\big{|}_{t_{M}=t_{M}^{1}}=0.$
Change from Regime 2 to 3 at $t_{M}^{2}$
$G_{n}^{3}-G_{n}^{2}\big{|}_{t_{M}=t_{M}^{2}}=G_{h}^{3}-G_{h}^{2}\big{|}_{t_{M}=t_{M}^{2}}=G_{W}^{3}-G_{W}^{2}\big{|}_{t_{M}=t_{M}^{2}}=0.$
Change from Regime 3 to 4 at $t_{M}^{3}$
$G_{n}^{4}-G_{n}^{3}\big{|}_{t_{M}=t_{M}^{3}}=0,$
$G_{h}^{4}-G_{h}^{3}\big{|}_{t_{M}=t_{M}^{3}}=G_{W}^{4}-G_{W}^{3}\big{|}_{t_{M}=t_{M}^{3}}=-\frac{8\delta\lambda(\lambda-1)^{2}(2\lambda-1)}{2H(\lambda-1)[H(3\lambda-1)(4\lambda-1)^{2}]}(1-\phi)\Pi<0.$
This completes the proof of Proposition 4.
A.5. Proof of Proposition 5
The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
We see
$\frac{\partial G_{n}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{h}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{W}^{1}}{\partial\phi}=0.$
Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
For the non-haven country, we see
$\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\Theta_{n}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{2\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{n}\equiv\phi[Ht_{M}(\lambda-1)(16\lambda^{2}-13\lambda+3)-\delta(\lambda-1)^{2}]+Ht_{M}(3\lambda-1)^{2}+\delta(\lambda-1)(16\lambda^{2}-19\lambda+5).$
The sign of the derivative is determined by that of $\Theta_{n}$. From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{1})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
For the haven country, we see
$\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\lambda\Theta_{h}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{h}\equiv\phi[Ht_{M}(\lambda-1)(5\lambda-3)-\delta(\lambda-1)^{2}]-Ht_{M}(3\lambda-1)^{2}+\delta(\lambda-1)(5\lambda-3).$
The sign of the derivative is determined by that of $\Theta_{h}$. From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{1})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
For the world, we get
$\frac{\partial G_{W}^{2}}{\partial\phi}=\frac{\Theta_{W}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{2\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
$\Theta_{W}$ $\equiv\phi[Ht_{M}(\lambda-1)(2\lambda-1)(13\lambda-3)-\delta(\lambda-1)^{2}(2\lambda+1)]-Ht_{M}(3\lambda-1)^{2}(2\lambda-1)$
$+\delta(\lambda-1)(26\lambda^{2}-25\lambda+5).$
The sign of the derivative is determined by that of $\Theta_{W}$. From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{1})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$. We see
$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\quad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\quad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
We get
$\frac{\partial G_{n}^{4}}{\partial\phi}=\frac{(\lambda-1)[2Ht_{M}(3\lambda-1)^{2}-\delta(\lambda-1)(8\lambda-3)]}{2H(3\lambda-1)^{2}}\Pi>0,$
$\frac{\partial G_{h}^{4}}{\partial\phi}=-\frac{\delta\lambda(\lambda-1)^{2}}{H(3\lambda-1)^{2}}\Pi<0,$
$\frac{\partial G_{W}^{4}}{\partial\phi}=\frac{(\lambda-1)[2Ht_{M}(3\lambda-1)^{2}-\delta(\lambda-1)(10\lambda-3)]}{2H(3\lambda-1)^{2}}\Pi>0,$
noting that $t_{M}>t_{M}^{3}>\frac{\delta(\lambda-1)(10\lambda-3)}{2H(3\lambda-1)^{2}}>\frac{\delta(\lambda-1)(8\lambda-3)}{2H(3\lambda-1)^{2}}$.
This completes the proof of Proposition 5.
Appendix B: Data for Figure 1
We select MNEs from the Orbis based on the following criteria.
  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
The numbers on which Figure 1 is based are given in Table B1. In column A, to calculate the pre-tax profits of all MNEs in a given year, we exclude those with missing revenues (“Operating revenue (Turnover)”) and pre-tax profits (“P/L before tax”). In column B, MNEs subject to the GMT are those with annual revenues of no less than 750 million EUR in at least two years of the last four years. Using column A and B, column C reports the share of MNEs subject to the GMT in terms of pre-tax profits and numbers.
Table B1: MNEs in the Orbis database
A: All MNEs B: MNEs $\geq$ 750mEUR C: Share $=$ B/A
Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
2018 2320 8656 2066 2333 0.89 0.27
2019 2430 8416 2216 2437 0.91 0.29
2020 1985 7920 1793 2228 0.90 0.28
2021 3564 8803 3239 2627 0.91 0.30
Source: Orbis database, own calculations.
Note: Pre-tax profits are in billion EUR.

### v6

@@
 A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
-Abstract
+A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
+IMPLICATIONS FOR TAX COMPETITION***Andreas Haufler, Seminar for Economic Policy, LMU Munich, Akademiestr.
+1, 80799 Munich, Germany.
+Email: Andreas.Haufler@econ.lmu.de.
+Hayato Kato, Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan.
+Email: hayato.kato@econ.osaka-u.ac.jp.
+ Financial support from the German Research Foundation through CRC TRR 190 (Haufler) and the JSPS KEKENHI (JP20H01495, JP22K13388, JP24H00014) (Kato) is gratefully acknowledged.
+This study is conducted as a part of the Project “Economic Policy Issues in the Global Economy” undertaken at the Research Institute of Economy, Trade and Industry (RIETI).
+Kato thanks the University of Munich for their hospitality while part of this paper was written.
+We thank Tomohiro Ara, Real Arai, Ruby Doeleman, Antoine Ferey, Emanuel Hansen, Jim Hines, Niels Johannesen, Makoto Hasegawa, Jota Ishikawa, Mick Keen, Sang-Hyun Kim, Yoshimasa Komoriya, Jonas Löbbing, Mohammed Mardan, Jakob Miethe, Kenichi Nishikata, Hirofumi Okoshi, Pascalis Raimondos, Dirk Schindler, Karl Schulz, Georg Thunecke, and Alfons Weichenrieder for many helpful comments and discussions.
+Thanks also to seminar and conference participants at the IIPF in Utah State, MPI Berlin, Goettingen, Osaka, Kyoto, Shimane, QUT Brisbane, Ryukoku, Gakushuin, CESifo (Munich & Venice Summer Institute), and RIETI for helpful comments.
+Andreas Haufler and Hayato Kato
 The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms.
-We analyze tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT.
+We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT.
 Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals.
 However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals.
-Our calibration exercise shows that the implementation of a 15% GMT rate results in a regime where only the havens adopt split tax rates.
-Upon GMT introduction, welfare and tax revenues fall in the tax havens but rise in the non-haven, yielding a positive net gain worldwide.
-Keywords: multinational firms; tax avoidance; global minimum tax; profit shifting
-JEL classification: F23, H25, H87
-1 Introduction
-Profit shifting by multinational enterprises (MNEs) has long been a major problem for corporate taxation worldwide.
-Using aggregate data, Tørsløv et al., (2023) estimate that more than one-third of all foreign-earned corporate profits of multinationals are shifted to tax havens.
-The important role of tax havens for profit shifting is confirmed in other studies using microdata of multinational affiliates/headquarters in France (Davies et al.,, 2018), the UK (Bilicka,, 2019), and the US (Guvenen et al.,, 2022).
-In response to the large revenue losses caused by profit shifting, the OECD has launched an action plan to fight base erosion in OECD countries, and in particular the profit shifting to tax havens (OECD,, 2013).
-A core development in this endeavor is the introduction of a global minimum tax (GMT), which is applied to large MNEs (OECD, 2020a, ; OECD, 2020b, ).
-The GMT with a tax rate of 15% has been agreed upon by a group of more than 130 countries in 2021 including many tax havens, and most of these countries are expected to have the GMT enacted by 2026.111The list of countries joining the agreement on the GMT, also known as Pillar 2, is found in OECD, (2021).
-The United States has already enacted, in its 2017 Tax Cuts and Jobs Act, a tax on Global Intangible Low-Tax Income (GILTI) that works in many respects like a GMT (see Chodorow-Reich et al.,, 2024 for details).
-The global revenue gains from a 15% GMT are estimated to be in the range of 155 to 192 billion USD, or 6.5 to 8.1% of global corporate tax revenues (Hugger et al.,, 2024, Section 8.2).
-[^2] Baraké et al., (2022, Table 1) break down the global revenue changes of the GMT by country and arrive at a revenue gain of around 55 to 67 billion EUR for the US, and a gain of similar magnitude for the EU as a whole.
-Estimates of both Hugger et al., (2024) and Baraké et al., (2022) are based on models where governments and multinational firms mechanically respond to the GMT.
-Wang, (2020) sets up a quantitative corporate tax framework that endogenizes the behavior of both firms and governments.
-Using similar frameworks, Ferrari et al., (2023) and Shen, (2024) quantify the effects of the GMT.
-While no econometric estimates exist for the global minimum tax yet, there are experiences from similar policies enacted at the national level.
-See Buettner and Poehnlein, (2024) for an analysis of the effects of a minimum tax rate for German municipalities.
-An important limitation of the GMT is, however, that it applies only to “large” multinationals, defined as multinational groups that have had no less than 750 million EUR in total annual revenues in at least two out of the last four years.
-[^3] Since 2016, multinational groups above this threshold size are also required to file individual reports on their activities, profits, and taxes paid in each of the countries in which they have a presence (country-by-country reporting).
-Fig.
-1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of all MNEs, which are responsible for about 90% of all MNE profits, have consolidated accounts above this threshold and are therefore covered by the GMT.
-[^4] This corresponds to the estimate of the OECD, which arrives at a coverage rate of 90% of profits, using the Orbis database and other data sources (OECD, 2020a, , # 505, p.
+Our calibration exercise shows that introducing a GMT rate of 15% results in a regime where only the havens adopt split tax rates.
+Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
+Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation.
+Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., , 2023), a pattern also documented in microdata for France (Davies et al., , 2018), the UK (Bilicka, , 2019), and the US (Guvenen et al., , 2022) In response to the resulting revenue losses, the OECD launched its Base Erosion and Profit Shifting (BEPS) initiative (OECD, , 2013), whose core outcome is the introduction of a global minimum tax (GMT), also known as Pillar Two.
+[^1] See Fuest and Neumeier, (2023); Hebous et al., (2024) for the institutional background.
+Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., , 2025; Ogawa and Tsuchiya, , 2025; Richter, , 2025).
+In 2021, more than 130 countries including many tax havens agreed on a 15% GMT; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT.
+[^2] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed October 1, 2025).
+Despite announcing after the June 2025 G7 summit that it would not fully comply with the GMT, the US had already implemented a similar tax on Global Intangible Low-Tax Income under the 2017 Tax Cuts and Jobs Act (Chodorow-Reich et al., , 2024).
+A limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years.
+[^3] Since 2016, such groups have been required to file country-by-country reports on their activities, profits, and taxes in each country of operation (Doeleman et al., , 2024, Section 2).
+The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, , 2022, p.14).
+According to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT.
+[^4] See Supplemental Appendix A.
+This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, , 2020, # 505, p.
 233).
-Nevertheless, a substantial amount of MNE profits exceeding 300 billion EUR remain outside the scope of the GMT.
-Moreover, the MNE sample in Orbis is not comprehensive and the database is known for oversampling large firms (Bajgar et al.,, 2020).
-Therefore, a coverage rate of the GMT of 90% of all MNE profits can be considered as an upper bound of the true profit share covered by the GMT.
-One implication of this incomplete coverage is that low-tax countries might respond to the introduction of the GMT by using a split corporate tax system, where a lower tax rate applies to firms below the GMT threshold.
-In fact, such a split is inherent in the regulation that tax havens are allowed to implement the GMT by raising their (low) general tax rate only for MNEs above the GMT threshold through a specific top-up tax.
-[^5] This is known as the Qualified Domestic Marginal Top-Up Tax (QDMTT).
-See Devereux, (2023) for a discussion.
-Tax havens are fully aware of this option.
-Ireland and Liechtenstein, for example, have already decided to keep their general tax rates at 12.5%, but top up the tax rate to 15% for affiliates of foreign MNEs above the threshold (Government of Ireland,, 2023; PwC,, 2024 for Liechtenstein).
-Hungary and Bulgaria have adopted a similarly split tax rate by joining the GMT agreement while maintaining their regular tax rates of 9% and 10%, respectively, for firms not covered by the GMT (Ernst & Young,, 2023 for Hungary; PwC,, 2023 for Bulgaria).
-[^6] The split corporate tax system is also expected to be introduced in Bermuda and Jersey (Government of Bermuda,, 2024; Government of Jersey,, 2023).
-These responses indicate that the effects of the GMT cannot be analyzed assuming a single, uniform corporate tax rate across all countries.
-Figure 1: GMT coverage
-Source: Orbis database, own calculations.
-Notes: The left bar in each year shows the share of MNEs covered by the GMT out of all MNEs.
-The right bar in each year shows the share of pre-tax profits earned by MNEs covered by GMT out of pre-tax profits by all MNEs.
-The MNEs covered by GMT are those whose annual revenues exceed $750$ million EUR in at least two of the last four years out of all MNEs.
-See Appendix B for details.
-Against this background, this paper analyzes under which conditions tax competition leads countries to differentiate their tax rates for large and small MNEs in response to the introduction of a GMT.
-We also study the welfare implications of such split tax system.
-Finally, we explore whether gradual reforms to the GMT—either through an increase in the GMT rate or an expansion of its coverage—enhance welfare for each country and for the world as a whole.
-To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits from there to a set of symmetric tax-haven countries.
-Countries maximize their welfare, a weighted sum of tax revenues and (for the non-haven) the post-tax income of MNEs’ owners, and they compete for the MNEs’ profits under the GMT with incomplete coverage.
-Both the non-haven and the havens can decide to commit, in the first stage, to set a uniform tax rate for all their multinationals, or not.
-In the second stage, all countries choose their tax rates, either uniform or split, where the tax rate applicable for large MNEs must be at or above the GMT level.
-In the third stage, multinationals make their profit-shifting choices.
-Our analysis shows that the introduction of a partial GMT leads to a sequence of tax competition regimes between the non-haven and the havens.
-If the GMT rate is binding for the tax havens, but still low enough, each of the tax havens will choose to commit to the GMT rate for all MNEs, in order to increase the (uniform) tax rate in the non-haven country.
-As the GMT rate is further increased, however, the non-cooperative tax equilibrium regime transitions to one where the havens split their tax rates by setting a GMT for large MNEs and a rate lower than the GMT for small MNEs.
-For sufficiently high GMT rates, a tax regime eventually emerges where the non-haven splits its tax rate as well.
-This is because a higher GMT rate limits competition for profits from large MNEs, leading the countries to focus instead on profits from small MNEs.
-The implications for welfare and tax revenues are as follows.
-At the regime-switching GMT rates, if one country starts to split its tax rate, the other country discretely loses welfare and tax revenues.
-In general, however, the non-haven always gains from gradual increases in the GMT rate, whereas the havens weakly gain only for sufficiently low levels of the GMT, and lose otherwise.
-Gradual increases in the coverage rate of the GMT have similar effects, with the important difference that they never reduce tax revenues in the non-haven country or globally.
-These results suggest that, although both countries benefit from the introduction of a moderate GMT rate, conflicts of interest will arise once the GMT rate or coverage is increased.
-In the final step, we calibrate our model using estimates of profit shifting from a global sample of countries (Tørsløv et al.,, 2023).
-In our benchmark calibration, the introduction of a $15\%$ GMT rate causes a regime shift: tax haven countries adopt split tax rates, while the non-haven maintains a uniform tax rate above the GMT.
-This result aligns with the actual responses observed in several tax haven countries, as described above.
-Regarding welfare and tax revenues, the introduction of the GMT benefits the non-haven while harming havens, leading to net welfare gains worldwide.
-This conflict of interest between the two groups of countries persists in the two gradual reforms likely to follow: increasing the GMT rate from $15\%$ to $18\%$ and expanding GMT coverage to all MNEs.
-However, the conflict is less pronounced for the latter reform.
-Our model contributes to the recent theoretical literature on the effects of tax competition in the presence of a GMT.
-In a model with many non-haven and haven countries, Johannesen, (2022) shows that introducing a GMT will always benefit the tax havens, because the GMT reduces the competition between them.
-In contrast, the welfare effects of the GMT are generally ambiguous for the non-haven country, because of the offsetting changes in tax revenues and firms’ post-tax profits.
-Hebous and Keen, (2023) also analyze the effects on welfare in a model where two countries differ in terms of both their size and their valuation of public spending.
-Their quantitative results indicate that GMT levels up to a range of 17 to 20% can constitute strict Pareto improvements, increasing welfare in both high-tax and low-tax countries.
-Hindriks and Nishimura, (2022) consider governments who choose not only tax rates but also tax enforcement levels.
-They show that the introduction of the GMT hinders tax enforcement cooperation between the high-tax and the low-tax country, making the high-tax country worse off.
-Finally, Janeba and Schjelderup, (2023) study a three-country model where two non-haven countries compete for real foreign direct investment (FDI) while simultaneously competing for profit shifting with a third, tax-haven country.
-In this setting the introduction of the GMT intensifies tax competition for real FDI.
-While tax revenues in the non-haven countries will still increase if countries compete via corporate tax rates, net revenue gains are zero if competition is via lump-sum subsidies.
-[^7] Using an international trade model of firms without profit-shifting motives, Nocco et al., (2024) examine a rich set of corporate taxes and subsidies on sales, production, and profits.
-Specifically, they identify the globally welfare-maximizing combination of non-discriminatory taxes and subsidies including a global minimum tax on firms’ profits.
-All these models have in common that the GMT must be levied uniformly on all firms, thus ignoring the existence of a threshold below which the GMT does not apply.
-In contrast, our model highlights the critical role of this threshold, as it allows countries to use different tax rates on firms with revenues above and below the threshold.
-This feature links our model to the literature on discriminatory tax competition, which examines whether tax revenues are higher in the non-cooperative tax equilibrium when countries may, or may not, set different tax rates on tax bases with different degrees of international mobility (Janeba and Peters,, 1999; Keen,, 2001; Janeba and Smart,, 2003).
-[^8] This literature has been extended to cover differences in country size (Bucovetsky and Haufler,, 2007), and imperfect competition in product markets (Gaigné and Wooton,, 2011).
-A related type of discrimination arises when profit shifting is monitored by governments in a deliberately loose way (Peralta et al.,, 2006).
-In our setting, tax discrimination will not be profitable for either country in the absence of the GMT, but it will arise in equilibrium if a sufficiently high level of the GMT is imposed.
-A distinct, but closely related literature has addressed the issue of whether the existence of tax havens—by allowing non-haven countries to tax-discriminate in favor of mobile, multinational firms—raises or reduces tax revenues and welfare in the non-haven countries (Slemrod and Wilson,, 2009; Hong and Smart,, 2010; Johannesen,, 2010; Elsayyad and Konrad,, 2012).
-Our analysis proceeds as follows.
-In Section 2 we set up our model of tax competition and characterize the sub-game perfect Nash equilibria that result in different tax regimes.
-Section 3 then turns to the welfare effects of introducing a GMT, and of increasing its tax rate and coverage.
-Section 4 calibrates our model to real-world data and quantifies its effects.
-Section 5 concludes.
+Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope.
+Moreover, Orbis tends to oversample large firms (Bajgar et al., , 2020), so the 90% coverage rate should be viewed as an upper bound.
+Does this incomplete coverage affect the effectiveness of the GMT?
+It does not if low-tax countries apply the GMT rate to all firms regardless of size.
+It does matter, however, if they differentiate their tax rates, applying the GMT rate to firms above the threshold and a lower rate to those below.
+Such a split is inherent in the regulation: tax havens may implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a top-up tax.
+[^5] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT) (Devereux, , 2023).
+This system has been adopted in Bahrain, Ireland, Switzerland and the UAE, among other countries with base statutory corporate tax rates below 15% (PwC, , 2025).
+Against this background, we ask when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT.
+Does a 15% GMT rate induce such a split, and how large are the resulting revenue and welfare gains?
+How do these outcomes change with broader GMT coverage?
+To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries.
+Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits.
+In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rates for all MNEs.
+In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate.
+The final stage involves MNEs’ profit-shifting decisions.
+Our analysis shows that introducing the GMT with partial coverage generates a sequence of non-cooperative tax equilibrium regimes between the non-haven and the havens.
+When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate.
+As the GMT rate rises, the havens begin to split their rates, applying the GMT to large MNEs and a lower rate to small MNEs.
+For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
+We calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms.
+A 15% GMT rate induces the havens to split their tax rates, which is consistent with observed responses in some tax havens including Ireland and Switzerland.
+It benefits the non-haven while harms the havens, generating net global revenue (welfare) gains of 136 billion USD (83 bUSD), comparable to existing estimates (Baraké et al., , 2022; Hugger et al., , 2024).
+If the GMT covers all firms, the 15% rate does not produce a split tax system but increases revenues and welfare in the non-haven and globally even more.
+This global revenue gain would have to be set against the additional compliance costs for multinationals, however.
+We contribute to the recent theoretical studies on tax competition under a GMT.
+They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, , 2022; Hebous and Keen, , 2023); real economic activities (Janeba and Schjelderup, , 2023; Schjelderup and Stähler, , 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, , 2025).
+All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT.
+Our model highlights this threshold, allowing countries to tax firms above and below it differently.
+[^6] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, , 1999; Keen, , 2001; Bucovetsky and Haufler, , 2007; Gaigné and Wooton, , 2011).
+Our simple quantitative exercise complements recent evaluations of corporate taxation using trade and multinational models (Wang, , 2020; Ferrari et al., , 2023; Shen, , 2024; Dyrda et al., , 2024), where taxes are assumed to be exogenous and/or uniform across firms.
+By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
+Two empirical papers (Bachas et al., , 2025; Clifford et al., , 2025) closely related to ours examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively.
+Their findings that GMT revenue gains depend strongly on its coverage emphasizes the relevance of our analysis.
 1 The model
 1.1 Setup
-We consider a tax competition model between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$.
-Tax competition arises between each of the tax havens and the non-haven.
-There are a large number of heterogeneous, multinational enterprises (MNEs) owned by individuals in the non-haven, each having a headquarters in the non-haven and an affiliate in each of the tax havens.
-The headquarters in the non-haven makes exogenous profits in this country, $\pi$, which are distributed in the range $[\underline{\pi},\infty)$, with a cumulative distribution function $F(\pi)$.
-Hence, all real activity occurs in the non-haven, and this is where the “true” profits accrue in its entirety.
-MNEs can shift a share $\theta\in[0,1]$ of their non-haven headquarters’ profits to the set of tax haven affiliates, in order to maximize their post-tax profits.
-[^9] This is the simplest possible setting for profit shifting.
-See e.g., Johannesen, (2022) or Krautheim and Schmidt-Eisenlohr, (2011) for a similar assumption.
-In particular, this setting avoids the possible non-existence of Nash equilibria that arise in models where the location of heterogeneous firms is endogenous (Baldwin and Okubo,, 2009; Davies and Eckel,, 2010; Haufler and Stähler,, 2013).
-We assume that the non-haven and all tax havens have signed the GMT agreement, and are therefore bound to set tax rates of at least the GMT rate, $t_{M}$.
-[^10] Devereux, (2023) discusses how the specific institutional setting of the GMT, in particular the Under-Taxed Payments Rule (UTPR), gives incentives to both non-haven countries and haven countries to join the GMT agreement, once this has been initiated by a critical number of large non-havens.
-Note also that we assume the tax base of the GMT to be the same as that of national corporate taxes, thus ignoring extra tax deductions (labelled “substance-based income exclusion”) under the GMT.
-See Schjelderup and Stähler, (2024) for a study focusing on the latter issue.
-However, the GMT applies only to MNEs that make exogenous profits no smaller than $\pi_{M}$, called “large MNEs” in our analysis.
-Therefore, countries remain free to set a lower tax rate than the GMT for all MNEs that make profits less than $\pi_{M}$, called “small MNEs.”
-The non-haven country maximizes a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of the heterogeneous multinationals, which are all owned by residents of the non-haven country.
-Since tax havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
-We consider a three-stage game.
-In the first stage, the non-haven and haven countries simultaneously choose whether to commit to a single tax rate for all MNEs, or not.
-[^11] This is related to the literature on endogenous timing, or leadership, in tax competition, where countries commit in a pre-play stage to set taxes early or late.
-See Kempf and Rota-Graziosi, (2010).
-To comply with the GMT agreement, this single tax rate has to be equal to, or higher than, the GMT rate.
-In our main model, we assume that the set of tax havens collectively decides on whether to set a single GMT rate, or split their tax rates.
-This assumption simplifies the algebra, but it does not change the qualitative nature of our results.
-[^12] We discuss the alternative of a decentralized first stage decision of the tax havens in Appendix A.2.
-In the second stage, all countries choose their tax rate(s) in a non-cooperative way.
-Countries that have committed to a single tax rate in the first stage set a single tax rate, whereas countries that have not committed choose different tax rates for large and small MNEs (where the tax rate on large MNEs has to be at least $t_{M}$).
-In the final stage, MNEs engage in profit shifting, given the tax rates in the non-haven and the symmetric tax havens.
-We solve this three-stage game by backward induction.
-The equilibrium concept we rely on is the sub-game perfect Nash equilibrium, which we call the tax-competition equilibrium.
+We consider tax competition between a non-haven country, indexed by $n$, and a finite set of $H$ symmetric tax-haven countries, indexed by $h$.
+There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens.
+Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven.
+[^7] Firm heterogeneity could result from (executives’) productivity (e.g., Melitz, , 2003; Lippi and Schivardi, , 2014).
+MNEs can shift some of these profits to the havens to maximise their post-tax profits.
+In the following, we assume collective decision-making by the havens.
+This assumption is relaxed in the quantitative analysis later (see also footnote 16) and further examined in Supplemental Appendix B.
+We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$.
+[^8] Devereux, (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates.
+The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
+The non-haven maximises a welfare function that consists of a weighted sum of tax revenues and the post-tax profits of MNEs, which are all owned by residents of the non-haven.
+Since havens do not have any ownership in the firms’ profits, the welfare of tax havens corresponds to their tax revenues.
+We analyse a three-stage game.
+In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs—at or above the GMT rate—or to allow differentiated rates.
+[^9] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 6 and in related work on endogenous timing in games (Huck et al., , 2002; Kempf and Rota-Graziosi, , 2010).
+The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates.
+The havens decide collectively on whether to unify or split their rates.
+In the second stage, countries set their tax rates non-cooperatively.
+Those that committed choose a single rate; others set different rates for large and small MNEs.
+In the final stage, MNEs shift profits given these tax rates.
+The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium, referred to as the tax-competition equilibrium.
 1.2 Profit shifting by multinationals
-Each MNE shifts a share $\theta_{h}$ of its exogenous pre-tax profits $\pi$ from their non-haven affiliate to each of the tax haven affiliates $h$.
-Firms incur transaction costs when shifting profits.
-These costs can either be thought of as concealment costs, which have to be incurred to hide the profit shifting from the non-haven country’s tax authorities, or as the expected fine that is to be paid when profit-shifting is detected and sanctioned.
-We assume that the costs of shifting $\theta_{h}\pi$ of profits to each haven $h$, given by $C_{h}(\cdot)$, are convex and, for analytical simplicity, quadratic in the level of profit-shifting, $\theta_{h}\pi$.
-Higher total (true) profits make it easier, however, to hide a given amount of profit-shifting.
-[^13] This specification is widely used in the literature on profit shifting (e.g., Hines and Rice,, 1994; Huizinga and Laeven,, 2008; Suárez Serrato,, 2018).
-Therefore, the shifting costs for a firm with profits $\pi$ to a haven $h$ are $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$.
-The exogenous parameter $\delta$ captures the ease with which profits can be shifted across countries.
-It incorporates any efforts that countries take to prevent profit-shifting.
-[^14] In Hindriks and Nishimura, (2022), this cost parameter is endogenously chosen by governments, either cooperatively or non-cooperatively.
-Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the tax haven countries, a MNE maximizes its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\Pi}$, by choosing the share of profit shifting to each haven:
+Each MNE shifts a share $\theta_{h}$ of its pre-tax profits $\pi$ from its non-haven headquarters to each tax haven affiliate $h$, and incurs transaction costs $C_{h}(\cdot)$ for doing so.
+These costs can be interpreted as concealment costs or as expected fines, if detected.
+We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$.
+[^10] This specification is standard in the literature (e.g., Hines and Rice, , 1994; Huizinga and Laeven, , 2008; Suárez Serrato, , 2018; Bilicka et al., , 2024).
+The parameter $\delta$ captures the difficulty of profit shifting.
+[^11] This can be higher, for example, due to taxpayer audit (Lediga et al., , 2023).
+Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^12] The GMT tax base equals that of national corporate taxes, ignoring additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler, (2024).
 $\max_{\{\theta_{h}\}_{h=1}^{H}}\ \widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)=(1-t_{n})\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi+\sum_{h=1}^{H}(1-t_{h})\theta_{h}\pi-\sum_{h=1}^{H}\frac{\delta(\theta_{h}\pi)^{2}}{2\pi}.$ (1)
-The MNEs’ optimal level profit shifting to each tax haven, $\theta_{h}$ is then:
+The optimal choice is
 $\theta_{h}=\frac{t_{n}-t_{h}}{\delta},$ (2)
-which is independent of $\pi$ and therefore holds for all MNEs simultaneously.
-[^15] In principle, our analysis can incorporate $\theta<0$, which occurs if $t_{n}<t_{h}$.
-However, this case will never arise in the tax-competition equilibrium we will see shortly.
-The optimal $\theta$ defined by (2) allows us to incorporate the MNEs’ adjustment to tax differentials in a very simple and compact form.
-It implies that MNEs of different size will respond to international tax differentials with the same tax base elasticity.
-Substituting (2) into (1) and using the symmetry of tax havens, the tax base of a MNE with profits $\pi$ in the non-haven country is $TB_{n}=(1-\sum_{h=1}^{H}\theta_{h})\pi=[1-H(t_{n}-t_{h})/\delta]\pi$.
-This yields a tax base elasticity of an individual MNE in the non-haven country such that $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, which is independent of $\pi$.
+which is independent of $\pi$ and therefore holds for all MNEs.
+The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity.
+Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, independent of $\pi$.
+Equal tax base elasticities remove countries’ incentives for tax discrimination across MNEs in the absence of the GMT.
+This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
 The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive.
@@
 [^13] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms.
-Applying a similar empirical strategy to US firms, Coles et al., (2022) also find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size.
-See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different size.
+Applying a similar empirical strategy to US firms, Coles et al., (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size.
+See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes.
 Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark.
-As a consequence, there is no reason per-se in our model that countries differentiate tax rates between MNEs of different size.
-However, as we will see below, a split tax regime can nevertheless occur in equilibrium as a result of introducing a GMT for large firms only.
 1.3 Governments’ tax setting choices
-Before examining the effects of GMT, we first solve the unconstrained optimization problem in the absence of a GMT.
-The non-haven country $n$ maximizes a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, where all MNE owners reside exclusively in the non-haven country.
-The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal valuation of public funds.
-The welfare premium on tax revenues, relative to private income, is a standard assumption in international taxation, which motivates positive tax rates in equilibrium.
-It can be given two alternative interpretations.
-The first is a redistributive (or political-economy) motive, when the government uses the corporate tax to redistribute income from capital owners to other households in the economy, which possess no earned income.
-The other interpretation is that corporate tax revenue is used to reduce other distortive taxes, and the (exogenous) deadweight loss associated with them.
-[^17] The latter argument assumes that there is a fixed excess burden of taxation in a country’s tax system (see e.g., Keen and Lahiri,, 1998).
-This is a plausible assumption in our setting, as we focus on the corporation tax, which accounts for less than 10% of total tax revenue in all developed countries.
-Either setting leads to a welfare gain for the non-haven country from levying a positive corporation tax on its MNEs.
-Finally, for $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues.
-This special case underlies, for example, the analyses in Hindriks and Nishimura, (2022) or Janeba and Schjelderup, (2023).
-The government of country $n$ therefore maximizes
-[…]
+Before examining the effects of the GMT, we first solve the unconstrained optimisation problem in the absence of the GMT.
+The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven.
+The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, , 2023).
+For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues (e.g., Janeba and Schjelderup, , 2023; Hindriks and Nishimura, , 2025).
+The government of the non-haven $n$ therefore maximises
+$G_{n}$ $=\underbrace{\int_{\underline{\pi}}^{\infty}\widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)dF}_{\text{Private benefit}}+\lambda\underbrace{\int_{\underline{\pi}}^{\infty}t_{n}\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi dF}_{\text{Tax revenues}}$
 $=(1-t_{n})\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi+\sum_{h=1}^{H}(1-t_{h})\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi-\sum_{h=1}^{H}\frac{\delta}{2}\left(\frac{t_{n}-t_{h}}{\delta}\right)^{2}\Pi$
-$\qquad\qquad+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
-[…]
-Each of the tax haven countries $h$ only cares about its tax revenues (multiplied by the same marginal value of public funds, $\lambda$, as for country $n$), as the affiliate located in $h$ is owned by their headquarters in the non-haven.
-This gives
-[…]
+$\hskip 18.49988pt\hskip 18.49988pt+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
+where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ is the total true profits of MNEs.
+Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
+$G_{h}=\lambda\int_{\underline{\pi}}^{\infty}t_{h}\theta_{h}\pi dF=\lambda t_{h}\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi.$ (4)
 To determine world welfare, we take a utilitarian approach and define $G_{W}\equiv G_{n}+\sum_{h=1}^{H}G_{h}$, using (1.3) and (4).
-Solving the first-order conditions of the two countries yields best responses: [^18] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
+Solving the first-order conditions of the two sets of countries yields best responses: [^14] The second-order conditions of both countries’ optimal tax problems are trivially satisfied.
 $t_{n}=\frac{(\lambda-1)\left(\sum_{h=1}^{H}t_{h}+\delta\right)}{H(2\lambda-1)},\qquad t_{h}=\frac{t_{n}}{2}.$ (5)
-Solving for the optimal tax rates in the unconstrained benchmark (“Regime 0”) then gives
+Solving for the optimal tax rates in the unconstrained benchmark, which we call ‘Regime 0,’ gives
 $t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)},\qquad t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}.$ (6)
-[…]
-A few comments on (6) are in order.
+We hereafter assume $\delta\in(0,\,3H/2)$ to ensure $t_{n}^{0}\in(0,1)$.
+Two comments on (6) are in order.
 First, tax rates in all countries are rising in the profit-shifting costs $\delta$ and falling in the number of tax havens $H$.
-Second, they are rising in the marginal value of public funds, $\lambda$.
-A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the valuation of tax revenues vis-à-vis the post-tax profits of MNEs.
+Second, they are rising in the MVPF, $\lambda$.
+A higher value of $\lambda$ raises the optimal tax rate in the non-haven by increasing the value of tax revenues vis-à-vis the post-tax profits of MNEs.
 The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well.
-Finally, each of the havens without an independent tax base will choose one-half of the tax rate of the non-haven.
-[^19] Note the important difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT.
-This contrast arises from different specifications of profit-shifting costs.
+[^15] Note the difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT.
 Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country.
-This winner-takes-all situation among the havens drives their equilibrium tax rates to zero.
-By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs.
-Hence each MNEs has an incentive to diversify its allocation of shifted profits across all havens.
-As we discuss in our calibration in Section 4, a setting with positive tax rates in haven countries in the unconstrained tax equilibrium is more in line with the empirical evidence.
-We can also derive the share of shifted profits in this unconstrained tax equilibrium. This is given by:
+By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
+The share of shifted profits in Regime 0 is given by
 $\sum_{h=1}^{H}\theta_{h}^{0}=H\cdot\left(\frac{t_{n}^{0}-t_{h}^{0}}{\delta}\right)=\frac{H}{\delta}\cdot\frac{\delta(\lambda-1)}{H(3\lambda-1)}=\frac{\lambda-1}{3\lambda-1}<\lim_{\lambda\to\infty}\sum_{h=1}^{H}\theta_{h}^{0}=\frac{1}{3}.$ (7)
-Hence, when the non-haven country maximizes tax revenues ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium.
+Hence, when the non-haven maximises tax revenues only ($\lambda\to\infty$), the international tax differential between the non-haven and the havens is maximal, resulting in one-third of all profits of MNEs being shifted to the havens in equilibrium.
 For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
+2 Theoretical results
 2.1 Tax regimes
-With a GMT in place, there is a first stage of the game in which each country chooses to commit or not to a single tax rate, which can either be the GMT rate or a tax rate above the GMT.
-If a country does not commit to a single tax rate, then it will choose to split tax rates in the second stage, and set different tax rates for large and small MNEs.
-MNEs with profits equal to or greater than the threshold $\pi_{M}$, i.e., $\pi\in[\pi_{M},\infty)$, are subject to the GMT requirement (large MNEs), whereas those with $\pi\in[\underline{\pi},\pi_{M})$ are not (small MNEs).
-We denote by $\phi$ the share of total profits earned by large MNEs, or the coverage rate of the GMT, which is exogenously given:
-[…]
+With a GMT in place, in the first stage of the game each country decides whether to commit to a single tax rate—equal to or above the GMT rate.
+Non-committing countries split rates in the second stage, setting different rates for large and small MNEs.
+Large MNEs with profits $\pi\in[\pi_{M},\infty)$ are subject to the GMT, while small MNEs with $\pi\in[\underline{\pi},\pi_{M})$ are not.
+Let $\phi$ denote the exogenous share of profits earned by large MNEs, i.e., the coverage rate of the GMT:
+$\phi\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\int_{\underline{\pi}}^{\infty}\pi dF}\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\Pi}.$ (8)
+Table 1. Possible Regimes of the Non-Cooperative Tax Game.
 Non-haven Haven
@@
 split tax rate Regime 4
-Table 1: Possible regimes of the non-cooperative tax game
-Table 1 shows that this sequence yields three potential choices for each of the two sets of countries.
+Table 1 shows possible combinations of tax policies of the non-haven and the havens.
 Of the nine potential tax regimes, four regimes cannot occur in equilibrium.
-In these potential regimes, if the non-haven sets a single GMT rate, the havens would choose a single tax that is at the same level or even higher.
-Hence tax revenues in each haven country would be zero.
-Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven country for small MNEs.
-Alternatively, if the haven sets a single non-GMT or a single GMT tax rate, the non-haven will never want to split its tax rate, because the tax base elasticity of its MNEs is independent of profits.
+In these regimes, if the non-haven adopts a single GMT rate, the havens set a uniform tax rate at the same level or higher and therefore earn no positive tax revenues.
+Clearly, this cannot be an equilibrium because each haven can secure strictly positive tax revenues by splitting its tax rate and underbidding the tax rate of the non-haven for small MNEs.
+Alternatively, if the havens set a single non-GMT or a single GMT rate, the non-haven never wants to split its tax rate, due to the constant tax base elasticity.
 Hence only five possible tax regimes remain.
-In Regime 0, the GMT rate is not binding for any country; this is the case of unconstrained tax competition analyzed in the previous section.
-As the GMT rate is continuously increased, the tax competition equilibrium passes through four other regimes.
-In the following, we first state the full characterization of the sub-game perfect Nash equilibrium of the three-stage tax competition game, before commenting on each of these regimes.
+Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier.
+As $t_{M}$ rises, the equilibrium moves through four additional regimes, characterised as follows.
 Proposition 1
 (Equilibrium regimes with binding GMT)
-Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal valuation of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$.
-As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterized by the following four regimes, with all symmetric haven countries choosing the same tax schedule.
+Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$.
+As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
   • (i)
@@
 The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
-    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\pi_{M},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\pi_{M},\infty).$
+    $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
   • (ii)
@@
     $\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
-Proof: See Appendix A.1.
-In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$.
-We call these the regime-switching rates.
-Our discussion below will focus on tax revenue effects.
-This is strictly correct only for the tax havens.
-It is also possible for the non-haven, however, because the marginal value of public funds, $\lambda$, is constant in our analysis, and it always exceeds the valuation of private income.
-All formal proofs are relegated to the Appendix.
-In Regime 1, where $t_{M}$ rises above the haven’s unconstrained tax rate $t_{h}^{0}$, the haven is forced to set $t_{M}$ for large MNEs.
-By contrast, the non-haven’s unconstrained rate is still above the GMT.
-As the non-haven is not bound by the GMT, it will continue to choose a single tax rate above the GMT for all MNEs.
-For the tax havens, the choice is therefore whether to apply the GMT also to small MNEs, or use a split tax schedule.
-By committing to a single GMT rate for all MNEs, the havens can induce the non-haven to set a higher uniform tax rate than in the unconstrained equilibrium.
-This argument is strongest when tax havens collectively decide on whether to commit to a single tax rate, as we assume in our main model.
-[^20] This is seen from the non-haven’s best response function (5), which is increasing in $t_{h}$.
-However, as we show in Appendix A.2, the same qualitative argument also holds for decentralised decision-making of tax havens.
-Havens therefore trade off the gains from inducing a higher uniform tax rate in the non-haven against the revenue losses from not undercutting the non-haven’s tax rate even more for small MNEs.
-In Regime 1, the first effect dominates and the havens find it profitable to commit to the GMT for all MNEs.
-As a result, all countries choose higher uniform tax rates in Regime 1, as compared to the unconstrained equilibrium without a GMT.
-In Regime 2, where $t_{M}$ rises above the regime-switching rate $t_{M}^{1}$, the non-haven is still not bound by the GMT, and will continue to set a single tax rate above the GMT for all MNEs.
+Proof: See Appendix 1.
+In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates.
+In Regime 1, when the GMT rate $t_{M}$ exceeds the havens’ unconstrained rate $t_{h}^{0}$, they must apply $t_{M}$ to large MNEs.
+The non-haven’s unconstrained rate remains above $t_{M}$, so it continues to set a single rate above $t_{M}$ for all MNEs.
+The havens then choose whether to apply $t_{M}$ also to small MNEs or use a split schedule.
+By committing to a single GMT rate, they can induce the non-haven to raise its uniform rate.
+This argument is strongest when the havens decide collectively, as assumed in our main model.
+[^16] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Supplemental Appendix B.
+In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens.
+The havens trade off gains from inducing a higher non-haven rate against revenue losses from not undercutting the non-haven’s rate even more for small MNEs.
+In Regime 1, the first effect dominates, so they commit to the GMT rate for all MNEs, leading all countries to set higher uniform tax rates than in Regime 0.
+In Regime 2, with $t_{M}$ above the regime-switching rate $t_{M}^{1}$, the non-haven remains unconstrained and sets a single rate above $t_{M}$ for all MNEs.
 For the havens, however, the trade-off described in Regime 1 above changes.
-As the GMT is now higher, there is less to gain from setting a uniform GMT rate for all MNEs, because the tax differential and thus shifted profits are reduced.
-[^21] This is seen from Regime 2 of Proposition 1, where the tax differential, $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases in $t_{M}$.
-At the same time, given the higher GMT rate, there is more to gain for the havens from undercutting the non-haven’s tax rate for small MNEs.
-In Regime 2, therefore, the havens will find it optimal to split their tax rate and set $t_{M}$ only for the large MNEs, but choose a lower tax rate for small MNEs.
-[^22] In practice, this split is done by choosing the lower tax rate for small MNEs as the general corporate tax rate, but topping up this tax rate for large MNEs.
-See footnote 5 in the introduction.
-Johannesen, (2022) has shown formally that it is always rational for tax havens to use the QDMTT up to the total tax level of the GMT, rather than leaving the top-up taxation to the home countries of the MNEs.
-Note, finally, that the upper boundary of Regime 2, $t_{M}^{2}$, is rising in the GMT coverage rate, $\phi$.
-This is because the non-haven country faces less competition from the havens (i.e., higher tax rates $t_{h}$) for large firms.
-If large MNEs make up a large part of the corporate tax base, the non-haven country will therefore be able to set a uniform tax rate above $t_{M}$ for a wider range of GMT rates.
-In Regime 3, where $t_{M}$ rises above the regime-switching rate $t_{M}^{2}$, the non-haven will no longer charge a tax rate above the GMT level, but it is bound by the GMT tax rate $t_{M}$ for large firms.
-By committing to set $t_{M}$ for all MNEs, it induces the haven countries to charge a higher tax rate for small MNEs.
-This argument is analogous to the one that applied to the tax havens in Regime 1.
-In equilibrium, therefore, the non-haven sets $t_{M}$ for all firms, rather than splitting its tax rate, and the havens respond by charging $t_{M}/2$ for small firms.
-Finally, in Regime 4, where $t_{M}>t_{M}^{3}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set the GMT rate for all firms.
-[^23] This is seen from Regime 3 of Proposition 1, where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases in $t_{M}$.
-To avoid large tax base losses from small MNEs, the non-haven therefore abandons its commitment and also undercuts the GMT for small MNEs.
+The higher $t_{M}$ reduces gains from a uniform rate, as the tax differential and thus shifted profits shrink.
+[^17] This is seen from Proposition 1(ii), where the tax differential $t_{n}^{2}-t_{h}^{2}=[\delta(\lambda-1)-\lambda Ht_{M}]/[H(2\lambda-1)]$ decreases with $t_{M}$.
+At the same time, it increases losses from not competing with the non-haven for small MNEs.
+Consequently, the havens split their rates, applying $t_{M}$ to large MNEs and a lower rate to small MNEs.
+In Regime 3, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the non-haven can no longer set a rate above $t_{M}$ but is bound by it for large MNEs.
+By committing to $t_{M}$ for all MNEs, it induces the havens to raise their rates for small MNEs, analogous to the argument in Regime 1.
+In equilibrium, the non-haven applies $t_{M}$ to all firms, while the havens set $t_{M}/2$ for small MNEs.
+Finally, in Regime 4, with $t_{M}$ above the regime-switching rate $t_{M}^{2}$, the international tax differential for small MNEs becomes very large, if the non-haven keeps its commitment to set $t_{M}$ for all MNEs.
+[^18] This is seen from Proposition 1(iii), where the tax differential $t_{n}^{3}-t_{h}^{3}=t_{M}/2$ increases with $t_{M}$.
+To limit tax base losses from small MNEs, the non-haven abandons its commitment and lowers its tax rate for them.
 Again, this argument is analogous to the one for the havens to end their commitment in Regime 2.
-In the Regime 4 equilibrium, both countries therefore split their tax rates and set their unconstrained tax rates for small MNEs, while adhering to the GMT for large MNEs.
-Figure 2: Equilibrium tax rates for different GMT rates
-Notes: The figure shows the equilibrium tax rates in the non-haven (a) and the haven (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes (c).
-Parameter values given in Table 2 affect the precise levels of tax rates and regime-switching points, but do not affect qualitatively the sequence and pattern of tax-competition equilibria.
-Figure 2 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of $t_{M}$.
+In Regime 4, both countries split their rates, applying unconstrained rates to small MNEs and the GMT rate to large MNEs.
+Fig.1. Equilibrium Tax Rates for Different GMT Rates.
+Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c).
+While the levels of tax rates and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.
+Figure 1 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$.
 The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b).
-The equilibrium regimes are summarized in panel (c).
-The solid curves give the tax rate(s) in the tax-competition equilibrium under the GMT; the dashed horizontal line is the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0).
-Tax rates in both countries generally rise as the GMT increases, but there are distinct patterns in the different regimes.
-At the switch from Regime 1 to 2, the haven countries start splitting their tax rates and their tax rate on small MNEs drops discontinuously.
+The equilibrium regimes are summarised in panel (c).
+The solid (dotted) curves are the tax rates for large MNEs (small MNEs) in the tax-competition equilibrium under the GMT; the dashed horizontal lines are the unconstrained equilibrium tax rate $t_{i}^{0}$ (in Regime 0).
+Tax rates in both countries generally rise as $t_{M}$ increases, but there are distinct patterns in the different regimes.
+At the switch from Regime 1 to 2, the havens begin to split their tax rates, and the rate for small MNEs falls discontinuously.
 At the switch from Regime 3 to 4, the non-haven splits its tax rate and discretely cut its tax rate for small MNEs to the level in the unconstrained Regime 0.
-The haven countries respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
-3 Welfare effects of the GMT
-Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of the introduction of the GMT and of two gradual reforms: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$.
-As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case.
-Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
-3.1 Introduction of the GMT
-We first compare welfare levels in the tax-competition equilibrium to those in the unconstrained equilibrium, Regime 0. This is summarized in:
-Proposition 2
-(GMT introduction)
-Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1, and assume the GMT coverage rate is not too low so that $\phi\in(3/4,1)$.
-Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
-  • (i)
-    a rise in welfare in the non-haven country for all Regimes 1 to 4.
-  • (ii)
-    a rise in welfare in the haven country for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
-  • (iii)
-    a rise in world welfare for all Regimes 1 to 4.
-Proof: See Appendix A.3.
-Introducing a small GMT slightly higher than the havens’ unconstrained rate $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the haven countries.
-At the same time, profit shifting is reduced in comparison to the no-GMT benchmark due to a smaller tax differential.
-The sum of these effects must certainly benefit the non-haven.
-It also benefits the tax havens, as long as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
-However, once the GMT rate exceeds a certain threshold, $t_{M}^{+}$, which still lies in Regime 1, the reduced tax differential and thus the fall in shifted profits will dominate from the perspective of the haven countries.
-Therefore, for all levels of the GMT above $t_{M}^{+}$, tax havens lose from the introduction of the GMT while the non-haven gains.
-[^24] This conflict of interest does not arise in (Johannesen,, 2022, Proposition 6), where havens gain from the introduction of a GMT at any level, as long as profit shifting is not completely eliminated.
-In his setting where MNE shift profits only to the lowest-tax country, all the tax havens end up setting their tax rate to zero.
-Therefore, revenues in the unconstrained tax competition equilibrium are zero for each haven, and any positive GMT rate generates a revenue increase for the representative haven.
-See also footnote (19).
-In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0.
-As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs.
-In this case, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.252525If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate will instead benefit the havens in Regime 3.
-See Appendix A.3 for details.
-Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes.
-These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
-3.2 Gradual reforms of the GMT
-We now turn to two likely reform options after the GMT has been introduced. We first study the effects of a gradually increasing GMT rate. The welfare effects of this reform are summarized in:
-Proposition 3
-(Gradual reform: increasing GMT rate $t_{M}$)
-Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1. A marginal increase in the GMT rate has the following effects:
-  • (i)
-    welfare in the non-haven country increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it discretely falls.
-  • (ii)
-    welfare in the haven country increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases for the remainder of Regime 1 and in Regime 2.
-It increases in Regime 3 and discretely falls at the last regime-switching point, $t_{M}^{3}$.
-It does not change in Regime 4.
-  • (iii)
-    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it discretely falls.
-Proof: See Appendices A.3 and A.4.
-As is true for the introduction of the GMT (Proposition 2), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium.
-An exception is the regime-switching point $t_{M}^{1}$ where the havens start splitting their tax rate and set a discretely lower tax rate for small MNEs.
-Small MNEs then shift more profits to the havens, hurting the non-haven.
-For the havens, the welfare (tax revenue) effects of an increase in the GMT depend on the negative effect of reduced profit-shifting opportunities and the positive effect of a higher tax rate in the non-haven.
-This leads to non-monotonic changes in the havens’ welfare.
-In Regime 1, the induced higher tax rate in the non-haven dominates initially, but as the GMT continues to rise, this is overcompensated by the reduced levels of profit shifting.
-The latter is also true throughout Regime 2.
-In Regime 3, the effect of a higher tax rate dominates again, as the non-haven commits to the uniform GMT rate in order to raise the havens’ tax rates.
-This regime, where commitment induces effective tax coordination, is similar to (the first part of) Regime 1.
-It ends at the regime-switching point $t_{M}^{3}$ where the non-haven starts splitting its tax rate.
-This leads to a discrete fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens.
-The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive.
-At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, where one of the countries starts splitting its tax rate and discretely lowers the tax rate on small MNEs, tax competition is tightened and world welfare falls at the margin.
-Next we examine the welfare effects of another likely reform option: a marginal increase in the GMT coverage rate $\phi$, while keeping the GMT rate $t_{M}$ fixed.
-As shown in (8), this is achieved by reducing the threshold profit level $\pi_{M}$ above which the GMT rate must be charged.
-The results are summarized in:
-Proposition 4
-(Gradual reform: increasing GMT coverage rate $\phi$)
-Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1.
-Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
-  • (i)
-    welfare in the non-haven is unaffected in Regime 1 and increases in Regimes 2 to 4.
-  • (ii)
-    welfare in the haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
-  • (iii)
-    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
-Proof: See Appendix A.5.
-In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates.
-Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates.
-Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies fewer small MNEs, reduces aggregate profit shifting.
-This unambiguously hurts the haven and benefits the non-haven.
-From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
-There are two exceptional cases where increasing $\phi$ causes a regime switch.
-The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rate.
-This regime switch benefits the non-haven and the world, while it has no effect on the havens.
-The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT.
-This latter regime switch, however, leaves welfare in all countries and the world unaffected.
-In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3).
-A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1.
-One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
+The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
+2.2 Welfare effects of the GMT
+We briefly summarise how the GMT introduction affects national and global welfare.
+Introducing a GMT rate $t_{M}$ slightly above the havens’ unconstrained one $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the havens.
+At the same time, profit shifting is reduced in comparison to the unconstrained Regime 0 due to a smaller tax differential.
+The sum of these effects must benefit the non-haven.
+It also benefits the havens, as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
+Once $t_{M}$ exceeds a certain threshold within Regime 1, the relationship reverses for the havens, as the loss of shifted profits dominates.
+Introducing any GMT rate above this threshold benefits the non-haven but harms the havens, with the net effect raising global welfare.
+Supplemental Appendix C provides a formal proposition and proof of these results.
+We also study the effects of marginal increases in the GMT tax rate $t_{M}$ and the GMT coverage rate $\phi$.
+An important result is that welfare in the non-haven and the world increases with $t_{M}$ in general, and the welfare gains become even larger as $\phi$ increases.
+A higher $\phi$ reduces the effectiveness of the split tax system and limits tax competition for small MNEs, thereby amplifying the welfare gains from raising $t_{M}$.
+See Supplemental Appendices D and E for formal discussions and Appendix F for graphical illustrations.
 3 Quantitative implications
-We now examine which regime can be expected under the currently imposed GMT rate.
-For this purpose, we use a calibrated version of our model and explore its quantitative effects.
-We calibrate our unconstrained model without GMT (Regime 0) to match basic data on international profit shifting.
-The results are summarized in Table 2.
-The coverage rate of the GMT is set at $\phi=0.9$, in line with OECD, 2020a and our own calculations from the ORBIS database (see Figure 1 and Appendix B).
-The marginal valuation of public funds (MVPF) is set at $\lambda=1.5$ in our calibration benchmark, corresponding to the benchmark value used in Hebous and Keen, (2023).
-The figures for total pre-tax profits of MNEs, $\Pi$, and the number of tax havens, $H$, are taken from Tørsløv et al., (2023) and the Country-by-Country Report Statistics (CbCR); see the notes to Table 2 for details.
-We then calculate the cost parameter of profit shifting, $\delta$, to exactly match the GDP-weighted average of effective tax rates in non-haven countries of $18.6\%$, as reported in Tørsløv et al., (2023).
-Among the non-targeted moments, our model somewhat underestimates the havens’ tax rate in the data, as well as the total profits shifted by MNEs and –accordingly– the non-havens’ loss of tax revenues.
-Overall, however, our simple model explains the data reasonably well.
+We now explore the quantitative importance of our theoretical results.
+We modify the model of the previous section by assuming that the havens decide tax policies in a decentralised manner (see footnote 16 and Supplemental Appendix B).
+The decentralised decision making leads to competition between the havens and thus makes it more likely that they split their tax rates.
+[^19] The new regime switching rate dividing Regimes 1 and 2 is smaller ($t_{M}^{b}<t_{M}^{1}$), and thus Regime 1 is less likely to emerge, as $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$.
+This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens, making gains from commitment smaller.
+We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2.
+The GMT coverage rate is set to $\phi=0.9$, following OECD, (2020) and our calculations from Orbis.
+[^20] See Supplemental Appendix A for details.
+The number of tax havens is set to $H=40$, as in Tørsløv et al., (2023).
+Combining their data with Country-by-Country Reporting (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD.
+The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the model-predicted non-haven tax rate and shifted profit share and their empirical counterparts.
+For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
+Table 2. Calibration of the Model.
 Parameter Source
-$\Pi=4,623$bUSD Total pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
-$\phi=0.9$ Profit share of MNEs covered by GMT OECD, 2020a , Orbis
+$\phi=0.9$ Profit share of MNEs covered by GMT OECD, (2020), Orbis
 $H=40$ Number of tax havens Tørsløv et al., (2023)
-$\lambda=1.5$ Marginal valuation of public funds Hebous and Keen, (2023)
-$\delta=27.4$ Cost of profit shifting Calibrated
+$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
+$\lambda=2.12$ Marginal value of public funds Calibrated
+$\delta=17.80$ Cost of profit shifting Calibrated
 Targeted moment Model Data Source
-Corp. tax rate of non-haven (Regime 0): $t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ $18.6\%$ $18.6\%$ Tørsløv et al., (2023)
+Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al., (2023)
+Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al., (2023), CbCR
 Non-targeted moment Model Data Source
-Corp. tax rate of haven (Regime 0): $t_{h}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
-MNE’s shifted profits: $\sum_{h}\theta_{h}^{0}\Pi$ $660$bUSD $969$bUSD Tørsløv et al., (2023)
-Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ $129$bUSD $247$bUSD Tørsløv et al., (2023)
-Table 2: Calibration of the model
+Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
+Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al., (2023)
 Notes: Data are from 2019.
-The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ”Table1” and ”Table U1” in the excel file ”1975-2019 updated estimates: Tables”).
+The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’).
 The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world.
@@
 With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
-Figure 3: Equilibrium welfare and shifted profits
-Notes: The figure shows the equilibrium levels of welfare normalized by the MVPF, $\lambda$, in the non-haven (a) and the sum of the $H=40$ haven countries (b), and the aggregate shifted profits (c), for different rates of GMT.
-Parameter values are given in Table 2.
-The solid curves in each panel are the equilibrium values under the partial GMT ($\phi=0.9$); the dashed horizontal lines are those in the unconstrained equilibrium (Regime 0).
-For this calibrated version of our model, the effects of introducing a GMT at varying tax rates $t_{M}$ on equilibrium tax schedules are shown in Figures 2, which we already discussed in Section 2.4.
-Figure 3 illustrates the effects on welfare in both non-haven and haven countries, and those on aggregate shifted profits, as compared to their counterparts in the unconstrained equilibrium (Regime 0).
-To make numbers comparable with tax revenues, the levels of welfare are normalized by the MVPF, $\lambda$.
-The most important result is that the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 2.
-In this regime, the havens find it profitable to split their tax rate and to set a lower tax rate for small MNEs as compared to large MNEs (Proposition 1(ii)).
-[^26] This result also holds if we assume that the first stage decision of whether to split tax rates or not is made by each tax haven individually (see Appendix A.2).
-In this case the range of GMT rates that support Regime 1 will be smaller than in our benchmark model, while the range of $t_{M}$ that supports an equilibrium in Regime 2 rises accordingly.
-This corresponds to the observations mentioned in the Introduction that several tax havens have announced to maintain their low statutory tax rate while topping up the taxation of large MNEs to comply with the GMT, known as the QDMTT (see footnote 5).
-The GMT introduction reduces the total amount of profits shifted, thereby increasing the non-haven’s welfare while reducing the havens’ welfare (tax revenues).
-Panel (a) of Table 3 further quantifies the welfare and revenue effects of a 15% GMT rate for our calibrated model, with the benchmark MVPF of $\lambda=1.5$.
-Welfare and tax revenue increases in the non-haven country clearly outweigh the revenue losses in the havens, leading to an increase in global welfare and tax revenues.
-Our calibrated gain in tax revenue in the non-haven (107 billion USD) is close to the summed revenue gains for Europe and the U.S.
-in Baraké et al., (2022) and worldwide tax revenue gains are in the same range as in Hugger et al., (2024) (see footnote 2).
-[^27] The percentage increases in corporate tax revenue in our model are not comparable to those in Baraké et al., (2022) and Hugger et al., (2024), as we do not model purely domestic firms.
-The welfare gain for the non-haven is smaller than its revenue gain, because the private income in the non-haven (the post-tax profits of MNEs net of profit-shifting costs) falls after the introduction of GMT.
-(a) MVPF: $\lambda=1.5$.  Regime 0 to 2 Non-haven Haven World
-Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
-Tax rate after GMT introduction $19.9\%$ for all $\begin{cases}9.9\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
-Welfare change in bUSD (change) $53$ ($1.6\%$) $-7$ ($-12.1\%$) $46$ ($1.4\%$)
-Revenue change in bUSD (change) $107$ ($14.5\%$) $-7$ ($-12.1\%$) $99$ ($12.4\%$)
-Shifted-profits change in bUSD (change) $267$ $(41.9\%)$ $-267$ $(-41.9\%)$ 00 $(0\%)$
-(b) MVPF: $\lambda=3$.  Regime 0 to 1 Non-haven Haven World
-Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
-Tax rate after GMT introduction $20.9\%$ for all $15.0\%$ for all —–
-Welfare change in bUSD (change) $122$ ($6.4\%$) $2$ ($2.0\%$) $124$ ($6.1\%$)
-Revenue change in bUSD (change) $168$ ($26.0\%$) $2$ ($2.0\%$) $170$ ($22.6\%$)
-Shifting-profits change in bUSD (change) $425$ $(36.8\%)$ $-425$ $(-36.8\%)$ 00 $(0\%)$
-(c) MVPF: $\lambda=10$.  Regime 0 to 1 Non-haven Haven World
-Tax rate before GMT introduction $18.6\%$ for all $9.3\%$ for all —–
-Tax rate after GMT introduction $21.3\%$ for all $15.0\%$ for all —–
-Welfare change in bUSD (change) $169$ ($17.3\%$) $12.4$ ($9.3\%$) $181$ ($16.3\%$)
-Revenue change in bUSD (change) $185$ ($31.1\%$) $12.4$ ($9.3\%$) $197$ ($27.1\%$)
-Shifting-profits change in bUSD (change) $463$ $(32.2\%)$ $-463$ $(-32.2\%)$ 00 $(0\%)$
-Table 3: Quantitative effects of introducing a $15\%$ GMT
-Notes: Each of panels (a) to (c) show the effects of a $15\%$ GMT introduction for different values of the MVPF, $\lambda\in\{1.5,3,10\}$.
-Results are based on the calibrated parameter values described in Table 2; the GMT coverage rate is $90\%$ ($\phi=0.9$) in particular.
-The welfare levels are normalized by the MVPF, $\lambda$.
-With $\lambda=1.5$, the $15\%$ GMT introduction leads the economy from Regime 0 to 2; with $\lambda\in\{3,10\}$, it leads instead to Regime 1.
-The haven refers to the sum of the $H=40$ haven-countries.
-Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
-One may argue that the MVPF is indeed higher than our benchmark value of $\lambda=1.5$, as evidenced by the very substantial efforts, discussed in the introduction, that non-haven countries take to increase their corporate tax revenues.
-Panels (b) and (c) of Table 3 therefore present the results of two robustness checks with high values of the MVPF: $\lambda=3$ and $\lambda=10$.
-The latter value, given in Panel (c) approaches a Leviathan government objective of tax revenue maximization.
-With these increased values of $\lambda$, the introduction of a $15\%$ GMT rate leads to a tax equilibrium in Regime 1, in which both countries employ a uniform tax rate on all MNEs.
-With the higher valuation of tax revenues in the non-haven country, the tax havens will find it profitable to commit to a uniform GMT rate for all MNEs, which in turn leads the non-haven to choose a higher uniform tax rate above the GMT, as compared to the benchmark in Panel (a).
-The GMT rate of $15\%$ therefore acts as a coordination device between the non-haven and the havens, benefitting both countries.
-[^28] These results depend on the assumption that tax havens make their first stage choice collectively.
-With decentralized first stage decisions of tax havens (cf.
-Appendix A.2), the upper boundary for a Regime 1 equilibrium, $t_{M}=t_{M}^{a}$, is calculated to be $9.37\%$, $9.42\%$, and $9.45\%$ for respective values of the marginal valuation of public funds, $\lambda=1.5$, $3$, and $10$.
-This implies that the introduction of a $15\%$ GMT rate leads to a Regime 2 equilibrium regardless of the level of $\lambda$.
-For $\lambda=1.5$, the quantitative effects under decentralized decision making of havens are the same as those in Panel (a) of Table 3.
-For $\lambda=3$ and $\lambda=10$, the effects differ from those in Panels (b) and (c) of Table 3, however, and yield a conflict of interests between the non-haven and the havens that is analogous to the one in Panel (a).
-Although observations at the time of writing somewhat favor the GMT resulting in split tax rates in the havens, corresponding to Regime 2, we leave open the question of which regime will eventually emerge.
-(a) Gains from $t_{M}=0.15$ to $t_{M}=0.18$ Non-haven Haven World
-Welfare change in bUSD (change) $34$ ($1.0\%$) $-17$ ($-31.0\%$) $18$ ($0.5\%$)
-Revenue change in bUSD (change) $59$ ($7.0\%$) $-17$ ($-31.0\%$) $42$ ($4.7\%$)
-Shifting-profits change in bUSD (change) $146$ $(37.9\%)$ $-146$ $(-37.9\%)$ 00 $(0\%)$
-(b) Gains from $\phi=0.9$ to $\phi=1.0$ Non-haven Haven World
-Welfare change in bUSD (change) $5.4$ ($0.2\%$) $-0.5$ ($-0.9\%$) $4.9$ ($0.1\%$)
-Revenue change in bUSD (change) $10.7$ ($1.3\%$) $-0.5$ ($-0.9\%$) $10.3$ ($1.1\%$)
-Shifting-profits change in bUSD (change) $27$ $(7.0\%)$ $-27$ $(-7.0\%)$ 00 $(0\%)$
-Table 4: Quantitative effects of gradual reforms
-Notes: Panel (a) shows gains/losses from an increase in the GMT rate from $t_{M}=0.15$ to $0.2$; panel (b) shows gains/losses from an increase in the GMT coverage rate from $\phi=0.9$ to $1.0$.
-Results are based on the calibrated parameter values described in Table 2; the MVPF is $\lambda=1.5$ in particular.
-The welfare levels are normalized by the MVPF, $\lambda$.
-The haven refers to the sum of the $H=40$ haven-countries.
-Due to rounding numbers, the sum of the welfare/revenue change of the non-haven and the haven in billion USD is not necessarily equal to that of the world.
-Finally, Table 4 provides a quantitative assessment of two gradual reforms that are likely to follow the newly introduced GMT rate of $15\%$ with incomplete coverage.
-The first reform is to increase the GMT rate from $15\%$ to $18\%$ while maintaining the coverage rate at $\phi=90\%$ (Panel (a)).
-The second is to broaden the GMT coverage from $90\%$ to $100\%$ while keeping the GMT rate at $15\%$ (Panel (b)).
-In both scenarios, we assume our benchmark MVPF of $\lambda=1.5$ and thus, the equilibrium remains in Regime 2 after the reforms.
-The results indicate that the non-haven gains more from the $18\%$ GMT rate ($+1.0\%$ in welfare; $+7.0\%$ in revenues) than from the full GMT coverage ($+0.2\%$ in welfare; $+1.3\%$ in revenues).
-The greater gains for the non-haven from the $18\%$ GMT rate are driven by the stronger reduction in profit shifting to the havens ($-37.9\%$), as compared to a full GMT coverage ($-7.0\%$).
-Hence the increase in the GMT rate creates a much sharper conflict of interest with the tax havens, as compared to the extension of GMT coverage.
-In this sense, covering more MNEs under the GMT may be a politically more feasible option than increasing the GMT rate.
-However, even a full coverage of the GMT will likely bring only moderate gains to the non-haven and to the world, while entailing higher administrative and compliance costs.
+Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3.
+The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields the equilibrium in Regime 2 (Table 3(a)): the havens split their tax rates and set a lower tax rate ($10.1\%$) for small MNEs as compared to large MNEs ($15.0\%$) (Proposition 1(ii)).
+[^21] Above the regime switching rate of $t_{M}^{b}=0.094$, Regime 2 prevails.
+This result aligns with observations that low-tax countries like Ireland plan to retain their base rate (12.5% in Ireland’s case) while applying top-up taxes on large MNEs under the GMT (QDMTT).
+Table 3. Quantitative Effects of GMT Reforms.
+(a) Gains from a 15% GMT under $\phi=0.9$ Non-haven Haven World
+Tax rate after GMT introduction $20.3\%$ for all $\begin{cases}10.1\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
+Welfare change in bUSD (% change) $87.2$ ($3.5\%$) 3.8-3.8 ($4.3\%$) $83.4$ ($3.2\%$)
+Revenue change in bUSD (% change) $139.5$ ($20.6\%$) 3.8-3.8 ($4.3\%$) $135.7$ ($17.7\%$)
+Profit change in bUSD (% change) $359.9$ $(37.5\%)$ 359.9-359.9 $(-37.5\%)$ 0 $(0\%)$
+(b) Gains from a 15% GMT under $\phi=1.0$ Non-haven Haven World
+Tax rate after GMT introduction $20.6\%$ for all $15\%$ for all —–
+Welfare change in bUSD (% change) $95.6$ ($3.8\%$) 3.6-3.6 ($4.1\%$) $92.0$ ($3.6\%$)
+Revenue change in bUSD (% change) $153.0$ ($22.6\%$) 3.6-3.6 ($4.1\%$) $149.3$ ($19.5\%$)
+Profit change in bUSD (% change) $393.0$ $(41.0\%)$ 393.0-393.0 $(-41.0\%)$ 0 $(0\%)$
+Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$) and panel (b), from $t_{M}=0$ to $0.15$ with full coverage $\phi=1.0$.
+The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven.
+Results are based on the calibrated parameter values described in Table 2.
+The welfare levels are normalised by the MVPF, $\lambda$.
+The ‘haven’ here refers to the sum of the $H=40$ haven countries.
+The non-haven’s revenue rises due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition.
+The calibrated revenue gain in the non-haven (139.5 bUSD) is close to the combined gains for Europe and the US in Baraké et al., (2022, Table 1) ($67+58=125$ bEUR), and the global revenue gain (135.7 bUSD) is comparable to Hugger et al., (2024, Section 8.2) (155–192 bUSD).
+Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues.
+[^22] In the havens without MNE headquarters, revenues coincide with welfare.
+Do the quantitative effects of introducing GMT change if the coverage is extended?
+The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) yields the non-splitting equilibrium in Regime 1 (Table 3(b)).
+By eliminating competition for small MNEs, the extended coverage increases revenue and welfare gains worldwide by 13.6 bUSD and 8.6 bUSD, respectively.
+Based on the estimates by Gaul et al., (2022), these costs would amount to about 10.5 bUSD for our sample, implying a small net revenue gain (+3.1+3.1 bUSD) but a net welfare loss (1.9-1.9 bUSD) globally.
+[^23] Using a survey of German MNEs (see Becker et al., , 2012 for details on their characteristics), Gaul et al., (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR.
+Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in Supplemental Appendix A), the total compliance costs of newly covered MNEs are 1.6 mEUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 bUSD.
+Clifford et al., (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
 4 Conclusion
-This paper has analyzed the effects of a global minimum tax (GMT) that is confined to large multinational enterprises (MNEs), thus leaving at least 10% of the global multinational tax base outside its scope.
-Using a simple model with profit shifting by heterogeneous MNEs, we have shown that this partial coverage of the GMT gives rise to a sequence of tax competition equilibria between a non-haven and a set of symmetric tax haven countries.
-In particular, introducing a low GMT rate that still binds the tax havens induces them to commit to a single GMT rate for all MNEs and thus results in higher welfare and tax revenues in both sets of countries.
-However, a further increase in the GMT rate leads first the tax havens, and then the non-haven, to split their tax rate for large and small MNEs, creating a conflict of interest between the two groups of countries.
-A similar conflict also arises from broadening the coverage of MNEs subject to the GMT.
-The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the haven countries split their tax rate and undercut the GMT for small MNEs (Regime 2).
-This finding corresponds to the observation that tax havens such as Ireland, Liechtenstein, Bermuda, or Bulgaria maintain their regular corporate tax rates below the level of 15%, but top up the taxation of large multinationals to the level of the GMT.
-In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the tax havens lose, resulting in a positive net gain globally.
-Our main objective in this paper has been to understand the implications of an incomplete coverage of the GMT.
-We have done so in a highly stylized model, in order to develop some sharp intuitions.
-With the insights gained, it would be fruitful to extend our basic model to incorporate further important aspects of the global minimum tax and the responses of multinational firms.
-A first extension would allow MNEs to adjust their real pre-tax profits in response to changes in tax rates, for example, by modifying their investment decision.
-A second extension would incorporate a splitting response of multinationals seeking to benefit from potentially lower tax rates on small MNEs.
-A third extension would consider the substance-based income exclusion, i.e., additional tax deductions for MNEs that increase investment and employment.
-We leave these extensions for further research.
-Appendix A: Proof of Propositions
-A.1. Proof of Proposition 1
+This paper has analysed the effects of a global minimum tax (GMT) confined to large multinational enterprises (MNEs), leaving about 10% of the global tax base outside its scope.
+This partial-coverage GMT generates a sequence of tax competition equilibria between a non-haven and tax havens.
+In particular, introducing a low GMT rate that binds the havens induces them to commit to the single GMT rate.
+However, a further increase in the GMT rate leads first the havens, and then the non-haven, to split their tax rates for large and small MNEs.
+The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the havens split their tax rates and undercut the GMT for small MNEs.
+This finding corresponds to the observation that tax havens such as Ireland and Switzerland maintain their base corporate tax rates below 15%, but top up taxes on large MNEs to the level of the GMT.
+In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the havens lose, resulting in a positive net gain globally.
+Appendix
+Appendix 1: Proof of Proposition 1
 Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$.
 A new equilibrium regime emerges when $t_{M}\geq t_{M}^{0}$.
-Since haven countries are all symmetric, we are concerned with the representative haven country $h$ in what follows.
+Since haven countries are all symmetric, we are concerned with a representative haven country $h$ in what follows.
 Regime 1: $t_{M}\in[t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
@@
 $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)},\qquad t_{h}^{1}=t_{M}.$ (A.1)
-$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\qquad G_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
-When the non-haven sets a single non-GMT rate and the haven splits its tax rates, the equilibrium tax rates are
+$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\hskip 18.49988ptG_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
+When the non-haven sets a single non-GMT rate and the haven splits its tax rate, the equilibrium tax rates are
 $t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]},\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$ (A.2)
-$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\qquad G_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
+$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\hskip 18.49988ptG_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
 Given the non-haven’s single non-GMT rate, the haven prefers to set a single GMT rate, if
@@
 Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate, and the haven sets a single GMT rate, given by (A.1).
-Table A1 shows the relevant payoffs in Regimes 1 and 2, where the GMT rate is not binding for the non-haven.
-Non-haven Haven
-single GMT rate split tax rate
-single non-GMT rate $G_{n}^{1},\quad G_{h}^{1}$ $G_{n}^{2},\quad G_{h}^{2}$
-single GMT rate
-split tax rate
-Table A1: Payoff matrix for Regimes 1 ($t_{M}\in[t_{M}^{1},t_{M}^{2}]$) and 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$)
-Notes: In Regimes 1 and 2, the GMT is binding only for the haven.
-The potential regimes where the non-haven does choose a single GMT rate or split tax rates are irrelevant because, whenever possible, it always chooses a single non-GMT rate that solves the unconstrained maximization problem.
 Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
-Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rates in response to a non-haven’s single non-GMT rate.
-The non-haven is still not bound by the GMT and continues to chooses its optimal unconstrained tax rate.
-Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rates, with tax rates given by (A.2).
+Since (A.3) becomes negative for $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, the haven splits its tax rate in response to a non-haven’s single non-GMT rate.
+The non-haven is still not bound by the GMT and continues to choose its unconstrained optimal tax rate.
+Therefore, the tax-competition equilibrium is that the non-haven sets a single non-GMT rate and the haven splits its tax rate (see (A.2)).
 Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
@@
 In response, the haven must always split its tax rate; otherwise it cannot obtain positive tax revenues.
-When the non-haven sets a single GMT rate and the haven splits its tax rates, the equilibrium tax rates and payoffs are
-$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
-$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\qquad G_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
+When the non-haven sets a single GMT rate and the haven splits its tax rate, the equilibrium tax rates and payoffs are
+$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\hskip 18.49988ptt_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
+$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\hskip 18.49988ptG_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
 When both the non-haven and the haven split their tax rates, the equilibrium tax rates are
-$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
-$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\qquad G_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
+$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\hskip 18.49988ptt_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
+$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\hskip 18.49988ptG_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
 Given the haven’s splitting rates, the non-haven prefers to set a single GMT rate, if
@@
 The strict equality holds at $t_{M}=t_{M}^{3}$ and it holds that $t_{M}>t_{M}^{2}>\frac{2\delta(\lambda-1)}{H(3\lambda-1)}$ for any $\lambda>1/3$.
-Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rates, as given in (A.4).
+Therefore, the tax-competition equilibrium is that the non-haven sets a single GMT rate, and the haven splits its tax rate, as given in (A.4).
 Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
-Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rates in response to the havens’ split rates.
-Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates, as given in (A.5).
-Non-haven Haven
-single GMT rate split tax rate
-single non-GMT rate
-single GMT rate $G_{n}^{3},\quad G_{h}^{3}$
-split tax rate $G_{n}^{4},\quad G_{h}^{4}$
-Table A2: Payoff matrix for Regimes 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$) and 4 ($t_{M}\in(t_{M}^{3},1]$)
-Notes: In Regimes 3 and 4, the GMT is binding for both the non-haven and the haven.
-Therefore, the regimes where the non-haven chooses a single-non GMT rate are irrelevant.
-The regimes where the haven sets a single GMT are also irrelevant because, by doing so, the haven would obtain non-positive welfare/tax revenues.
-A.2. Decentralized decision of tax havens
-In our main model, we have assumed that tax havens collectively decide on whether to set a single GMT rate, or to split their tax rate.
-We here examine an alternative, decentralized decision process in the first stage of our game where each single tax haven decides whether to split or not, given the other havens’ decisions.
-To simplify this analysis, we assume that countries can split their tax rate only if they are bound by the GMT.
-This reflects the fact that preferential tax regimes for specific tax bases are not allowed under the OECD’s action plan to fight base erosion.
-Suppose then that $t_{M}\in[t_{M}^{0},t_{M}^{2}]$.
-Out of the total of $H$ tax havens, there are $H_{M}$ havens that commit to a single GMT rate while the remaining $H_{S}=H-H_{M}$ havens split their tax rate.
-When the non-haven sets a single non-GMT rate and a single haven $h$ sets a uniform GMT rate, the equilibrium tax rates are
-$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{1}=t_{M},$ (A.7)
-which is the counterpart of (A.1). Let $G_{h}^{1}(H_{M})$ be the haven’s equilibrium payoff with commitment.
-If the single haven $h$ splits its tax rate instead, whereas the non-haven sets a uniform non-GMT rate, the equilibrium tax rates are
-$t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\qquad t_{h}^{2}=\begin{cases}t_{n}(H_{M})/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},$ (A.8)
-which is the counterpart of (A.2). Let $G_{h}^{2}(H_{M})$ be the haven’s equilibrium payoff under splitting.
-We then ask if a single haven has an incentive to choose a uniform GMT rate given that all the other $H-1$ havens split their tax rates.
-A single haven $h$ has a unilateral incentive to do so, if the following holds:
-$\Delta G_{h}^{12}(H_{M}=1)$ $\equiv G_{h}^{1}(H_{M}=1)-G_{h}^{2}(H_{M}=0)$
-$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta}{\delta H^{2}[\lambda(3+\phi)-(1+\phi)]^{2}[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]}\geq 0,$ (A.9)
-where $\Theta\equiv A-Bt_{M},\quad A\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
-$B\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
-$\text{and}\qquad Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0\ \ \text{or}\ \ t_{M}\geq\frac{\delta(\lambda-1)}{H(3\lambda-1)}\equiv t_{M}^{0}.$
-It is straightforward to show that $\Theta>0$ holds at $t_{M}^{0}$ and hence, by continuity, also for $t_{M}$ marginally above $t_{M}^{0}$.
-Therefore, there must be a positive range of $t_{M}$, with $t_{M}^{0}<t_{M}\leq t_{M}^{a}\equiv A/B$ for which the inequality (A.9) holds for a haven $h$.
-By symmetry, this argument applies to all havens.
-For a Regime 1 equilibrium with commitment to occur, a further condition is that no single haven $h$ must have an incentive to split, given that all the other $H_{M}-1$ havens commit to a single GMT rate.
-A single haven $h$ has no unilateral incentive to split its tax rate, if the following holds:
-$\Delta G_{h}^{12}(H_{M}=H)$ $\equiv G_{h}^{1}(H_{M}=H)-G_{h}^{2}(H_{M}=H-1)$
-$=\frac{(1-\phi)[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta^{\prime}}{\delta H(2\lambda-1)[\phi(\lambda-1)+\lambda(4H-1)-(2H-1)]^{2}}\geq 0,$ (A.10)
-where $\Theta^{\prime}\equiv C-Dt_{M},\quad C\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
-$D\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
-and $t_{M}\geq t_{M}^{0}$ holds again.
-The equivalent condition is $t_{M}\leq t_{M}^{b}\equiv C/D$.
-We can immediately see $\Delta G_{h}^{12}(H_{M}=H)<0$ for $t_{M}\in(t_{M}^{b},t_{M}^{2}]$ and check $t_{M}^{a}<t_{M}^{b}<t_{M}^{1}<t_{M}^{2}$:
-$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{\Theta^{\prime\prime}[(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}]}>0,$
-$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}}{\Theta^{\prime\prime}[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0,$
-$\Theta^{\prime\prime}\equiv 3(3H-1)\lambda^{2}-2(3H-2)\lambda+H-1+\phi(\lambda-1)[\lambda(3H-5)-(H-3)]>0.$
-We can conclude that for $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, all the havens set a single GMT rate and the non-haven sets a single non-GMT rate; for $t_{M}\in(t_{M}^{a},t_{M}^{2}]$, all the havens split their tax rate and the non-haven sets a single non-GMT rate.
-The former regime corresponds to Regime 1 and the latter one to Regime 2 in the text.
-The only difference is that the range of Regime 1 is smaller when havens make their first stage decision in a decentralized way, implying that the havens’ commitment is less likely.
-This is intuitive because the commitment of a single haven does not induce the non-haven to increase its tax rate by as much as the simultaneous commitment of many havens.
-Finally, in the remaining range of the GMT, $t_{M}\in(t_{M}^{2},1]$, the difference between collective vs.
-decentralized decisions of havens plays no role any more.
-In either setting, the tax havens always have an incentive to split their tax rate, whether the non-haven sets a single GMT rate (Regime 3), or splits its tax rate as well (Regime 4).
-A.3. Proofs of Propositions 3 and 4
+Since (A.6) becomes negative for $t_{M}\in(t_{M}^{3},1]$, the non-haven splits its tax rate in response to the haven’s split tax rates.
+Therefore, the tax-competition equilibrium is that both the non-haven and the haven split their tax rates.
+∎
+Supplemental Appendix of
+‘A Global Minimum Tax for Large Firms Only:
+Implications for Tax Competition’
+Andreas Haufler [^1] Seminar for Economic Policy, LMU Munich, Akademiestr.
+1, 80799 Munich, Germany.
+Email: Andreas.Haufler@econ.lmu.de.
+Hayato Kato [^2] Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan.
+Email: hayato.kato@econ.osaka-u.ac.jp
+October, 2025
+Contents
+Appendix A MNEs Subject to the GMT
+We select from the Orbis database MNEs satisfying the following conditions:
+  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
+  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
+  • Non-missing revenues (‘Operating revenue (Turnover)’) and pre-tax profits (‘P/L before tax’).
+Column (a) of Table A.1 reports the pre-tax profits and the total number of MNEs in each year.
+Column (b) restricts the sample to MNEs subject to the GMT, i.e., those with annual revenues of at least 750 million EUR in two or more of the past four years.
+Using columns (a) and (b), column (c) shows the share of MNEs subject to the GMT in terms of the pre-tax profits and the number of MNEs.
+Table A.1. MNEs in the Orbis Database.
+(a) All MNEs (b) MNEs $\geq$ 750mEUR (c) Share == B/A
+Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
+2018 2320 8656 2066 2333 0.89 0.27
+2019 2430 8416 2216 2437 0.91 0.29
+2020 1985 7920 1793 2228 0.90 0.28
+2021 3564 8803 3239 2627 0.91 0.30
+Source: Orbis database, own calculations.
+Note: Pre-tax profits are in billion EUR.
+Appendix B Decentralized Decision-Making by Tax Havens
+In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate.
+Here, we consider an alternative, decentralized decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others.
+In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions.
+First, countries can split their tax rates only if they are subject to the GMT.
+This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019).
+Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates.
+This allows us to focus on the problem of a single haven.
+Out of the total $H$ tax havens, $H_{M}$ havens commit to a single GMT rate, while the remaining $H-H_{M}$ havens split their tax rates.
+When the non-haven sets a single non-GMT rate and a haven $h$ sets a single GMT rate, their equilibrium tax rates are respectively
+$\begin{split}&t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\\ &t_{h}^{1}(H_{M})=t_{M},\end{split}$
+Let $G_{h}^{1}(H_{M})$ denote the equilibrium payoff of the haven $h$ with commitment.
+We note that as in the text, setting a single non-GMT rate is the non-haven’s dominant strategy because if $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, it is never constrained by the GMT: $t_{n}(H_{M})\leq t_{n}(H)=t_{n}^{2}\leq t_{M}$.
+If instead the haven $h$ splits its tax rate while the other havens and the non-haven keep their rates unchanged, the number of havens that maintain commitment falls to $H_{M}-1$.
+The equilibrium tax rates of the non-haven and the haven $h$ are respectively
+$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (B.1)
+Let $G_{h}^{2}(H_{M}-1)$ be the equilibrium payoff of the haven $h$ under splitting.
+The difference between the haven $h$’s payoffs in the two cases is
+$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
+$\ \ =\frac{(1-\phi)\Pi[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta(H_{M})}{\delta\left[\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})\right]\left[(\lambda-1)\left\{\phi(H-H_{M}+1)+H+H_{M}-1\right\}+2H\lambda\right]^{2}},$
+where
+$\Theta(H_{M})\equiv AH_{M}+B,$
+$A\equiv(\lambda-1)(1-\phi)\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right],$
+$B\equiv H[\phi(\lambda-1)+3\lambda-1]\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right]-2t_{M}(\lambda-1)^{2}(1-\phi).$
+If $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$, the haven $h$ always sets a single GMT rate, regardless of how many other havens it believes set a single GMT rate.
+Under our assumption of symmetric belief formation, all havens end up setting a single GMT rate, and the equilibrium coincides with Regime 1 in the text.
+Conversely, if $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$, all havens end up splitting their tax rates, and the equilibrium coincides with Regime 2 in the text.
+Since the denominator of $\Delta G_{h}^{12}(H_{M})$ is positive and $Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0$ for $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, the sign of $\Delta G_{h}^{12}(H_{M})$ is determined by that of $\Theta(H_{M})$.
+Considering that $\Theta(H_{M})$ is linear in $H_{M}$, the conditions for $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$ reduce to the following two conditions:
+$\Theta(1)=A\cdot 1+B=C-Dt_{M}\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{a}\equiv C/D,$
+$\Theta(H)=AH+B=2(E-Ft_{M})\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{b}\equiv E/F,$
+where
+$C\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
+$D\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
+$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
+$F\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
+We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
+$t_{M}^{a}-t_{M}^{0}=\frac{2\delta(\lambda-1)^{2}[\phi(\lambda-1)+3\lambda-1]}{D(3\lambda-1)}>0,$
+$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
+$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}[\phi(H+1)(\lambda-1)+\lambda(3H-1)-(H-1)]}{F[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0.$
+We can also confirm $\Theta(H_{M})=AH_{M}+B$ increases with $H_{M}$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$, or equivalently $A>0$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$.
+Since $A>0$ if $t_{M}<t_{M}^{c}\equiv\delta(\lambda-1)/[H(3\lambda-1)-2(\lambda-1)]$, it suffices to show $t_{M}^{b}<t_{M}^{c}$:
+$t_{M}^{c}-t_{M}^{b}=\frac{\delta(\lambda-1)[\lambda^{2}+(H-1)(2\lambda-1)-\phi(\lambda-1)^{2}]}{F[H(3\lambda-2)-(H-1)]}>0,$
+which is indeed positive.
+Using these results, we can thus conclude that (i) if $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$; (ii) if $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$; and (iii) if $t_{M}\in(t_{M}^{a},t_{M}^{b}]$, there exists a positive integer $H_{M}^{*}\in\{2,\dots,H\}$ such that $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,\dots,H_{M}^{*}-1\}$ and $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{H_{M}^{*},\dots,H\}$.
+In cases (i) and (ii), the equilibrium is unique.
+In case (iii), there can be at most three equilibria.
+First, if each haven believes that fewer than $H_{M}^{*}$ havens set a single GMT rate, all havens end up splitting their tax rates, i.e., $H_{M}=0$.
+Second, if each haven believes that more than $H_{M}^{*}-1$ havens set a single GMT rate, all havens end up doing so, i.e., $H_{M}=H$.
+Third, if $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, then $H_{M}=H_{M}^{*}$ havens set a single GMT rate while the remaining $H-H_{M}^{*}$ havens split their tax rates, consistent with their beliefs.
+The discussion is summarized as follows.
+If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)).
+If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)).
+If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge.
+If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (B.1)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (B.1)).
+A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralized manner, implying that they are less likely to commit to a single GMT rate.
+This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
+Appendix C Welfare Effects of the Introduction of GMT
+The following proposition characterizes the welfare effects of introducing a binding GMT rate, compared with the unconstrained equilibrium (Regime 0).
+Proposition C.1
+(GMT introduction)
+Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text, and assume the GMT coverage rate is sufficiently high, such that $\phi\in(3/4,1)$.
+Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
+  • (i)
+    a rise in welfare in the non-haven for all Regimes 1 to 4.
+  • (ii)
+    a rise in welfare in the (representative) haven for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
+  • (iii)
+    a rise in world welfare for all Regimes 1 to 4.
+These results hold for any value of the marginal value of public funds (MVPF), $\lambda\in(1,\infty)$.
+In the limit as $\lambda\to\infty$, countries care only about tax revenues.
+Hence, qualitatively identical results apply to tax revenues.
+Proof: In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
 Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
@@
 $G_{h}^{1}-G_{h}^{0}=\frac{\lambda\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(2\lambda-1)-Ht_{M}\lambda(3\lambda-1)\right]}{\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi\equiv\Delta$
-$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\end{cases},$
+$\Delta\ \begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}\equiv t_{M}^{+}\\ <0&\text{if}\ t_{M}>t_{M}^{+}\end{cases},$
 $G_{W}^{1}-G_{W}^{0}=\frac{\left[Ht_{M}(3\lambda-1)-\delta(\lambda-1)\right]\left[\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)-Ht_{M}\lambda^{2}(3\lambda-1)\right]}{2\delta H(2\lambda-1)(3\lambda-1)^{2}}\Pi>0,$
@@
 $t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(11\lambda^{2}-10\lambda+2)}{H\lambda^{2}(3\lambda-1)}.$
-The marginal effect of the GMT is
-$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,\qquad\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\\ <0&\text{if}\ t_{M}>\frac{\delta(\lambda-1)}{2H\lambda}\end{cases},$
+In Regime 1, the welfare effect of the introduction a GMT rate is positive in the non-haven and the world.
+In the haven, it is positive if $t_{M}\in[t_{M}^{0},t_{M}^{+})$ and negative if $t_{M}\in(t_{M}^{+},t_{M}^{1}]$.
+A marginal increase in $t_{M}$ leads to
+$\frac{\partial G_{n}^{1}}{\partial t_{M}}=\frac{H\lambda^{2}t_{M}+\delta(\lambda-1)^{2}}{\delta(2\lambda-1)}\Pi>0,$
+$\frac{\partial G_{h}^{1}}{\partial t_{M}}=\frac{\lambda[\delta(\lambda-1)-2H\lambda t_{M}]}{\delta(2\lambda-1)}\Pi\begin{cases}\geq 0&\text{if}\ t_{M}\leq\frac{\delta(\lambda-1)}{2H\lambda}\equiv t_{M}^{++}\\ <0&\text{if}\ t_{M}>t_{M}^{++}\end{cases},$
 $\frac{\partial G_{W}^{1}}{\partial t_{M}}=\frac{\delta(\lambda-1)(2\lambda-1)-H\lambda^{2}t_{M}}{\delta(2\lambda-1)}\Pi>0,$
@@
 $t_{M}\leq t_{M}^{1}<\frac{\delta(\lambda-1)(2\lambda-1)}{2H\lambda}.$
-In Regime 1, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
-On the other hand, the welfare level of the haven rises if $t_{M}<\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$ after the GMT introduction and it falls if $t_{M}>\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$.
-Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
-For the non-haven country, we see
+In Regime 1, the welfare effect of a marginal increase in the GMT rate is positive in the non-haven and the world.
+In the haven, it is positive if $t_{M}\in[t_{M}^{0},\delta(\lambda-1)/(2H\lambda))$ and negative if $t_{M}\in(\delta(\lambda-1)/(2H\lambda),t_{M}^{1}]$.
+Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
+For the non-haven, we see
 $\frac{\partial G_{n}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{1}}=\frac{(\lambda-1)[8\lambda^{3}-5\lambda^{2}-2\lambda+1+\phi(\lambda-1)^{2}]}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi>0,$
@@
 which holds for any $\lambda>1$ and $\phi\in(0,1)$.
-For the haven country, we get
+For the haven, we get
 $\frac{\partial G_{h}^{2}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{1}}=-\frac{2\lambda(2\lambda-1)(\lambda-1)^{2}}{[3\lambda-1+\phi(\lambda-1)][\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}\phi\Pi<0,$
@@
 which holds for all $\lambda>1$ and $\phi\in(0,1)$.
-In Regime 2, the welfare levels of the non-haven and the world therefore rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
-On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
-Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
-For the non-haven country, we have
+In Regime 2, both the introduction and a marginal increase of the GMT rate raise welfare in the non-haven and the world, but reduce it in the haven.
+Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
+For the non-haven, we have
 $\frac{\partial G_{n}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{(\lambda-1)[2\lambda-1+\phi(2\lambda+1)]}{2[3\lambda-1-\phi(\lambda-1)]}\Pi>0,$
@@
 $G_{n}^{3}-G_{n}^{0}\big|_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(16\lambda^{3}-13\lambda^{2}+3)-\phi(\lambda-1)^{2}(8\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
-For the haven country, we see
+For the haven, we see
 $\frac{\partial G_{h}^{3}}{\partial t_{M}}\bigg|_{t_{M}=t_{M}^{2}}=\frac{\lambda(\lambda-1)}{3\lambda-1-\phi(\lambda-1)}(1-\phi)\Pi>0,$
@@
 $G_{W}^{3}-G_{W}^{0}\big|_{t_{M}=t_{M}^{2}}=\frac{\delta(\lambda-1)^{2}[(3\lambda-1)(14\lambda^{2}-15\lambda+3)-\phi(\lambda-1)^{2}(10\lambda-3)]}{2H(3\lambda-1)^{2}[3\lambda-1-\phi(\lambda-1)]^{2}}\phi\Pi>0.$
-In Regime 3, the welfare levels of the non-haven, the haven, and the world all rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
+In Regime 3, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven.
+A marginal increase in the GMT rate raises welfare in the non-haven, the haven and the world.
 Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
-For the non-haven country, we see
+For the non-haven, we see
 $G_{n}^{4}-G_{n}^{0}\big|_{t_{M}=t_{M}^{3}}=\frac{\delta(\lambda-1)^{2}(8\lambda-3)^{2}}{2H(3\lambda-1)^{2}(4\lambda-1)}\phi\Pi>0,$
@@
 With these observations and the fact that $G_{n}^{4}-G_{n}^{0}$ is linear in $t_{M}$, we conclude that $G_{n}^{4}-G_{n}^{0}>0$ for $t_{M}\in(t_{M}^{3},1]$.
-For the haven country, we have
+For the haven, we have
 $G_{h}^{4}-G_{h}^{0}=-\frac{\delta\lambda(\lambda-1)^{2}}{H(3\lambda-1)^{2}}\phi\Pi>0,$
@@
 Combining these results leads to $G_{W}^{4}-G_{W}^{0}>0$ and $\partial G_{W}^{4}/\partial t_{M}>0$.
-In Regime 4, the welfare levels of the non-haven and the world rise after the GMT introduction and they increase along with a gradual increase in the GMT rate.
-On the other hand, the welfare level of the haven declines after the GMT introduction and it is independent of $t_{M}$.
-This proves Proposition 3 and a part of Proposition 4. The proof of the results on the regime-switching points is given in Appendix 3.
-A.4. Regime-switching points
+In Regime 4, the introduction of a GMT rate increases welfare in the non-haven and the world, while reducing it in the haven.
+A marginal increase in the GMT rate raises welfare in the non-haven and the world, but has no effect on welfare in the haven.
+∎
+Appendix D Welfare Effects of Increasing the GMT Rate
+The following proposition characterises the welfare effects of a marginal increase in the GMT rate $t_{M}$.
+Proposition D.1
+(Increase in the GMT rate $t_{M}$)
+Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. A marginal increase in the GMT rate has the following effects:
+  • (i)
+    welfare in the non-haven increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it falls discontinuously.
+  • (ii)
+    welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2.
+It increases in Regime 3, falls discontinuously at the regime-switching point $t_{M}^{3}$, and remains unchanged in Regime 4.
+  • (iii)
+    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
+As is true for the introduction of the GMT (Proposition C.1 in Section C of this Supplemental Appendix), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium.
+An exception is the regime-switching point $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs.
+Small MNEs then shift more profits to the havens, hurting the non-haven.
+For havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven.
+These opposing forces generate a non-monotonic welfare response.
+In Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{1}]$), the positive effect dominates initially (when $t_{M}<t_{M}^{++}$), but as the GMT rate rises further, the negative effect prevails (when $t_{M}\geq t_{M}^{++}$).
+In Regime 2 ($t_{M}\in(t_{M}^{1},t_{M}^{2}]$), the negative effect dominates throughout.
+In Regime 3 ($t_{M}\in(t_{M}^{2},t_{M}^{3}]$), the positive effect of a higher tax rate dominates again, as the non-haven commits to a uniform GMT rate in order to raise the havens’ tax rates.
+This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$).
+It ends at the regime-switching point $t_{M}^{3}$ at which the non-haven starts splitting its tax rate.
+This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens.
+In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits.
+Consequently, changes in the GMT rate no longer affect the havens’ welfare.
+The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive.
+At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs.
+Therefore, tax competition is tightened and world welfare falls at the margin.
+In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
+Proof:  In the proof of Proposition C.1 in Section C, we already showed the welfare effects of a marginal increase in the GMT rate within each regime.
+Here we show the welfare effects at the regime switching points.
 At the three regime switching GMT rates, $t_{M}^{1}$, $t_{M}^{2}$ and $t_{M}^{3}$, welfare in the non-haven, the haven, and the world changes as follows.
@@
 $G_{h}^{4}-G_{h}^{3}\big|_{t_{M}=t_{M}^{3}}=G_{W}^{4}-G_{W}^{3}\big|_{t_{M}=t_{M}^{3}}=-\frac{8\delta\lambda(\lambda-1)^{2}(2\lambda-1)}{2H(\lambda-1)[H(3\lambda-1)(4\lambda-1)^{2}]}(1-\phi)\Pi<0.$
-This completes the proof of Proposition 4.
-A.5. Proof of Proposition 5
-The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
+This completes the proof of Proposition D.1. ∎
+Appendix E Welfare Effects of Increasing the GMT Coverage Rate
+The following proposition characterizes the welfare effects of a marginal increase in the GMT coverage rate $\phi$.
+Proposition E.1
+(Increasing in the GMT coverage rate $\phi$)
+Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text.
+Except for special cases where a regime switch occurs, a marginal increase in the GMT coverage rate of $\phi\in(0,1)$ has the following effects:
+  • (i)
+    welfare in the non-haven is unaffected in Regime 1 and increases in Regimes 2 to 4.
+  • (ii)
+    welfare in the (representative) haven is unaffected in Regime 1 and decreases in Regimes 2 to 4.
+  • (iii)
+    world welfare is unaffected in Regime 1 and increases in Regimes 2 to 4.
+In Regime 1, an increase in $\phi$ has no effect on welfare in any country, because all countries levy uniform tax rates.
+Changing $\phi$ does have effects, however, in the other regimes, where one or both countries split their tax rates.
+Since the equilibrium tax differential is always greater for small MNEs than for large MNEs, a higher $\phi$, which implies more large MNEs subject to the GMT, reduces aggregate profit shifting.
+This unambiguously hurts the haven and benefits the non-haven.
+From the global perspective, a higher GMT coverage is always desirable, as it reduces pressure on tax competition for the profits of small MNEs.
+There are two exceptional cases where increasing $\phi$ causes a regime switch.
+The first is at $t_{M}=t_{M}^{1}+\epsilon$ with $\epsilon$ being a small positive number, where a rise in $\phi$ pushes Regime 2 back to Regime 1 by weakening the havens’ incentive to split their tax rates.
+This regime switch benefits the non-haven and the world, while it has no effect on the havens.
+The other is at $t_{M}=t_{M}^{2}+\epsilon$, where a higher $\phi$ triggers a switch from Regime 3 to Regime 2, inducing the non-haven to set a tax rate above the GMT.
+This latter regime switch, however, leaves welfare in all countries and the world unaffected.
+In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3).
+A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1.
+One notable difference from the effects of a higher $t_{M}$ is that a higher $\phi$ can never harm the non-haven country and the world, even if it induces a regime switch.
+In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
+Proof:  The effects of an increase in the GMT coverage rate, $\phi$, on welfare in the non-haven, the haven, and the world are given as follows.
 Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
@@
 $\frac{\partial G_{n}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{h}^{1}}{\partial\phi}=0,\quad\frac{\partial G_{W}^{1}}{\partial\phi}=0.$
-Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
-For the non-haven country, we see
+Regime 2: $t_{M}\in(t_{M}^{1},t_{M}^{2}]$, where $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$ and $t_{M}^{2}=\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$.
+For the non-haven, we see
 $\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\Theta_{n}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{2\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
@@
 The sign of the derivative is determined by that of $\Theta_{n}$.
-From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{1})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
-For the haven country, we see
+From $\Theta_{n}(t_{M}=t_{M}^{1})>0$, $\Theta_{n}(t_{M}=t_{M}^{2})>0$, and the fact that $\Theta_{n}(t_{M})$ is linear in $t_{M}$, it follows that $\Theta_{n}>0$.
+For the haven, we see
 $\frac{\partial G_{h}^{2}}{\partial\phi}=\frac{\lambda\Theta_{h}[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]}{\delta H[3\lambda-1+\phi(\lambda-1)]^{3}}\Pi,$
@@
 The sign of the derivative is determined by that of $\Theta_{h}$.
-From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{1})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
+From $\Theta_{h}(t_{M}=t_{M}^{1})<0$, $\Theta_{h}(t_{M}=t_{M}^{2})<0$, and the linearity of $\Theta_{h}(t_{M})$ in $t_{M}$, it follows that $\Theta_{h}<0$.
 For the world, we get
@@
 The sign of the derivative is determined by that of $\Theta_{W}$.
-From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{1})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
+From $\Theta_{W}(t_{M}=t_{M}^{1})>0$, $\Theta_{W}(t_{M}=t_{M}^{2})>0$, and the linearity of $\Theta_{W}(t_{M})$ in $t_{M}$, it follows that $\Theta_{W}>0$.
 Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
 We see
-$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\quad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\quad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
+$\frac{\partial G_{n}^{3}}{\partial\phi}=\frac{H(4\lambda-1)t_{M}^{2}}{8\delta}\Pi>0,\qquad\frac{\partial G_{h}^{3}}{\partial\phi}=-\frac{H\lambda t_{M}^{2}}{8\delta}\Pi<0,\qquad\frac{\partial G_{W}^{3}}{\partial\phi}=\frac{H(2\lambda-1)t_{M}^{2}}{4\delta}\Pi>0,$
 Regime 4: $t_{M}\in(t_{M}^{3},1]$, where $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
@@
 noting that $t_{M}>t_{M}^{3}>\frac{\delta(\lambda-1)(10\lambda-3)}{2H(3\lambda-1)^{2}}>\frac{\delta(\lambda-1)(8\lambda-3)}{2H(3\lambda-1)^{2}}$.
-This completes the proof of Proposition 5.
-Appendix B: Data for Figure 1
-We select MNEs from the Orbis based on the following criteria.
-  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
-  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
-The numbers on which Figure 1 is based are given in Table B1.
-In column A, to calculate the pre-tax profits of all MNEs in a given year, we exclude those with missing revenues (“Operating revenue (Turnover)”) and pre-tax profits (“P/L before tax”).
-In column B, MNEs subject to the GMT are those with annual revenues of no less than 750 million EUR in at least two years of the last four years.
-Using column A and B, column C reports the share of MNEs subject to the GMT in terms of pre-tax profits and numbers.
-Table B1: MNEs in the Orbis database
-A: All MNEs B: MNEs $\geq$ 750mEUR C: Share $=$ B/A
-Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
-2018 2320 8656 2066 2333 0.89 0.27
-2019 2430 8416 2216 2437 0.91 0.29
-2020 1985 7920 1793 2228 0.90 0.28
-2021 3564 8803 3239 2627 0.91 0.30
-Source: Orbis database, own calculations.
-Note: Pre-tax profits are in billion EUR.
+This completes the proof of Proposition E.1. ∎
+Appendix F Supplementary Figures
+Figure F.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
+The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those under full GMT coverage (dashed curves).
+The equilibrium tax rates under full GMT are as follows.
+Regime 0: $t_{M}\in\left[0,t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)}\right)$.  Each of the non-haven and the haven sets its own single non-GMT rate:
+$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Regime 1f: $t_{M}\in\left[t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)},t_{M}^{f}\equiv\frac{\delta(\lambda-1)}{H\lambda}\right]$.
+The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
+$t_{n}^{1f}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Regime 2f: $t_{M}\in(t_{M}^{f},1]$.  Both the non-haven and the haven choose a single GMT rate:
+$t_{n}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the haven under full GMT.
+In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
+We note $t_{M}^{f}>t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$, given $\phi\in(0,1)$ and $\lambda>1$.
+That is, the equilibrium outcomes under partial GMT begins to diverge from those under full GMT once $t_{M}$ enters the range of Regime 1.
+This is particularly evident in panels (a) and (c).
+Fig. F.1. Equilibrium Welfare and Shifted Profits.
+Notes: The figure shows the equilibrium levels of welfare in the non-haven ($G_{n}$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$.
+The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0).
+While the levels of welfare, shifted profits and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.

### v7

@@
 A Global Minimum Tax for Large Firms Only:
-A GLOBAL MINIMUM TAX FOR LARGE FIRMS ONLY:
-IMPLICATIONS FOR TAX COMPETITION***Andreas Haufler, Seminar for Economic Policy, LMU Munich, Akademiestr.
-1, 80799 Munich, Germany.
-Email: Andreas.Haufler@econ.lmu.de.
-Hayato Kato, Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan.
-Email: hayato.kato@econ.osaka-u.ac.jp.
- Financial support from the German Research Foundation through CRC TRR 190 (Haufler) and the JSPS KEKENHI (JP20H01495, JP22K13388, JP24H00014) (Kato) is gratefully acknowledged.
-This study is conducted as a part of the Project “Economic Policy Issues in the Global Economy” undertaken at the Research Institute of Economy, Trade and Industry (RIETI).
-Kato thanks the University of Munich for their hospitality while part of this paper was written.
-We thank Tomohiro Ara, Real Arai, Ruby Doeleman, Antoine Ferey, Emanuel Hansen, Jim Hines, Niels Johannesen, Makoto Hasegawa, Jota Ishikawa, Mick Keen, Sang-Hyun Kim, Yoshimasa Komoriya, Jonas Löbbing, Mohammed Mardan, Jakob Miethe, Kenichi Nishikata, Hirofumi Okoshi, Pascalis Raimondos, Dirk Schindler, Karl Schulz, Georg Thunecke, and Alfons Weichenrieder for many helpful comments and discussions.
-Thanks also to seminar and conference participants at the IIPF in Utah State, MPI Berlin, Goettingen, Osaka, Kyoto, Shimane, QUT Brisbane, Ryukoku, Gakushuin, CESifo (Munich & Venice Summer Institute), and RIETI for helpful comments.
-Andreas Haufler and Hayato Kato
+Abstract
 The Global Minimum Tax (GMT) is applied only to firms above a certain size threshold, permitting countries to set differential tax rates for small and large firms.
 We analyse tax competition among multiple tax havens and a non-haven country for heterogeneous multinationals to evaluate the effects of this partial coverage of GMT.
-Upon the introduction of a moderately low GMT rate, the havens commit to the single uniform GMT rate for all multinationals.
+Upon the introduction of a low but binding GMT rate, the havens commit to the single uniform GMT rate for all multinationals.
 However, gradual increases in the GMT rate induce the havens, and subsequently the non-haven, to adopt discriminatory, lower tax rates for small multinationals.
-Our calibration exercise shows that introducing a GMT rate of 15% results in a regime where only the havens adopt split tax rates.
+Our calibration exercise shows that introducing a GMT rate close to 15% results in a regime where only the havens adopt split tax rates.
 Welfare and tax revenues fall in the havens but rise in the non-haven, yielding a positive net gain worldwide.
+JEL classification: F23; F25; H87
+Keywords: Multinational firms; Tax avoidance; Global minimum tax; Profit shifting
+1 Introduction
 Profit shifting by multinational enterprises (MNEs) remains a central challenge for international corporate taxation.
-Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., , 2023), a pattern also documented in microdata for France (Davies et al., , 2018), the UK (Bilicka, , 2019), and the US (Guvenen et al., , 2022) In response to the resulting revenue losses, the OECD launched its Base Erosion and Profit Shifting (BEPS) initiative (OECD, , 2013), whose core outcome is the introduction of a global minimum tax (GMT), also known as Pillar Two.
-[^1] See Fuest and Neumeier, (2023); Hebous et al., (2024) for the institutional background.
-Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., , 2025; Ogawa and Tsuchiya, , 2025; Richter, , 2025).
-In 2021, more than 130 countries including many tax havens agreed on a 15% GMT; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT.
-[^2] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed October 1, 2025).
-Despite announcing after the June 2025 G7 summit that it would not fully comply with the GMT, the US had already implemented a similar tax on Global Intangible Low-Tax Income under the 2017 Tax Cuts and Jobs Act (Chodorow-Reich et al., , 2024).
-A limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years.
-[^3] Since 2016, such groups have been required to file country-by-country reports on their activities, profits, and taxes in each country of operation (Doeleman et al., , 2024, Section 2).
-The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, , 2022, p.14).
-According to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT.
-[^4] See Supplemental Appendix A.
-This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, , 2020, # 505, p.
+Aggregate estimates suggest that more than one-third of foreign-earned corporate profits are shifted to tax havens (Tørsløv et al., 2023).
+[^1] Combining micro- and macro-level data on portfolio investment, Coppola et al.
+(2021) highlight the role of tax havens as conduits through which firms access international capital markets.
+See Camarero et al.
+(2025) for an application of their dataset in the context of the EU.
+The important role of tax havens for profit shifting is confirmed in other studies using microdata of multinational affiliates in France (Davies et al., 2018), the U.K.
+(Bilicka, 2019), and the U.S.
+(Guvenen et al., 2022).
+In response to the large revenue losses caused by profit shifting, the OECD has launched an action plan to fight base erosion in OECD countries, and in particular the profit shifting to tax havens (OECD, 2013).
+A core development in this endeavor is the introduction of a global minimum tax (GMT), also known as Pillar Two.
+[^2] See Fuest and Neumeier (2023); Hebous et al.
+(2024) for the institutional background.
+Another outcome of the BEPS initiative is Pillar One, which reallocates taxing rights from the country of a MNE’s residence to the market jurisdiction (Mukunoki et al., 2025; Ogawa and Tsuchiya, 2025; Richter, 2025).
+Pillar One aims to address the tax challenges posed by the digital economy (see Lassmann et al., 2025 for how multinational digital platforms respond to corporate taxes).
+In 2021, more than 130 countries including many tax havens agreed on a 15% GMT rate; by August 2025, 65 countries had drafted or enacted legislation implementing the GMT.
+[^3] See Tax Foundation, ‘The Latest on the Global Tax Agreement’: https://taxfoundation.org/blog/global-tax-agreement/ (accessed January 29, 2026).
+The United States had already implemented, in its 2017 Tax Cuts and Jobs Act, a tax on Global Intangible Low-Tax Income (GILTI) that works in many respects like a GMT (Chodorow-Reich et al., 2024).
+An important limitation of the GMT is that it applies only to ‘large’ multinationals, defined as groups with at least 750 million EUR in annual revenues in two of the last four years.
+[^4] Since 2016, such groups have been required to file country-by-country reports (CbCR) on their activities, profits, and taxes in each country of operation (Doeleman et al., 2024, Section 2).
+The revenue threshold reflects the balance between limiting profit shifting and minimising compliance costs (OECD, 2022, p.14).
+Fig.
+1 shows that, according to the Orbis database of Bureau van Dijk, about 30% of MNEs, accounting for roughly 90% of total profits, exceeded this threshold in 2018–2021 and were thus covered by the GMT.
+[^5] This figure aligns with the OECD’s estimate of 90% profit coverage (OECD, 2020, # 505, p.
 233).
 Nevertheless, over 300 billion EUR of MNEs’ profits remain outside its scope.
-Moreover, Orbis tends to oversample large firms (Bajgar et al., , 2020), so the 90% coverage rate should be viewed as an upper bound.
-Does this incomplete coverage affect the effectiveness of the GMT?
-It does not if low-tax countries apply the GMT rate to all firms regardless of size.
-It does matter, however, if they differentiate their tax rates, applying the GMT rate to firms above the threshold and a lower rate to those below.
-Such a split is inherent in the regulation: tax havens may implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a top-up tax.
-[^5] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT) (Devereux, , 2023).
-This system has been adopted in Bahrain, Ireland, Switzerland and the UAE, among other countries with base statutory corporate tax rates below 15% (PwC, , 2025).
-Against this background, we ask when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT.
-Does a 15% GMT rate induce such a split, and how large are the resulting revenue and welfare gains?
-How do these outcomes change with broader GMT coverage?
+Moreover, Orbis tends to oversample large firms (Bajgar et al., 2020), and hence the 90% coverage rate should be seen as an upper bound.
+One implication of this incomplete coverage is that low-tax countries might respond to the introduction of the GMT by using a split corporate tax system, where a lower tax rate applies to firms below the GMT threshold.
+In fact, such a split is inherent in the regulation: tax havens can implement the GMT by raising their (low) base tax rates only for MNEs above the threshold through a specific top-up tax.
+[^6] This is known as the Qualified Domestic Minimum Top-Up Tax (QDMTT).
+See Devereux (2023) for a discussion.
+Tax havens are fully aware of this option.
+Ireland, for example, has decided to keep its general tax rate at 12.5%, but top up the tax rate to 15% for affiliates of foreign MNEs above the threshold (Government of Ireland, 2023).
+Several other major low-tax countries, including Bahrain, Bulgaria, Hungary, Switzerland and the UAE have adopted similar split tax systems by joining the GMT agreement while maintaining their base statutory corporate tax rates below 15% (PwC, 2025).
+[^7] The latest corporate tax rates around the world can be found in ‘Tax Foundation, ‘Corporate Tax Rates Around the World, 2025’: https://taxfoundation.org/data/all/global/corporate-tax-rates-by-country-2025/ (accessed January 29, 2026).
+Note that not all low-tax countries have adopted a split tax system: Cyprus and Gibraltar, for example, raised their base rates from 12.5% to 15% for both large and small MNEs, effective in 2026 and 2024, respectively.
+See Shanda Consult, ‘The Cyprus Tax Reform 2026: A Strategic Evolution for Business and Compliance’: https://shandaconsult.com/cyprus-tax-reform-2026 and PwC, ‘Gibraltar—Taxes on corporate income’: https://taxsummaries.pwc.com/gibraltar/corporate/taxes-on-corporate-income (accessed January 29, 2026).
+Fig. 1. GMT coverage.
+Source: Orbis database, own calculations.
+Notes: The left bar in each year shows the share of MNEs covered by the GMT out of all MNEs.
+The right bar in each year shows the share of pre-tax profits earned by MNEs covered by GMT out of pre-tax profits by all MNEs.
+The MNEs covered by GMT are those whose annual revenues exceed $750$ million EUR in at least two of the last four years out of all MNEs.
+See A for details.
+These responses show that the effects of the GMT cannot be analysed assuming a single, uniform corporate tax rate across all countries.
+Against this background, this paper asks when tax competition leads countries to differentiate tax rates for large and small MNEs under the GMT, whether a 15% GMT rate induces such a split, and how large are the resulting revenue and welfare gains.
 To answer these questions, we build a simple model where heterogeneous MNEs headquartered in a non-haven country shift profits to tax-haven countries.
 Countries maximise welfare—a weighted sum of tax revenues and, for the non-haven, the post-tax income of MNEs’ owners—and compete for the MNEs’ profits.
-In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rates for all MNEs.
+In the first stage, both the non-haven and the havens decide whether to commit to a uniform tax rate for all MNEs.
 In the second stage, countries set tax rates, either uniform or split, with large MNEs taxed at least at the GMT rate.
@@
 When the GMT rate is binding but low, the havens commit to it for all MNEs to raise the non-haven’s uniform tax rate.
-As the GMT rate rises, the havens begin to split their rates, applying the GMT to large MNEs and a lower rate to small MNEs.
+As the GMT rate rises, the havens start splitting their rates, applying the GMT to large MNEs and a lower rate to small MNEs.
 For sufficiently high GMT rates, the non-haven also splits its rate, as limited competition for large MNEs’ profits shifts the focus towards small MNEs.
-We calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms.
-A 15% GMT rate induces the havens to split their tax rates, which is consistent with observed responses in some tax havens including Ireland and Switzerland.
-It benefits the non-haven while harms the havens, generating net global revenue (welfare) gains of 136 billion USD (83 bUSD), comparable to existing estimates (Baraké et al., , 2022; Hugger et al., , 2024).
-If the GMT covers all firms, the 15% rate does not produce a split tax system but increases revenues and welfare in the non-haven and globally even more.
-This global revenue gain would have to be set against the additional compliance costs for multinationals, however.
-We contribute to the recent theoretical studies on tax competition under a GMT.
-They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, , 2022; Hebous and Keen, , 2023); real economic activities (Janeba and Schjelderup, , 2023; Schjelderup and Stähler, , 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, , 2025).
+We then calibrate our model using estimates of profit shifting and examine the quantitative effects of GMT reforms.
+For our basic model, a 15% GMT rate leads both the non-haven and the havens to choose uniform rates for all firms, whereas a 16% GMT rate induces the havens to split their tax rates.
+In an extended model where tax havens choose their tax rates in a decentralized way, a 15% GMT rate is sufficient to induce a split tax system for the havens in equilibrium.
+This splitting result is consistent with the observed responses in several tax havens, including Ireland and Switzerland, as discussed earlier.
+Our quantitative results also show that introducing a 15% GMT rate benefits the non-haven while harming the havens, generating net global revenue gains of 149–156 billion USD (and welfare gains of 92–96 billion USD).
+These revenue gains are comparable to existing estimates in the literature (Baraké et al., 2022; Hugger et al., 2024).
+Our analysis contributes to the recent theoretical studies on tax competition under a GMT.
+They ask how introducing a GMT affects national and global revenues and welfare in the presence of country heterogeneity (Johannesen, 2022; Hebous and Keen, 2023); real economic activities (Janeba and Schjelderup, 2023; Schjelderup and Stähler, 2024); and endogenous tax-enforcement levels (Hindriks and Nishimura, 2025).
 All these models assume the GMT applies uniformly, ignoring the firm-size threshold below which firms are exempt from the GMT.
 Our model highlights this threshold, allowing countries to tax firms above and below it differently.
-[^6] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, , 1999; Keen, , 2001; Bucovetsky and Haufler, , 2007; Gaigné and Wooton, , 2011).
-Our simple quantitative exercise complements recent evaluations of corporate taxation using trade and multinational models (Wang, , 2020; Ferrari et al., , 2023; Shen, , 2024; Dyrda et al., , 2024), where taxes are assumed to be exogenous and/or uniform across firms.
-By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
-Two empirical papers (Bachas et al., , 2025; Clifford et al., , 2025) closely related to ours examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively.
-Their findings that GMT revenue gains depend strongly on its coverage emphasizes the relevance of our analysis.
+[^8] This links our model to the literature on discriminatory tax competition, where countries can tax firms differently by their mobility (Janeba and Peters, 1999; Keen, 2001; Peralta et al., 2006; Gaigné and Wooton, 2011; Haufler and Runkel, 2012).
+Our simple quantitative exercise yields results that are comparable to more detailed estimates of the revenue effects of a GMT (Baraké et al., 2022; Hugger et al., 2024).
+It also complements recent evaluations of corporate taxation using trade and multinational models (Wang, 2020; Ferrari et al., 2023; Shen, 2024; Dyrda et al., 2024; Santacreu, 2026), where taxes are assumed to be exogenous and/or uniform across firms.By contrast, we abstract from firms’ real activities and focus on the strategic role of size-dependent taxes.
+One of the few papers offering empirical evidence on minimum tax regimes is Buettner and Poehnlein (2024), who study the effects of introducing a minimum local business tax rate in Germany.
+Finally, two recent empirical papers (Bachas et al., 2025; Clifford et al., 2025) examine which firms fall within the scope of the GMT using firm/affiliate level data for 16 countries and Germany, respectively.
+Their findings that GMT revenue gains depend strongly on its coverage emphasize the relevance of our analysis.
+We proceed as follows.
+In Section 2 we set up our model of tax competition.
+Section 3 characterises the sub-game perfect Nash equilibria that result in different tax regimes.
+Section 4 turns to the welfare effects of introducing a GMT, and of increasing its tax rate and coverage.
+Section 5 calibrates our model to real-world data and quantifies its effects.
+Section 6 concludes.
 2 The model
@@
 There is a continuum of heterogeneous MNEs which are headquartered in the non-haven and have affiliates in all havens.
-Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven.
-[^7] Firm heterogeneity could result from (executives’) productivity (e.g., Melitz, , 2003; Lippi and Schivardi, , 2014).
-MNEs can shift some of these profits to the havens to maximise their post-tax profits.
-In the following, we assume collective decision-making by the havens.
-This assumption is relaxed in the quantitative analysis later (see also footnote 16) and further examined in Supplemental Appendix B.
+[^9] To focus on the aggregate level of profit shifting relevant for countries’ corporate tax revenues, we abstract from the endogenous location choice of MNEs.
+On the evolution of US-owned MNEs over time and space, see Guvenen et al.
+(2022) for profit shifting and Garetto et al.
+(2024) for real economic activities.
+Headquarters earn (‘true’) exogenous profits $\pi\in[\underline{\pi},\infty)$ with distribution $F(\pi)$, arising from real activity in the non-haven.MNEs can shift some of these profits to the havens to maximise their post-tax profits.
+In the following, we assume (i) true profits $\pi$ independent of taxes and (ii) the collective decision-making by the havens for analytical simplicity.
+These assumptions are relaxed in the quantitative analysis in Section 5.2, where we explore the robustness of our results.
 We assume the non-haven and all havens have signed the GMT agreement and must set tax rates no lower than the GMT rate $t_{M}$.
-[^8] Devereux, (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates.
+[^10] Devereux (2023) analyses how the GMT’s institutional design, especially the Under-Taxed Payments Rule (UTPR), incentivizes non-haven and haven countries to join once a critical mass of large non-havens participates.
 The GMT applies only to MNEs with profits $\pi\geq\pi_{M}$ (‘large MNEs’), allowing countries to set lower rates for firms with $\pi<\pi_{M}$ (‘small MNEs’).
@@
 We analyse a three-stage game.
-In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs—at or above the GMT rate—or to allow differentiated rates.
-[^9] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 6 and in related work on endogenous timing in games (Huck et al., , 2002; Kempf and Rota-Graziosi, , 2010).
+In the first stage, the non-haven and havens simultaneously decide whether to commit to a single tax rate for all MNEs that is at or above the GMT rate, or to allow differentiated rates.
+[^11] The ability to commit, a realistic feature of international tax relations, is standard both in the studies cited in footnote 8 and in related work on endogenous timing in games (Huck et al., 2002; Kempf and Rota-Graziosi, 2010).
 The choice between a uniform and a split tax system is more fundamental, and less frequently revised, than the choice of tax rates.
@@
 In the final stage, MNEs shift profits given these tax rates.
-The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium, referred to as the tax-competition equilibrium.
+The game is solved by backward induction, yielding a sub-game perfect Nash equilibrium.
+We also refer to this as the tax-competition equilibrium.
 2.2 Profit shifting by multinationals
@@
 We assume the costs are quadratic in the level of profit shifting but decrease with the size of a firm: $C_{h}(\theta_{h}\pi)=\delta(\theta_{h}\pi)^{2}/(2\pi)=\delta\theta_{h}^{2}\pi/2$, resulting in total shifting costs of $\sum_{h=1}^{H}C_{h}(\theta_{h}\pi)$.
-[^10] This specification is standard in the literature (e.g., Hines and Rice, , 1994; Huizinga and Laeven, , 2008; Suárez Serrato, , 2018; Bilicka et al., , 2024).
+[^12] This specification is standard in the literature (e.g., Hines and Rice, 1994; Huizinga and Laeven, 2008; Suárez Serrato, 2018; Bilicka et al., 2024).
 The parameter $\delta$ captures the difficulty of profit shifting.
-[^11] This can be higher, for example, due to taxpayer audit (Lediga et al., , 2023).
-Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^12] The GMT tax base equals that of national corporate taxes, ignoring additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler, (2024).
+Given the tax rates $t_{n}$ and $t_{h}$ respectively in the non-haven and each of the havens, an MNE maximises its global post-tax profits net of profit-shifting costs, denoted by $\widetilde{\pi}$, by choosing the share of profit shifting $\theta_{h}$ to each haven: [^13] The GMT tax base equals that of national corporate taxes.
+We thus ignore additional deductions (‘substance-based income exclusion’), an issue examined by Schjelderup and Stähler (2024).
 $\max_{\{\theta_{h}\}_{h=1}^{H}}\ \widetilde{\pi}\left(\{\theta_{h}\}_{h=1}^{H};\pi\right)=(1-t_{n})\left(1-\sum_{h=1}^{H}\theta_{h}\right)\pi+\sum_{h=1}^{H}(1-t_{h})\theta_{h}\pi-\sum_{h=1}^{H}\frac{\delta(\theta_{h}\pi)^{2}}{2\pi}.$ (1)
@@
 The optimal $\theta_{h}$ in (2) implies that MNEs of different sizes share the same tax base elasticity.
-Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, independent of $\pi$.
-Equal tax base elasticities remove countries’ incentives for tax discrimination across MNEs in the absence of the GMT.
+Substituting (2) into (1) and using the symmetry of havens, the tax base of an MNE with $\pi$ in the non-haven becomes $TB_{n}=[1-H(t_{n}-t_{h})/\delta]\pi$, yielding $\varepsilon_{n}=-\frac{dTB_{n}/TB_{n}}{dt_{n}/t_{n}}=\frac{t_{n}}{\delta-H(t_{n}-t_{h})}$, which is independent of $\pi$.
+Equal tax base elasticities thus remove countries’ incentives to tax-discriminate between MNEs in the absence of the GMT.
 This allows us to focus on the splitting responses to a GMT that is applied only to large firms.
-The empirical evidence on the relationship between the size of firms and their responsiveness to tax is inconclusive.
+The empirical evidence on the relationship between the size of firms and their responsiveness to tax is not conclusive.
 Media coverage of tax avoidance by very large multinationals gives an impression that bigger firms have a higher tax-base elasticity.
 However, there are several studies which find the opposite result that the tax-base elasticity is indeed higher for small firms (not necessarily MNEs).
-[^13] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al., (2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms.
-Applying a similar empirical strategy to US firms, Coles et al., (2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size.
-See also Auliffe et al., (2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes.
-Therefore, assuming the tax base elasticity to be independent of firm size may be considered a useful and not unrealistic benchmark.
+[^14] Making use of bunching in the distribution of taxable income of firms in the UK, Devereux et al.
+(2014) find that small-sized firms change their taxable income more significantly in response to statutory tax rate changes, as compared to medium-sized firms.
+Applying a similar empirical strategy to US firms, Coles et al.
+(2022) find that the elasticity of taxable income with respect to effective tax rates is monotonically decreasing in firm size.
+See also Auliffe et al.
+(2023) and the references therein for the heterogeneous tax elasticities of tangible asset investments among firms of different sizes.
+A detailed analysis of French MNEs using Country-by-Country Reports (CbCR) empirically supports our result in eq.
+(2) that firms of different size shift roughly the same share of their profits (Aliprandi et al., 2025, Figure 5).
+This is compatible with the finding that large MNEs are responsible for most of the total amount of profit shifting (Clifford et al., 2025), given their large share of total profits.
+In summary, assuming the tax base elasticity to be independent of firm size seems to be a useful and empirically plausible baseline.
 2.3 Governments’ tax setting choices
@@
 The non-haven country $n$ maximises a weighted sum of private income and tax revenues, where private income are the profits (net of taxes and profit-shifting costs) of all MNEs, whose owners reside exclusively in the non-haven.
-The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, , 2023).
-For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximizing its tax revenues (e.g., Janeba and Schjelderup, , 2023; Hindriks and Nishimura, , 2025).
+The non-haven’s tax revenues are weighted with a factor $\lambda>1$, which represents the marginal value of public funds (MVPF) (Hebous and Keen, 2023).
+For $\lambda\rightarrow\infty$, we obtain the special case of a Leviathan government that is solely interested in maximising its tax revenues (e.g., Janeba and Schjelderup, 2023; Hindriks and Nishimura, 2025).
 The government of the non-haven $n$ therefore maximises
@@
 $\hskip 18.49988pt\hskip 18.49988pt+\lambda t_{n}\left(1-\frac{Ht_{n}-\sum_{h=1}^{H}t_{h}}{\delta}\right)\Pi,$ (3)
-where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ is the total true profits of MNEs.
-Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by their headquarters in the non-haven. This gives
+where $\Pi\equiv\int_{\underline{\pi}}^{\infty}\pi dF$ are the total true profits of MNEs.
+Each of the havens $h$ only cares about its tax revenues (multiplied by the MVPF $\lambda$), as the affiliate located in $h$ is owned by its headquarters in the non-haven. This gives
 $G_{h}=\lambda\int_{\underline{\pi}}^{\infty}t_{h}\theta_{h}\pi dF=\lambda t_{h}\left(\frac{t_{n}-t_{h}}{\delta}\right)\Pi.$ (4)
@@
 The higher tax rate in the non-haven in turn leads each of the havens to raise its tax rate as well.
-[^15] Note the difference between this result and the analysis in Johannesen, (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT.
-Johannesen, (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country.
+[^16] Note the difference between this result and the analysis in Johannesen (2022), where the tax rate in each of the tax havens is zero in the absence of the GMT.
+Johannesen (2022) assumes that shifting profits from a high-tax to a low-tax country incurs zero cost so that MNEs shift their profits only to the lowest-tax country.
 By contrast, we assume that profit shifting from the high-tax country to a given low-tax country has convex costs; hence MNEs diversify their allocation of shifted profits across all havens.
@@
 For lower values of $\lambda$, the equilibrium tax differential is reduced, and the share of shifted profits is accordingly lower.
-2 Theoretical results
-2.1 Tax regimes
+3 Equilibrium tax regimes
 With a GMT in place, in the first stage of the game each country decides whether to commit to a single tax rate—equal to or above the GMT rate.
@@
 $\phi\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\int_{\underline{\pi}}^{\infty}\pi dF}\equiv\dfrac{\int_{\pi_{M}}^{\infty}\pi dF}{\Pi}.$ (8)
-Table 1. Possible Regimes of the Non-Cooperative Tax Game.
+Table 1
+Possible regimes of the non-cooperative tax game.
 Non-haven Haven
@@
 split tax rate Regime 4
-Table 1 shows possible combinations of tax policies of the non-haven and the havens.
+Table 1 shows possible combinations of tax policies in the non-haven and the havens.
 Of the nine potential tax regimes, four regimes cannot occur in equilibrium.
@@
 Regime 0 is the unconstrained tax-competition equilibrium, where the GMT rate $t_{M}$ is non-binding in all countries, as discussed earlier.
-As $t_{M}$ rises, the equilibrium moves through four additional regimes, characterised as follows.
-Proposition 1
-(Equilibrium regimes with binding GMT)
-Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$.
+As $t_{M}$ rises, the equilibrium moves through four additional regimes, which are characterised as follows.
+Proposition 1.
+(Equilibrium regimes with binding GMT)  Consider a GMT rate $t_{M}\geq t_{M}^{0}\equiv t_{h}^{0}$, a GMT coverage rate $\phi\in(0,1)$, a marginal value of public funds $\lambda>1$, and a cost parameter of profit shifting $\delta\in(0,3H/2)$.
 As $t_{M}$ is continuously increased, the tax-competition equilibrium is characterised by the following four regimes, with all symmetric havens choosing the same tax schedule.
@@
     $\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$
-Proof: See Appendix 1.
+Proof: See B.1.
 In the following we go through the different regimes, which are divided by the three threshold GMT rates $\{t_{M}^{i}\}_{i=1,2,3}$. We call these the regime-switching rates.
@@
 This argument is strongest when the havens decide collectively, as assumed in our main model.
-[^16] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Supplemental Appendix B.
+[^17] Qualitatively, the same results hold when the havens make their first-stage decisions in a decentralised manner, as shown in Section 5.2.
 In this case, the parameter range of Regime 1 is smaller, since the commitment of a single haven induces the non-haven to raise its tax rate less than the simultaneous commitment of multiple havens.
@@
 Notes: The figure shows the equilibrium tax rates in the non-haven in panel (a) and the haven in (b) for different GMT rates $t_{M}$, and the summary of equilibrium regimes in (c).
-While the levels of tax rates and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.
-Figure 1 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$.
+While the levels of tax rates and regime-switching rates vary with parameter values, the qualitative patterns remain unchanged.
+Fig.
+2 shows the equilibrium tax rates in the different regime-specific equilibria for varying values of the GMT rate $t_{M}$.
 The non-haven’s tax rates are given in panel (a), while the havens’ tax rates are in panel (b).
@@
 The havens respond to that by also discontinuously reducing their tax rates on small MNEs to the Regime 0 level.
-2.2 Welfare effects of the GMT
-We briefly summarise how the GMT introduction affects national and global welfare.
-Introducing a GMT rate $t_{M}$ slightly above the havens’ unconstrained one $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the havens.
-At the same time, profit shifting is reduced in comparison to the unconstrained Regime 0 due to a smaller tax differential.
-The sum of these effects must benefit the non-haven.
-It also benefits the havens, as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
-Once $t_{M}$ exceeds a certain threshold within Regime 1, the relationship reverses for the havens, as the loss of shifted profits dominates.
-Introducing any GMT rate above this threshold benefits the non-haven but harms the havens, with the net effect raising global welfare.
-Supplemental Appendix C provides a formal proposition and proof of these results.
-We also study the effects of marginal increases in the GMT tax rate $t_{M}$ and the GMT coverage rate $\phi$.
-An important result is that welfare in the non-haven and the world increases with $t_{M}$ in general, and the welfare gains become even larger as $\phi$ increases.
-A higher $\phi$ reduces the effectiveness of the split tax system and limits tax competition for small MNEs, thereby amplifying the welfare gains from raising $t_{M}$.
-See Supplemental Appendices D and E for formal discussions and Appendix F for graphical illustrations.
+4 Welfare effects of the GMT
+Having fully described the tax-competition equilibrium in each regime, we can now turn to the welfare effects of introducing the GMT.
+As in Proposition 1, the qualitative nature of the propositions below holds independently of the marginal valuation of public funds, $\lambda$, and it includes revenue maximization ($\lambda\rightarrow\infty$) as a special case.
+Our discussion of the results will therefore focus on the implications that the GMT has on tax revenues.
+We compare welfare levels in the tax-competition equilibrium to those in the unconstrained equilibrium, Regime 0. The results are summarized in:
+Proposition 2.
+(GMT introduction)  Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1, and assume the GMT coverage rate is not too low so that $\phi\in(3/4,1)$.
+Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
+  • (i)
+    a rise in welfare in the non-haven country for all Regimes 1 to 4.
+  • (ii)
+    a rise in welfare in the haven country for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
+  • (iii)
+    a rise in world welfare for all Regimes 1 to 4.
+Proof: See B.2.
+Introducing a small GMT slightly higher than the havens’ unconstrained rate $t_{h}^{0}$ leads to an equilibrium in Regime 1, with higher tax rates in both the non-haven and the haven countries.
+At the same time, profit shifting is reduced in comparison to the no-GMT benchmark due to a smaller tax differential.
+The sum of these effects must certainly benefit the non-haven.
+It also benefits the havens, as long as the gain from higher rates exceeds the loss of the havens’ tax base, i.e., shifted profits.
+However, once the GMT rate exceeds a certain threshold, $t_{M}^{+}$, which still lies in Regime 1, the reduced tax differential and thus the fall in shifted profits will dominate from the perspective of the haven countries.
+Therefore, for all levels of the GMT above $t_{M}^{+}$, tax havens lose from the introduction of the GMT while the non-haven gains.
+[^20] This conflict of interest does not arise in (Johannesen, 2022, Proposition 6), where havens gain from the introduction of a GMT at any level, as long as profit shifting is not completely eliminated.
+In his setting where MNEs shift profits only to the lowest-tax country, all the havens end up setting their tax rate to zero.
+Therefore, revenues in the unconstrained tax competition equilibrium are zero for each haven, and any positive GMT rate generates a revenue increase for the representative haven.
+See also footnote 16.
+In the remaining Regimes 2 to 4, the tax differential for large MNEs is always smaller than the one in Regime 0.
+As long as large MNEs covered by the GMT account for a sufficiently high share of aggregate profits, $\phi\in(3/4,1)$, the havens lose from the introduction of $t_{M}$ higher than $t_{M}^{+}$ due to the reduced profit shifting of large MNEs.
+Hence, the effects of introducing the GMT on tax revenues and welfare in the non-haven and the haven countries remain opposed to each other throughout Regimes 2 to 4.212121If small MNEs not covered by the GMT account for a sufficiently large share of aggregate profits so that $\phi\in(0,3/4)$, then the introduction of the GMT rate would instead benefit the havens in Regime 3.
+See B.2 for details.
+Finally, the gains to the non-haven exceed the losses (if any) to the havens, so that the GMT introduction raises tax revenues and welfare worldwide in all regimes.
+These aggregate welfare gains result from the higher tax rates in both countries and the reduced levels of profit shifting.
+Fig.
+C.1 in C illustrates the welfare effects summarised in Proposition 2.
+After the GMT has been introduced, two further reform options exist: (i) a marginal increase in the GMT rate, $t_{M}$, and (ii) a marginal increase in the GMT coverage rate, $\phi$.
+An increase in $t_{M}$ generally increases welfare in the non-haven and the world, except at the regime-switching rates where either of the non-haven or the havens introduce a split tax system.
+For the non-haven country, the welfare effects of marginal increases in the GMT rate are non-monotonous.
+A higher $\phi$ has non-negative welfare effects for the non-haven and the world in all regimes, and it has non-positive welfare effects for the havens.
+Intuitively, a rise in $\phi$ limits tax competition for small MNEs, and therefore amplifies the welfare effects arising from the GMT.
+We fully study these reforms in D.1 and D.2 respectively.
 5 Quantitative implications
-We now explore the quantitative importance of our theoretical results.
-We modify the model of the previous section by assuming that the havens decide tax policies in a decentralised manner (see footnote 16 and Supplemental Appendix B).
-The decentralised decision making leads to competition between the havens and thus makes it more likely that they split their tax rates.
-[^19] The new regime switching rate dividing Regimes 1 and 2 is smaller ($t_{M}^{b}<t_{M}^{1}$), and thus Regime 1 is less likely to emerge, as $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$.
-This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens, making gains from commitment smaller.
+5.1 Basic model
 We calibrate the unconstrained model with no GMT (Regime 0) to match international profit-shifting data, as summarised in Table 2.
-The GMT coverage rate is set to $\phi=0.9$, following OECD, (2020) and our calculations from Orbis.
-[^20] See Supplemental Appendix A for details.
-The number of tax havens is set to $H=40$, as in Tørsløv et al., (2023).
-Combining their data with Country-by-Country Reporting (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD.
-The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the model-predicted non-haven tax rate and shifted profit share and their empirical counterparts.
+The GMT coverage rate is set to $\phi=0.9$, following OECD (2020) and our calculations in Table 1.
+The number of tax havens is set to $H=40$, as in Tørsløv et al.
+(2023).
+Combining their data with Country-by-Country Reports (CbCR) yields total MNE pre-tax profits of $\Pi=4,623$ billion USD.
+The marginal value of public funds (MVPF), $\lambda$, and the cost parameter of profit shifting, $\delta$, are jointly calibrated to minimise the squared distance between the two model-predicted moments— the non-haven tax rate $t_{n}^{0}$ and the shifted profit share $\sum_{h}\theta_{h}^{0}$—and their empirical counterparts.
 For non-targeted moments, our model slightly underestimates the havens’ tax rate, $t_{h}^{0}$, and the non-haven’s revenue losses, $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$, but overall fits the data reasonably well.
-Table 2. Calibration of the Model.
+Table 2
+Calibration of the model.
 Parameter Source
-$\phi=0.9$ Profit share of MNEs covered by GMT OECD, (2020), Orbis
-$H=40$ Number of tax havens Tørsløv et al., (2023)
-$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al., (2023), CbCR
-$\lambda=2.12$ Marginal value of public funds Calibrated
-$\delta=17.80$ Cost of profit shifting Calibrated
+$\phi=0.9$ Profit share of MNEs covered by GMT OECD (2020), Orbis
+$H=40$ Number of tax havens Tørsløv et al. (2023)
+$\Pi=4,623$bUSD Pre-tax profits of MNEs Tørsløv et al. (2023), CbCR
+$\lambda=2.1$ Marginal value of public funds Calibrated
+$\delta=17.8$ Cost of profit shifting Calibrated
 Targeted moment Model Data Source
-Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al., (2023)
-Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al., (2023), CbCR
+Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al. (2023)
+Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al. (2023), CbCR
 Non-targeted moment Model Data Source
-Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al., (2023)
-Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al., (2023)
+Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al. (2023)
+Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al. (2023)
 Notes: Data are from 2019.
-The definitions of non-haven and haven countries follow Tørsløv et al., (2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’).
+The definitions of non-haven and haven countries follow Tørsløv et al.
+(2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables’).
 The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world.
 There are $40$ haven countries across the world.
-The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates.
-Tørsløv et al., (2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7).
+The corporate tax rates for the representative the non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates.
+Tørsløv et al.
+(2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7).
 Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs.
@@
 Using this calibrated version of our model, we examine the quantitative effects of GMT reforms, as summarised in Table 3.
-The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields the equilibrium in Regime 2 (Table 3(a)): the havens split their tax rates and set a lower tax rate ($10.1\%$) for small MNEs as compared to large MNEs ($15.0\%$) (Proposition 1(ii)).
-[^21] Above the regime switching rate of $t_{M}^{b}=0.094$, Regime 2 prevails.
-This result aligns with observations that low-tax countries like Ireland plan to retain their base rate (12.5% in Ireland’s case) while applying top-up taxes on large MNEs under the GMT (QDMTT).
-Table 3. Quantitative Effects of GMT Reforms.
-(a) Gains from a 15% GMT under $\phi=0.9$ Non-haven Haven World
+The key result is that a 15% GMT rate with partial coverage ($\phi=0.9$) yields an equilibrium in Regime 1 (Table 3(a)), whereas a 16% GMT rate leads to an equilibrium in Regime 2 (Table 3(b)).
+At $t_{M}=0.16$, the havens split their tax rates and set a lower tax rate ($10.3\%$) for small MNEs as compared to large MNEs (16.0%) (Proposition 1(ii)).
+Indeed, the threshold GMT rate is $t_{M}^{1}=0.156$, above which Regime 2 prevails.
+As we will see below, the threshold level for the GMT to induce an equilibrium in Regime 2 is further reduced, if the havens set their tax rates in a decentralized way.
+Table 3
+Quantitative effects of GMT reforms.
+(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World
+Tax rate after GMT introduction $20.5\%$ for all $15.0\%$ for all —–
+Welfare change in bUSD (% change) $95.6$ ($3.8\%$) 3.6-3.6 ($4.1\%$) $92.0$ ($3.6\%$)
+Revenue change in bUSD (% change) $153.0$ ($22.6\%$) 3.6-3.6 ($4.1\%$) $149.3$ ($19.5\%$)
+Profit change in bUSD (% change) $393.0$ $(41.0\%)$ 393.0-393.0 $(-41.0\%)$ 0 $(0\%)$
+(b) Gains from a 16% GMT with $\phi=0.9$ Non-haven Haven World
+Tax rate after GMT introduction $20.6\%$ for all $\begin{cases}10.3\%&\text{for small}\\ 16.0\%&\text{for large}\end{cases}$ —–
+Welfare change in bUSD (% change) $104.4$ ($4.2\%$) 8.8-8.8 ($9.9\%$) $95.6$ ($3.7\%$)
+Revenue change in bUSD (% change) $165.0$ ($24.4\%$) 8.8-8.8 ($9.9\%$) $156.2$ ($20.3\%$)
+Profit change in bUSD (% change) $422.3$ $(44.0\%)$ 422.3-422.3 $(-44.0\%)$ 0 $(0\%)$
+[…]
+Tax rate after GMT introduction $20.8\%$ for all $16.0\%$ for all —–
+Welfare change in bUSD (% change) $114.5$ ($4.6\%$) 8.9-8.9 ($10.0\%$) $105.6$ ($4.1\%$)
+Revenue change in bUSD (% change) $181.1$ ($26.8\%$) 8.9-8.9 ($10.0\%$) $172.2$ ($22.5\%$)
+Profit change in bUSD (% change) $461.2$ $(48.1\%)$ 461.2-461.2 $(-48.1\%)$ 0 $(0\%)$
+Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$); panel (b), from $t_{M}=0$ to $0.16$ with $\phi=0.9$; and panel (c), from $t_{M}=0$ to $0.16$ with full coverage ($\phi=1.0$).
+The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven.
+The regime-switching rate between Regimes 1 and 2 is $t_{M}^{1}=15.6\%$.
+Results are based on the calibrated parameter values described in Table 2.
+The welfare levels are normalised by the MVPF, $\lambda$.
+The ‘haven’ here refers to the sum of the $H=40$ haven countries.
+For both $t_{M}=0.15$ and $0.16$, the non-haven’s revenues rise due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition.
+The calibrated revenue gain in the non-haven (153–165 billion USD) is close to the combined gains for Europe and the U.S.
+in Baraké et al.
+(2022, Table 1) ($67+58=125$ billion EUR) and worldwide gains (149–156 billion USD) are in the same range as in Hugger et al.
+(2024, Section 8.2) (155–192 billion USD).
+Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues.
+[^22] In the havens without MNE headquarters, revenues coincide with welfare.
+Do the quantitative effects of introducing GMT change if the coverage is extended?
+The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) are unchanged since both the non-haven and the havens set a single tax rate for all MNEs, implying the same outcomes as in Table 3(a).
+By contrast, introducing $t_{M}=0.16$ with full coverage yields the non-splitting equilibrium in Regime 1 (Table 3(c)), rather than the splitting one in Regime 2 (Table 3(b)).
+By eliminating competition for small MNEs, the extended coverage increases revenue by about 16 billion USD and welfare by about 10 billion USD, both in the non-haven and worldwide.
+[^23] Comparing Tables 3(b) and 3(c), extending $\phi$ from 0.9 to 1.0 increases the non-haven’s revenues by 181.1165.0=16.1181.1-165.0=16.1 billion USD and their welfare by 114.5104.4=10.1114.5-104.4=10.1 billion USD.
+Corresponding numbers for the world are 172.2156.2=16.0172.2-156.2=16.0 billion USD and 105.695.6=10.0105.6-95.6=10.0 billion USD, respectively.
+These gains should be compared with the estimated compliance costs for small MNEs, however.
+Applying the results of Gaul et al.
+(2022) to our setting, the additional compliance costs are in the range of 10.5 billion USD.
+[^24] Using a survey of German MNEs (see Becker et al., 2012 for details on their characteristics), Gaul et al.
+(2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR.
+Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in A), the total compliance costs of newly covered MNEs are 1.6 million EUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 billion USD.
+Subtracting this amount from the gains of a full-coverage GMT results in net revenue gains of 5.5 billion USD, but a small net welfare loss of –0.5 billion USD globally.
+[^25] Clifford et al.
+(2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
+5.2 Extensions
+In this section we provide two extensions of our basic model and confirm the robustness of our quantitative findings.
+5.2.1 Decentralised decision-making by tax havens
+We modify the basic model by allowing the havens to decide tax policies in a decentralised manner.
+The decentralised decision-making leads to competition between the havens and thus makes it more likely that they split their tax rates.
+[^26] The new regime-switching rate dividing Regimes 1 and 2, denoted by $t_{M}^{b}$, is smaller than that under collective decision-making, i.e., $t_{M}^{b}<t_{M}^{1}$.
+Consequently, the parameter range in which Regime 1 emerges is narrower: $t_{M}^{0},t_{M}^{b})\subset[t_{M}^{0},t_{M}^{1})$.
+This is because the commitment to a single GMT rate by a single haven does not induce the non-haven to raise its tax rate as much as the collective commitment by multiple havens.
+This reduces the gains from committing to a single GMT rate for each haven, and increases the range of GMT rates that lead to a split tax equilibrium.
+A formal analysis of decentralized decision-making by tax havens is provided in E.1.
+Using the calibrated parameter values in Table 2, we reassess the quantitative effects of the GMT reforms in Table 4.272727We can use the same calibrated parameter values as in the collective decision-making case because the analytical expressions for the targeted calibration moments, i.e., the non-haven’s rate $t_{n}^{0}$ and the share of shifted profits $\sum_{h}\theta_{h}^{0}$ in Regime 0, remain unchanged.
+Table 4(a) reports the results for a 15% GMT rate with partial coverage ($\phi=0.9$).
+In contrast to the collective decision-making case in Table 3(a), this reform yields an equilibrium in Regime 2: havens split their tax rates, setting a lower rate for small MNEs (10.1%) than for large MNEs (15.0%) (see Proposition 1(ii)).
+This result — a GMT rate of 15% leading the havens to adopt a split tax system — aligns with observations that low-tax countries like Ireland have announced to keep their base rate unchanged (in Ireland’s case at 12.5%) while applying additional taxes on large MNEs to comply with the GMT.
+As a result of this response by the havens, the revenue and welfare gains from introducing a 15% GMT rate are smaller than under collective decision-making (Table 4(a) vs.
+Table 3(a)).
+Increasing the coverage rate from 90% to 100% at the 15% GMT rate eliminates the split tax system, as shown in Table 4(b), and brings revenue and welfare gains for both the non-haven and the world.
+[^28] Comparing Tables 4(a) and 4(b), extending $\phi$ from 0.9 to 1.0 increases the non-haven’s revenues by 153.0139.5=13.5153.0-139.5=13.5 billion USD and their welfare by 95.687.2=8.495.6-87.2=8.4 billion USD.
+Corresponding revenue and welfare gains for the world are 149.3135.7=13.6149.3-135.7=13.6 billion USD and 92.083.4=8.692.0-83.4=8.6 billion USD, respectively.
+These numbers are comparable to those under collective decision-making (see the last paragraph of Section 5.1).
+Table 4
+GMT reforms under decentralised decision-making by tax havens.
+(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World
 Tax rate after GMT introduction $20.3\%$ for all $\begin{cases}10.1\%&\text{for small}\\ 15.0\%&\text{for large}\end{cases}$ —–
@@
 Profit change in bUSD (% change) $359.9$ $(37.5\%)$ 359.9-359.9 $(-37.5\%)$ 0 $(0\%)$
-(b) Gains from a 15% GMT under $\phi=1.0$ Non-haven Haven World
+(b) Gains from a 15% GMT with $\phi=1.0$ Non-haven Haven World
 Tax rate after GMT introduction $20.6\%$ for all $15\%$ for all —–
@@
 The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven.
+The regime-switching rate between Regimes 1 and 2 is $t_{M}^{b}=9.3\%$ (see also footnote 26).
 Results are based on the calibrated parameter values described in Table 2.
@@
 The ‘haven’ here refers to the sum of the $H=40$ haven countries.
-The non-haven’s revenue rises due to weaker tax competition (higher tax rates) and reduced profit shifting, while the havens’ revenues fall as the loss of shifted profits outweighs the gains from milder tax competition.
-The calibrated revenue gain in the non-haven (139.5 bUSD) is close to the combined gains for Europe and the US in Baraké et al., (2022, Table 1) ($67+58=125$ bEUR), and the global revenue gain (135.7 bUSD) is comparable to Hugger et al., (2024, Section 8.2) (155–192 bUSD).
-Since falling shifted profits reduce MNEs’ private benefits, the non-haven’s welfare rises less than their revenues.
-[^22] In the havens without MNE headquarters, revenues coincide with welfare.
-Do the quantitative effects of introducing GMT change if the coverage is extended?
-The effects of $t_{M}=0.15$ with full coverage ($\phi=1.0$) yields the non-splitting equilibrium in Regime 1 (Table 3(b)).
-By eliminating competition for small MNEs, the extended coverage increases revenue and welfare gains worldwide by 13.6 bUSD and 8.6 bUSD, respectively.
-Based on the estimates by Gaul et al., (2022), these costs would amount to about 10.5 bUSD for our sample, implying a small net revenue gain (+3.1+3.1 bUSD) but a net welfare loss (1.9-1.9 bUSD) globally.
-[^23] Using a survey of German MNEs (see Becker et al., , 2012 for details on their characteristics), Gaul et al., (2022) estimate the average one-time expenses of implementing GMT per firm at 1.6 million EUR.
-Applying this to the number of small MNEs not covered by GMT in our database ($\approx$6,000; see Table A.1 in Supplemental Appendix A), the total compliance costs of newly covered MNEs are 1.6 mEUR $\times$ 6,000 $\times$ 1.1 USD/EUR $\approx$ 10.5 bUSD.
-Clifford et al., (2025) also estimate size-dependent compliance costs for German MNEs and similarly find that an increased coverage of the GMT would provide only limited gains for Germany.
+5.2.2 Real responses of MNEs’ profits to taxes
+We extend the basic model to incorporate real responses of MNEs to taxes.
+To keep the model simple, we assume that the pre-tax profit of an MNE decreases with the non-haven’s tax rate according to $\pi=\pi^{b}/(1+t_{n})$, where the baseline pre-tax profit, $\pi^{b}\in[\underline{\pi},\infty)$, follows a cumulative distribution function.
+The total pre-tax profits are then
+$\Pi=\int_{\underline{\pi}}^{\infty}\frac{\pi^{b}}{1+t_{n}}dF=\frac{\Pi^{b}}{1+t_{n}},$
+where $\Pi^{b}\equiv\int_{\underline{\pi}}^{\infty}\pi^{b}dF$ are the total baseline pre-tax profits.
+This negative effect of taxes on pre-tax profits generally reduces optimal corporate tax rates in both countries, compared to the basic model with pre-tax profits independent of taxes.
+To reassess the quantitative effects of the GMT reforms, we recalibrate the parameter values because the non-haven’s tax rate in Regime 0, one of the targeted calibration moments, differs from that in the basic model.
+Employing the same calibration strategy as before, the new calibration yields an MVPF of $\lambda=6.0$ as shown in Table E.1 in E.2, compared to $\lambda=2.1$ in Table 2.
+When MNEs’ profits respond to taxes, the non-haven’s unconstrained tax rate is lower than in the absence of such responses.
+To match this lower rate with its data counterpart, the MVPF, which captures the relative importance of tax revenues relative to MNEs’ private benefits, must be higher.
+A higher MVPF raises equilibrium tax rates in both countries.
+In particular, the regime-switching rate between Regimes 1 and 2 is 15.9%, slightly higher than 15.6% in the basic model, above which the havens start splitting their tax rates.
+The quantitative magnitudes of the GMT reforms are reported in Table E.2.
+in E.2 and are similar to those in Table 3.
+For example, the introduction of a 15% GMT rate with partial coverage increases welfare in the non-haven and the world by 108.9 and 106.3 billion USD respectively (vs.
+95.6 and 92.0 billion USD in Table 3(a)), and reduces the havens’ welfare by 2.5 billion USD (vs.
+3.6 billion USD in Table 3(a)).
 6 Conclusion
-This paper has analysed the effects of a global minimum tax (GMT) confined to large multinational enterprises (MNEs), leaving about 10% of the global tax base outside its scope.
-This partial-coverage GMT generates a sequence of tax competition equilibria between a non-haven and tax havens.
-In particular, introducing a low GMT rate that binds the havens induces them to commit to the single GMT rate.
-However, a further increase in the GMT rate leads first the havens, and then the non-haven, to split their tax rates for large and small MNEs.
-The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, the most likely regime is one where the non-haven sets a uniform tax rate above the GMT for all MNEs, whereas the havens split their tax rates and undercut the GMT for small MNEs.
-This finding corresponds to the observation that tax havens such as Ireland and Switzerland maintain their base corporate tax rates below 15%, but top up taxes on large MNEs to the level of the GMT.
-In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the havens lose, resulting in a positive net gain globally.
+This paper has analysed the effects of a global minimum tax (GMT) that is confined to large multinational enterprises (MNEs), thus leaving at least 10% of the global multinational tax base outside its scope.
+Using a simple model with profit shifting by heterogeneous MNEs, we have shown that this partial coverage of the GMT gives rise to a sequence of tax competition equilibria between a non-haven and a set of symmetric tax haven countries.
+In particular, introducing a low GMT rate that still binds the tax havens induces them to commit to a single GMT rate for all MNEs and thus results in higher welfare and tax revenues in both sets of countries.
+However, a further increase in the GMT rate leads first the tax havens, and then the non-haven, to split their tax rates for large and small MNEs, creating a conflict of interest between the two groups of countries.
+The calibrated version of our model suggests that, upon the introduction of the current GMT rate of 15%, both the non-haven and the havens set a uniform tax rate (Regime 1), while introducing a 16% GMT rate leads the havens to split their tax rates and undercut the GMT for small MNEs (Regime 2).
+We have also checked, under alternative model specifications, that a GMT rate close to 15% results in a split tax system in the haven countries.
+Although it is still too early to draw definitive conclusions given the ongoing implementation of the GMT, this finding is consistent with the observation that low-tax countries such as Ireland, Hungary, and Switzerland maintain their statutory corporate tax rates below 15%, but top up the taxation of large MNEs to the level of the GMT.
+In terms of welfare and tax revenues, our calibrated model predicts that the non-haven gains while the tax havens lose, resulting in a positive net gain globally.
+Our main objective in this paper has been to understand the implications of an incomplete coverage of the GMT.
+We have done so in a highly stylized model, in order to develop sharp intuitions.
+With the insights gained, it would be fruitful to extend our basic model to incorporate further important aspects of the GMT and the responses of MNEs.
+One extension would incorporate a splitting response of MNEs, whereby they restructure into smaller entities to benefit from potentially lower tax rates on small MNEs.
+Another extension would consider the substance-based income exclusion, i.e., additional tax deductions for MNEs that increase investment and employment.
+We leave these extensions for further research.
 Appendix
-Appendix 1: Proof of Proposition 1
+Appendix A Data for Fig. 1
+We select from the Orbis database MNEs satisfying the following conditions:
+  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
+  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
+  • Non-missing revenues (‘Operating revenue (Turnover)’) and pre-tax profits (‘P/L before tax’).
+Column (a) of Table A.1 reports the pre-tax profits and the total number of MNEs in each year.
+Column (b) restricts the sample to MNEs subject to the GMT, i.e., those with annual revenues of at least 750 million EUR in two or more of the past four years.
+Using columns (a) and (b), column (c) shows the share of MNEs subject to the GMT in terms of the pre-tax profits and the number of MNEs.
+Table A.1
+MNEs in the Orbis database.
+(a) All MNEs (b) MNEs $\geq$ 750mEUR (c) Share == B/A
+Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
+2018 2320 8656 2066 2333 0.89 0.27
+2019 2430 8416 2216 2437 0.91 0.29
+2020 1985 7920 1793 2228 0.90 0.28
+2021 3564 8803 3239 2627 0.91 0.30
+Source: Orbis database, own calculations.
+Note: Pre-tax profits are in billion EUR.
+Appendix B Proofs of Propositions
+B.1 Proof of Proposition 1
 Regime 0 can be an equilibrium as long as the haven’s unconstrained rate is greater than the GMT rate, $t_{M}<t_{h}^{0}=t_{M}^{0}\equiv\delta(\lambda-1)/[H(3\lambda-1)]$.
@@
 $t_{n}^{1}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)},\qquad t_{h}^{1}=t_{M}.$ (A.1)
-$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\hskip 18.49988ptG_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
+$\begin{split}&G_{n}^{1}\equiv G_{n}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}),\qquad G_{h}^{1}\equiv G_{h}(t_{n}=t_{n}^{1},t_{h}=t_{h}^{1}).\end{split}$
 When the non-haven sets a single non-GMT rate and the haven splits its tax rate, the equilibrium tax rates are
 $t_{n}^{2}=\frac{2(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]},\qquad t_{h}^{2}=\begin{cases}\frac{(\lambda-1)(\phi Ht_{M}+\delta)}{H[\lambda(3+\phi)-(1+\phi)]}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.$ (A.2)
-$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\hskip 18.49988ptG_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
+$\begin{split}&G_{n}^{2}\equiv G_{n}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}),\qquad G_{h}^{2}\equiv G_{h}(t_{n}=t_{n}^{2},t_{h}=t_{h}^{2}).\end{split}$
 Given the non-haven’s single non-GMT rate, the haven prefers to set a single GMT rate, if
@@
 Regime 3: $t_{M}\in(t_{M}^{2},t_{M}^{3}]$, where $t_{M}^{2}\equiv\frac{2\delta(\lambda-1)}{H[3\lambda-1-\phi(\lambda-1)]}$ and $t_{M}^{3}=\frac{2\delta(\lambda-1)(8\lambda-3)}{H(3\lambda-1)(4\lambda-1)}$.
-For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximizing rate in response to the haven’s splitting rates.
+For $t_{M}>t_{M}^{2}$, the non-haven’s single-non GMT rate, given by (A.2), is bound by the GMT and it cannot choose the unconstrained maximising rate in response to the haven’s splitting rates.
 The non-haven thus chooses either a single GMT rate or a splitting tax schedule.
@@
 When the non-haven sets a single GMT rate and the haven splits its tax rate, the equilibrium tax rates and payoffs are
-$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\hskip 18.49988ptt_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
-$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\hskip 18.49988ptG_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
+$\begin{split}t_{n}^{3}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{3}=\begin{cases}t_{M}/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.4)
+$\begin{split}G_{n}^{3}\equiv G_{n}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}),\qquad G_{h}^{3}\equiv G_{h}(t_{n}=t_{n}^{3},t_{h}=t_{h}^{3}).\end{split}$
 When both the non-haven and the haven split their tax rates, the equilibrium tax rates are
-$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\hskip 18.49988ptt_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
-$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\hskip 18.49988ptG_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
+$\begin{split}t_{n}^{4}=\begin{cases}\frac{2\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\qquad t_{h}^{4}=\begin{cases}\frac{\delta(\lambda-1)}{H(3\lambda-1)}&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases}.\end{split}$ (A.5)
+$\begin{split}G_{n}^{4}\equiv G_{n}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}),\qquad G_{h}^{4}\equiv G_{h}(t_{n}=t_{n}^{4},t_{h}=t_{h}^{4}).\end{split}$
 Given the haven’s splitting rates, the non-haven prefers to set a single GMT rate, if
@@
 ∎
-Supplemental Appendix of
-‘A Global Minimum Tax for Large Firms Only:
-Implications for Tax Competition’
-Andreas Haufler [^1] Seminar for Economic Policy, LMU Munich, Akademiestr.
-1, 80799 Munich, Germany.
-Email: Andreas.Haufler@econ.lmu.de.
-Hayato Kato [^2] Graduate School of Economics, the University of Osaka, 1-7 Machikaneyama, Toyonaka, Osaka 560-0043, Japan.
-Email: hayato.kato@econ.osaka-u.ac.jp
-October, 2025
-Contents
-Appendix A MNEs Subject to the GMT
-We select from the Orbis database MNEs satisfying the following conditions:
-  • Global Ultimate Owner with foreign subsidiaries. The threshold ownership is 50.01%.
-  • C1: MNEs report only consolidated accounts, not unconsolidated accounts.
-  • Non-missing revenues (‘Operating revenue (Turnover)’) and pre-tax profits (‘P/L before tax’).
-Column (a) of Table A.1 reports the pre-tax profits and the total number of MNEs in each year.
-Column (b) restricts the sample to MNEs subject to the GMT, i.e., those with annual revenues of at least 750 million EUR in two or more of the past four years.
-Using columns (a) and (b), column (c) shows the share of MNEs subject to the GMT in terms of the pre-tax profits and the number of MNEs.
-Table A.1. MNEs in the Orbis Database.
-(a) All MNEs (b) MNEs $\geq$ 750mEUR (c) Share == B/A
-Year Pre-tax profits Number Pre-tax profits Number Pre-tax profits Number
-2018 2320 8656 2066 2333 0.89 0.27
-2019 2430 8416 2216 2437 0.91 0.29
-2020 1985 7920 1793 2228 0.90 0.28
-2021 3564 8803 3239 2627 0.91 0.30
-Source: Orbis database, own calculations.
-Note: Pre-tax profits are in billion EUR.
-Appendix B Decentralized Decision-Making by Tax Havens
-In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate.
-Here, we consider an alternative, decentralized decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others.
-In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions.
-First, countries can split their tax rates only if they are subject to the GMT.
-This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019).
-Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates.
-This allows us to focus on the problem of a single haven.
-Out of the total $H$ tax havens, $H_{M}$ havens commit to a single GMT rate, while the remaining $H-H_{M}$ havens split their tax rates.
-When the non-haven sets a single non-GMT rate and a haven $h$ sets a single GMT rate, their equilibrium tax rates are respectively
-$\begin{split}&t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\\ &t_{h}^{1}(H_{M})=t_{M},\end{split}$
-Let $G_{h}^{1}(H_{M})$ denote the equilibrium payoff of the haven $h$ with commitment.
-We note that as in the text, setting a single non-GMT rate is the non-haven’s dominant strategy because if $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, it is never constrained by the GMT: $t_{n}(H_{M})\leq t_{n}(H)=t_{n}^{2}\leq t_{M}$.
-If instead the haven $h$ splits its tax rate while the other havens and the non-haven keep their rates unchanged, the number of havens that maintain commitment falls to $H_{M}-1$.
-The equilibrium tax rates of the non-haven and the haven $h$ are respectively
-$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (B.1)
-Let $G_{h}^{2}(H_{M}-1)$ be the equilibrium payoff of the haven $h$ under splitting.
-The difference between the haven $h$’s payoffs in the two cases is
-$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
-$\ \ =\frac{(1-\phi)\Pi[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta(H_{M})}{\delta\left[\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})\right]\left[(\lambda-1)\left\{\phi(H-H_{M}+1)+H+H_{M}-1\right\}+2H\lambda\right]^{2}},$
-where
-$\Theta(H_{M})\equiv AH_{M}+B,$
-$A\equiv(\lambda-1)(1-\phi)\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right],$
-$B\equiv H[\phi(\lambda-1)+3\lambda-1]\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right]-2t_{M}(\lambda-1)^{2}(1-\phi).$
-If $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$, the haven $h$ always sets a single GMT rate, regardless of how many other havens it believes set a single GMT rate.
-Under our assumption of symmetric belief formation, all havens end up setting a single GMT rate, and the equilibrium coincides with Regime 1 in the text.
-Conversely, if $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$, all havens end up splitting their tax rates, and the equilibrium coincides with Regime 2 in the text.
-Since the denominator of $\Delta G_{h}^{12}(H_{M})$ is positive and $Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0$ for $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, the sign of $\Delta G_{h}^{12}(H_{M})$ is determined by that of $\Theta(H_{M})$.
-Considering that $\Theta(H_{M})$ is linear in $H_{M}$, the conditions for $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$ reduce to the following two conditions:
-$\Theta(1)=A\cdot 1+B=C-Dt_{M}\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{a}\equiv C/D,$
-$\Theta(H)=AH+B=2(E-Ft_{M})\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{b}\equiv E/F,$
-where
-$C\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
-$D\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
-$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
-$F\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
-We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
-$t_{M}^{a}-t_{M}^{0}=\frac{2\delta(\lambda-1)^{2}[\phi(\lambda-1)+3\lambda-1]}{D(3\lambda-1)}>0,$
-$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
-$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}[\phi(H+1)(\lambda-1)+\lambda(3H-1)-(H-1)]}{F[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0.$
-We can also confirm $\Theta(H_{M})=AH_{M}+B$ increases with $H_{M}$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$, or equivalently $A>0$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$.
-Since $A>0$ if $t_{M}<t_{M}^{c}\equiv\delta(\lambda-1)/[H(3\lambda-1)-2(\lambda-1)]$, it suffices to show $t_{M}^{b}<t_{M}^{c}$:
-$t_{M}^{c}-t_{M}^{b}=\frac{\delta(\lambda-1)[\lambda^{2}+(H-1)(2\lambda-1)-\phi(\lambda-1)^{2}]}{F[H(3\lambda-2)-(H-1)]}>0,$
-which is indeed positive.
-Using these results, we can thus conclude that (i) if $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$; (ii) if $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$; and (iii) if $t_{M}\in(t_{M}^{a},t_{M}^{b}]$, there exists a positive integer $H_{M}^{*}\in\{2,\dots,H\}$ such that $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,\dots,H_{M}^{*}-1\}$ and $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{H_{M}^{*},\dots,H\}$.
-In cases (i) and (ii), the equilibrium is unique.
-In case (iii), there can be at most three equilibria.
-First, if each haven believes that fewer than $H_{M}^{*}$ havens set a single GMT rate, all havens end up splitting their tax rates, i.e., $H_{M}=0$.
-Second, if each haven believes that more than $H_{M}^{*}-1$ havens set a single GMT rate, all havens end up doing so, i.e., $H_{M}=H$.
-Third, if $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, then $H_{M}=H_{M}^{*}$ havens set a single GMT rate while the remaining $H-H_{M}^{*}$ havens split their tax rates, consistent with their beliefs.
-The discussion is summarized as follows.
-If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)).
-If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)).
-If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge.
-If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (B.1)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (B.1)).
-A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralized manner, implying that they are less likely to commit to a single GMT rate.
-This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
-Appendix C Welfare Effects of the Introduction of GMT
-The following proposition characterizes the welfare effects of introducing a binding GMT rate, compared with the unconstrained equilibrium (Regime 0).
-Proposition C.1
-(GMT introduction)
-Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text, and assume the GMT coverage rate is sufficiently high, such that $\phi\in(3/4,1)$.
-Compared to the unconstrained equilibrium (Regime 0), the introduction of a binding GMT rate leads to:
-  • (i)
-    a rise in welfare in the non-haven for all Regimes 1 to 4.
-  • (ii)
-    a rise in welfare in the (representative) haven for the first part of Regime 1 ($t_{M}<t_{M}^{+}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H\lambda(3\lambda-1)}$), but a fall in welfare in the rest of Regime 1 and in Regimes 2 to 4 ($t_{M}>t_{M}^{+}$).
-  • (iii)
-    a rise in world welfare for all Regimes 1 to 4.
-These results hold for any value of the marginal value of public funds (MVPF), $\lambda\in(1,\infty)$.
-In the limit as $\lambda\to\infty$, countries care only about tax revenues.
-Hence, qualitatively identical results apply to tax revenues.
-Proof: In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
+B.2 Proof of Proposition 2
+In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
 Regime 1: $t_{M}\in(t_{M}^{0},t_{M}^{1}]$, where $t_{M}^{0}=\frac{\delta(\lambda-1)}{H(3\lambda-1)}$ and $t_{M}^{1}=\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$.
@@
 ∎
-Appendix D Welfare Effects of Increasing the GMT Rate
+Appendix C Equilibrium welfare and shifted profits
+Fig. C.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
+The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those with full GMT coverage (dashed curves).
+The equilibrium tax rates under full GMT are as follows.
+Regime 0: $t_{M}\in\left[0,t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)}\right)$.  Each of the non-haven and the haven sets its own single non-GMT rate:
+$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Regime 1f: $t_{M}\in\left[t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)},t_{M}^{f}\equiv\frac{\delta(\lambda-1)}{H\lambda}\right]$.
+The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
+$t_{n}^{1f}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Regime 2f: $t_{M}\in(t_{M}^{f},1]$.  Both the non-haven and the haven choose a single GMT rate:
+$t_{n}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
+Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the havens under full GMT.
+In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
+We note $t_{M}^{f}>t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$, given $\phi\in(0,1)$ and $\lambda>1$.
+That is, the equilibrium outcomes under partial GMT begins to diverge from those under full GMT once $t_{M}$ enters the range of Regime 1.
+This is particularly evident in panels (a) and (c).
+Fig. C.1. Equilibrium welfare and shifted profits.
+Notes: The figure shows the equilibrium levels of welfare normalised by the marginal value of public funds (MVPF) $\lambda$ in the non-haven ($G_{n}/\lambda$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}/\lambda$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$.
+The vertical axis is in billion USD.
+The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0).
+Results are based on the calibrated parameter values described in Table 2.
+Appendix D Welfare effects of gradual GMT reforms
+D.1 Welfare effects of increasing the GMT rate
 The following proposition characterises the welfare effects of a marginal increase in the GMT rate $t_{M}$.
-Proposition D.1
-(Increase in the GMT rate $t_{M}$)
+Proposition D.1. (Increase in the GMT rate $t_{M}$)
 Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text. A marginal increase in the GMT rate has the following effects:
   • (i)
-    welfare in the non-haven increases in all regimes, except at the regime-switching point $t_{M}^{1}$, where it falls discontinuously.
+    welfare in the non-haven increases in all regimes, except at the regime-switching rate $t_{M}^{1}$, where it falls discontinuously.
   • (ii)
     welfare in the (representative) haven increases in Regime 1 as long as $t_{M}<t_{M}^{++}\equiv\frac{\delta(\lambda-1)}{2H\lambda}$, then decreases thereafter in Regime 1 and in Regime 2.
-It increases in Regime 3, falls discontinuously at the regime-switching point $t_{M}^{3}$, and remains unchanged in Regime 4.
+It increases in Regime 3, falls discontinuously at the regime-switching rate $t_{M}^{3}$, and remains unchanged in Regime 4.
   • (iii)
-    welfare in the world increases in all regimes, except at the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
-As is true for the introduction of the GMT (Proposition C.1 in Section C of this Supplemental Appendix), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium.
-An exception is the regime-switching point $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs.
+    welfare in the world increases in all regimes, except at the two regime-switching rates $t_{M}^{1}$ and $t_{M}^{3}$, where it falls discontinuously.
+As is true for the introduction of the GMT (Proposition 1), the non-haven generally gains from a gradual increase in the GMT rate $t_{M}$, as this reduces profit-shifting to the havens in equilibrium.
+An exception is the regime-switching rate $t_{M}^{1}$ where the havens being to split their tax rates and set a discretely lower tax rate for small MNEs.
 Small MNEs then shift more profits to the havens, hurting the non-haven.
-For havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven.
+For the havens, the welfare (i.e., tax revenue) effects of an increase in the GMT rate reflect two opposing forces: a negative effect from reduced profit-shifting opportunities and a positive effect from a higher tax rate in the non-haven.
 These opposing forces generate a non-monotonic welfare response.
@@
 This regime, where commitment induces effective tax coordination, is similar to the first part of Regime 1 ($t_{M}\in[t_{M}^{0},t_{M}^{++})$).
-It ends at the regime-switching point $t_{M}^{3}$ at which the non-haven starts splitting its tax rate.
-This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the tax havens.
+It ends at the regime-switching rate $t_{M}^{3}$ at which the non-haven starts splitting its tax rate.
+This leads to a discontinuous fall in the non-haven’s tax rate on small MNEs, which must hurt the havens.
 In Regime 4 ($t_{M}\in(t_{M}^{3},1]$), both the non-haven and the havens apply the same GMT rate to large MNEs, preventing them from shifting profits.
@@
 The marginal effects of $t_{M}$ on world welfare are dominated by the effects on the non-haven, and are therefore generally positive.
-At the two regime-switching points $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs.
+At the two regime-switching rates $t_{M}^{1}$ and $t_{M}^{3}$, however, one of the countries begins to split its tax rate and discontinuously lowers the tax rate on small MNEs.
 Therefore, tax competition is tightened and world welfare falls at the margin.
 In the following proof, since the haven countries are all symmetric, we are concerned with a representative haven $h$.
-Proof:  In the proof of Proposition C.1 in Section C, we already showed the welfare effects of a marginal increase in the GMT rate within each regime.
-Here we show the welfare effects at the regime switching points.
+Proof:  In B.2, we already showed the welfare effects of a marginal increase in the GMT rate within each regime. Here we show the welfare effects at the regime switching rates.
 At the three regime switching GMT rates, $t_{M}^{1}$, $t_{M}^{2}$ and $t_{M}^{3}$, welfare in the non-haven, the haven, and the world changes as follows.
@@
 $G_{h}^{4}-G_{h}^{3}\big|_{t_{M}=t_{M}^{3}}=G_{W}^{4}-G_{W}^{3}\big|_{t_{M}=t_{M}^{3}}=-\frac{8\delta\lambda(\lambda-1)^{2}(2\lambda-1)}{2H(\lambda-1)[H(3\lambda-1)(4\lambda-1)^{2}]}(1-\phi)\Pi<0.$
-This completes the proof of Proposition D.1. ∎
-Appendix E Welfare Effects of Increasing the GMT Coverage Rate
+This completes the proof of Proposition C.1. ∎
+D.2 Welfare effects of increasing the GMT coverage rate
 The following proposition characterizes the welfare effects of a marginal increase in the GMT coverage rate $\phi$.
-Proposition E.1
-(Increasing in the GMT coverage rate $\phi$)
+Proposition D.2. (Increasing in the GMT coverage rate $\phi$)
 Consider the tax-competition equilibrium with GMT, as summarized in Proposition 1 in the text.
@@
 This latter regime switch, however, leaves welfare in all countries and the world unaffected.
-In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition 3).
+In summary, the welfare effects of an increase in the GMT coverage rate $\phi$ are similar to those of an increase in the GMT rate $t_{M}$ (Proposition C.1).
 A higher $\phi$ has opposing welfare effects in non-haven and haven countries in Regimes 2 to 4, but not in Regime 1.
@@
 noting that $t_{M}>t_{M}^{3}>\frac{\delta(\lambda-1)(10\lambda-3)}{2H(3\lambda-1)^{2}}>\frac{\delta(\lambda-1)(8\lambda-3)}{2H(3\lambda-1)^{2}}$.
-This completes the proof of Proposition E.1. ∎
-Appendix F Supplementary Figures
-Figure F.1 shows the welfare levels of the non-haven (panel (a)) and the sum of the havens (panel (b)), and the shifted profits for different GMT rates $t_{M}$.
-The tax-competition equilibrium outcomes under partial GMT coverage (solid curves) are compared with those under full GMT coverage (dashed curves).
-The equilibrium tax rates under full GMT are as follows.
-Regime 0: $t_{M}\in\left[0,t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)}\right)$.  Each of the non-haven and the haven sets its own single non-GMT rate:
-$t_{n}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{0}=\frac{2\delta(\lambda-1)}{H(3\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
-Regime 1f: $t_{M}\in\left[t_{M}^{0}\equiv\frac{\delta(\lambda-1)}{H(3\lambda-1)},t_{M}^{f}\equiv\frac{\delta(\lambda-1)}{H\lambda}\right]$.
-The non-haven chooses a single non-GMT rate, and the (representative) haven chooses a single GMT rate:
-$t_{n}^{1f}=\frac{(\lambda-1)(Ht_{M}+\delta)}{H(2\lambda-1)}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{1f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
-Regime 2f: $t_{M}\in(t_{M}^{f},1]$.  Both the non-haven and the haven choose a single GMT rate:
-$t_{n}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty),\qquad t_{h}^{2f}=t_{M}\ \ \text{for}\ \pi\in[\underline{\pi},\infty).$
-Let $G_{n}^{f}$ and $G_{h}^{f}$ be the equilibrium payoffs of the non-haven and the haven under full GMT.
-In Regime 0 ($t_{M}\in[0,t_{M}^{0})$), the equilibrium tax rates are identical under full and partial GMT coverage, as neither country is constrained by the GMT.
-We note $t_{M}^{f}>t_{M}^{1}\equiv\frac{\delta(\lambda-1)(2\lambda-1)}{H[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}$, given $\phi\in(0,1)$ and $\lambda>1$.
-That is, the equilibrium outcomes under partial GMT begins to diverge from those under full GMT once $t_{M}$ enters the range of Regime 1.
-This is particularly evident in panels (a) and (c).
-Fig. F.1. Equilibrium Welfare and Shifted Profits.
-Notes: The figure shows the equilibrium levels of welfare in the non-haven ($G_{n}$) in panel (a) and the sum of the $H=40$ havens ($\sum_{h=1}^{H}G_{h}$) in (b), and the aggregate shifted profits ($\sum_{h=1}^{H}\theta_{h}\Pi$) in (c), for different GMT rates $t_{M}$.
-The solid curves in each panel are the equilibrium values under the partial GMT ($\phi\in(3/4,1)$); the dashed curves are those under the full GMT ($\phi=1$); horizontal lines are those under no GMT (Regime 0).
-While the levels of welfare, shifted profits and regime-switching points vary with parameter values, the qualitative patterns remain unchanged.
+This completes the proof of Proposition C.2. ∎
+Appendix E Extensions of the basic model
+E.1 Decentralised decision-making by tax havens
+In the text, we assumed that a group of tax-haven countries collectively decide whether to adopt a single GMT rate or to split their tax rate.
+Here, we consider an alternative, decentralised decision-making process in the first stage of our game, where each tax haven individually decides whether to split, given the decisions of the others.
+In the following, we focus on $t_{M}\in[t_{M}^{0},t_{M}^{2}]$ and impose two assumptions.
+First, countries can split their tax rates only if they are subject to the GMT.
+This assumption reflects the fact that preferential tax regimes for specific tax bases are not permitted under the OECD’s action plan to address base erosion and profit shifting (OECD, 2019).
+Second, each haven forms the same belief about how the other havens and the non-haven choose their tax rates.
+This allows us to focus on the problem of a single haven.
+Out of the total $H$ tax havens, $H_{M}$ havens commit to a single GMT rate, while the remaining $H-H_{M}$ havens split their tax rates.
+When the non-haven sets a single non-GMT rate and a haven $h$ sets a single GMT rate, their equilibrium tax rates are respectively
+$\begin{split}&t_{n}(H_{M})=\frac{2(\lambda-1)[\phi t_{M}(H-H_{M})+H_{M}t_{M}+\delta]}{\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})},\\ &t_{h}^{1}(H_{M})=t_{M},\end{split}$
+Let $G_{h}^{1}(H_{M})$ denote the equilibrium payoff of the haven $h$ with commitment.
+We note that as in the text, setting a single non-GMT rate is the non-haven’s dominant strategy because if $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, it is never constrained by the GMT: $t_{n}(H_{M})\leq t_{n}(H)=t_{n}^{2}\leq t_{M}$.
+If instead the haven $h$ splits its tax rate while the other havens and the non-haven keep their rates unchanged, the number of havens that maintain commitment falls to $H_{M}-1$.
+The equilibrium tax rates of the non-haven and the haven $h$ are respectively
+$\begin{split}&t_{n}(H_{M}-1)=\frac{2(\lambda-1)\left[t_{M}\left\{H\phi+(H_{M}-1)(1-\phi)\right\}+\delta\right]}{(\lambda-1)[\phi(H-H_{M}+1)+H+H_{M}-1]+2H\lambda},\\ &t_{h}^{2}(H_{M}-1)=\begin{cases}t_{n}(H_{M}-1)/2&\text{for}\ \pi\in[\underline{\pi},\pi_{M})\\ t_{M}&\text{for}\ \pi\in[\pi_{M},\infty)\end{cases},\end{split}$ (A.7)
+Let $G_{h}^{2}(H_{M}-1)$ be the equilibrium payoff of the haven $h$ under splitting.
+The difference between the haven $h$’s payoffs in the two cases is
+$\Delta G_{h}^{12}(H_{M})\equiv G_{h}^{1}(H_{M})-G_{h}^{2}(H_{M}-1)$
+$\ \ =\frac{(1-\phi)\Pi[Ht_{M}(3\lambda-1)-\delta(\lambda-1)]\Theta(H_{M})}{\delta\left[\phi(\lambda-1)(H-H_{M})+\lambda(3H+H_{M})-(H+H_{M})\right]\left[(\lambda-1)\left\{\phi(H-H_{M}+1)+H+H_{M}-1\right\}+2H\lambda\right]^{2}},$
+where
+$\Theta(H_{M})\equiv AH_{M}+B,$
+$A\equiv(\lambda-1)(1-\phi)\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right],$
+$B\equiv H[\phi(\lambda-1)+3\lambda-1]\left[\delta(\lambda-1)-t_{M}\left\{H(3\lambda-1)-2(\lambda-1)\right\}\right]-2t_{M}(\lambda-1)^{2}(1-\phi).$
+If $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$, the haven $h$ always sets a single GMT rate, regardless of how many other havens it believes set a single GMT rate.
+Under our assumption of symmetric belief formation, all havens end up setting a single GMT rate, and the equilibrium coincides with Regime 1 in the text.
+Conversely, if $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$, all havens end up splitting their tax rates, and the equilibrium coincides with Regime 2 in the text.
+Since the denominator of $\Delta G_{h}^{12}(H_{M})$ is positive and $Ht_{M}(3\lambda-1)-\delta(\lambda-1)\geq 0$ for $t_{M}\in[t_{M}^{0},t_{M}^{2}]$, the sign of $\Delta G_{h}^{12}(H_{M})$ is determined by that of $\Theta(H_{M})$.
+Considering that $\Theta(H_{M})$ is linear in $H_{M}$, the conditions for $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$ reduce to the following two conditions:
+$\Theta(1)=A\cdot 1+B=C-Dt_{M}\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{a}\equiv C/D,$
+$\Theta(H)=AH+B=2(E-Ft_{M})\geq 0$ $\Leftrightarrow\ t_{M}\leq t_{M}^{b}\equiv E/F,$
+where
+$C\equiv\delta(\lambda-1)[\phi(\lambda-1)(H-1)+\lambda(3H+1)-(H+1)]>0,$
+$D\equiv H\phi(\lambda-1)[\lambda(3H-5)-(H-3)]+H(3\lambda-1)[\lambda(3H-1)-(H-1)]>0,$
+$E\equiv\delta H(2\lambda-1)(\lambda-1)>0,$
+$F\equiv(6H^{2}-4H+1)\lambda^{2}-(5H^{2}-6H+2)\lambda+(H-1)^{2}-\phi(\lambda-1)^{2}>0.$
+We can check $t_{M}^{0}<t_{M}^{a}<t_{M}^{b}<t_{M}^{1}$:
+$t_{M}^{a}-t_{M}^{0}=\frac{2\delta(\lambda-1)^{2}[\phi(\lambda-1)+3\lambda-1]}{D(3\lambda-1)}>0,$
+$t_{M}^{b}-t_{M}^{a}=\frac{\delta(H-1)(1-\phi)^{2}(\lambda-1)^{4}}{DF}>0,$
+$t_{M}^{1}-t_{M}^{b}=\frac{\delta(H-1)(2\lambda-1)(\lambda-1)^{2}[\phi(H+1)(\lambda-1)+\lambda(3H-1)-(H-1)]}{F[\lambda(3\lambda-1)-\phi(\lambda-1)^{2}]}>0.$
+We can also confirm $\Theta(H_{M})=AH_{M}+B$ increases with $H_{M}$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$, or equivalently $A>0$ for $t_{M}\in[t_{M}^{0},t_{M}^{b}]$.
+Since $A>0$ if $t_{M}<t_{M}^{c}\equiv\delta(\lambda-1)/[H(3\lambda-1)-2(\lambda-1)]$, it suffices to show $t_{M}^{b}<t_{M}^{c}$:
+$t_{M}^{c}-t_{M}^{b}=\frac{\delta(\lambda-1)[\lambda^{2}+(H-1)(2\lambda-1)-\phi(\lambda-1)^{2}]}{F[H(3\lambda-2)-(H-1)]}>0,$
+which is indeed positive.
+Using these results, we can thus conclude that (i) if $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{1,2,\dots,H\}$; (ii) if $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,2,\dots,H\}$; and (iii) if $t_{M}\in(t_{M}^{a},t_{M}^{b}]$, there exists a positive integer $H_{M}^{*}\in\{2,\dots,H\}$ such that $\Delta G_{h}^{12}(H_{M})<0$ for $H_{M}\in\{1,\dots,H_{M}^{*}-1\}$ and $\Delta G_{h}^{12}(H_{M})\geq 0$ for $H_{M}\in\{H_{M}^{*},\dots,H\}$.
+In cases (i) and (ii), the equilibrium is unique.
+In case (iii), there can be at most three equilibria.
+First, if each haven believes that fewer than $H_{M}^{*}$ havens set a single GMT rate, all havens end up splitting their tax rates, i.e., $H_{M}=0$.
+Second, if each haven believes that more than $H_{M}^{*}-1$ havens set a single GMT rate, all havens end up doing so, i.e., $H_{M}=H$.
+Third, if $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, then $H_{M}=H_{M}^{*}$ havens set a single GMT rate while the remaining $H-H_{M}^{*}$ havens split their tax rates, consistent with their beliefs.
+The discussion is summarised as follows.
+If $t_{M}\in[t_{M}^{0},t_{M}^{a}]$, the non-haven chooses a single non-GMT rate, and each haven chooses a single GMT rate, as in Regime 1 in the text (Proposition 1(i)).
+If $t_{M}\in(t_{M}^{b},t_{M}^{2}]$, the non-haven sets a single non-GMT rate, and each haven splits its tax rate and chooses a GMT rate for large MNEs, but a lower rate than the GMT for small MNEs, as in Regime 2 in the text (Proposition 1(ii)).
+If $t_{M}\in[t_{M}^{a},t_{M}^{b})$, both Regimes 1 and 2 can be equilibria and either may emerge.
+If $t_{M}\in[t_{M}^{a},t_{M}^{b})$ and in addition $H_{M}$ coincides with $H_{M}^{*}$ at which $\Delta G_{h}^{12}(H_{M}=H_{M}^{*})=0$, the non-haven sets a single non-GMT rate ($t_{n}(H_{M}^{*})$ in (A.7)), and there are $H_{M}^{*}$ havens setting a single GMT rate $t_{M}$ and $H-H_{M}^{*}$ havens splitting their tax rates ($t_{h}^{2}(H_{M}^{*})$ in (A.7)).
+A clear difference from Proposition 1 is that the range where only Regime 1 occurs is smaller ($t_{M}^{0},t_{M}^{a})\subset[t_{M}^{0},t_{M}^{1})$) when havens make their first stage decision in a decentralised manner, implying that they are less likely to commit to a single GMT rate.
+This is intuitive because the commitment of a haven induces the non-haven to raise its tax rate less than the simultaneous commitment of many havens.
+E.2 Real responses of MNEs’ profits to taxes
+As described in the text, we allow the pre-tax profits of MNEs, $\pi$, to vary in taxes.
+Specifically, we assume $\pi=\pi^{b}/(1+t_{n})$, where the baseline pre-tax profit, $\pi^{b}\in[\underline{\pi},\infty)$, follows a cumulative distribution function $F(\pi^{b})$.
+The total pre-tax profits are
+$\Pi=\int_{\underline{\pi}}^{\infty}\frac{\pi^{b}}{1+t_{n}}dF=\frac{\Pi^{b}}{1+t_{n}},$
+where $\Pi^{b}\equiv\int_{\underline{\pi}}^{\infty}\pi^{b}dF$ is the total baseline pre-tax profits.
+We experimented with an alternative specification, $\pi=\pi^{b}(1-t_{n})$, and found that the quantitative implications of introducing a 15% GMT rate are similar to those here.
+We adopt the calibration strategy described in Section 5 of the text.
+Specifically, we calibrate the unconstrained model without GMT (Regime 0) to match key data on international profit shifting.
+We set the GMT coverage rate to $\phi=0.9$ and the number of haven countries to $H=40$, as in Section 5.
+We jointly choose the MVPF $\lambda$ and the profit-shifting cost parameter $\delta$ to minimize the sum of squared distances between the model-implied non-haven tax rate and shifted profits and their empirical counterparts, $(t_{n}^{0}-0.186)^{2}+(\sum_{h}\theta_{h}^{0}-0.209)^{2}$, where
+$t_{n}^{0}=\frac{2\left[\sqrt{H\left[(16\delta+9H)\lambda^{2}-(38\delta-6H)\lambda+12\delta+H\right]}-H(3\lambda-1)\right]}{H(8\lambda-3)},\qquad t_{h}^{0}=\frac{t_{n}^{0}}{2},$
+$\theta_{h}^{0}=(t_{n}^{0}-t_{h}^{0})/\delta.$
+The calibration results are given in Table E.1 and the quantitative effects of GMT reforms in Table E.2.
+The results are similar to those in Section 5.
+As discussed in the text, the only notable difference is the higher calibrated value of the MVPF, $\lambda=6.0$ (vs.
+$\lambda=2.1$ in Table 2).
+Otherwise, calibrated values and the magnitude of the quantitative findings are similar to those in the basic model.
+Table E.1
+Calibration of the model with MNEs’ pre-tax profits dependent on taxes.
+Parameter Source
+$\phi=0.9$ Profit share of MNEs covered by GMT OECD (2020), Orbis
+$H=40$ Number of tax havens Tørsløv et al. (2023)
+$\Pi^{b}=$5,480bUSD Baseline pre-tax profits of MNEs Calibrated
+$\lambda=6.0$ Marginal valuation of public funds Calibrated
+$\delta=17.7$ Cost of profit shifting Calibrated
+Targeted moment Model Data Source
+Pre-tax profits of MNEs: $\Pi=\Pi^{b}/(1+t_{n}^{0})$ 4,623bUSD 4,623bUSD Tørsløv et al. (2023), CbCR
+Corp. tax rate of non-haven: $t_{n}^{0}$ $\% 18.618.6\%$ Tørsløv et al. (2023)
+Share of shifted profits: $\sum_{h}\theta_{h}^{0}$ $\% 20.920.9\%$ Tørsløv et al. (2023), CbCR
+Non-targeted moment Model Data Source
+Corp. tax rate of haven: $t_{h}^{0}$ $9.3\%$ $13.7\%$ Tørsløv et al. (2023)
+Non-haven’s revenue loss: $t_{n}^{0}\sum_{h}\theta_{h}^{0}\Pi$ 180180bUSD 247247bUSD Tørsløv et al. (2023)
+Notes: Data are from 2019.
+The definitions of non-haven and haven countries follow Tørsløv et al.
+(2023) and the data is from the authors’ website: https://missingprofits.world (corresponding figures in the spreadsheets of ‘Table 1’ and ‘Table U1’ in the excel file ‘1975-2019 updated estimates: Tables‘).
+The non-haven countries are $30$ OECD countries, $7$ major developing countries and the rest of the world.
+There are $40$ haven countries across the world.
+The corporate tax rates for the representative non-haven and haven countries are calculated as the GDP-weighted averages of their respective effective tax rates.
+Tørsløv et al.
+(2023) report the pre-tax profits of foreign-owned MNEs only, which do not include those of domestically-owned MNEs (see their Section 3.1.2 on p.7).
+Using the CbCR Statistics of the EU Tax Observatory (https://www.taxobservatory.eu/repository/the-cbcr-explorer/), we compute the pre-tax profit ratio of domestically-owned MNEs to foreign-owned MNEs.
+With this ratio ($0.785$) and the pre-tax profits of foreign-owned MNEs ($2,590$ billion USD), we obtain the total pre-tax profits of foreign- and domestically-owned MNEs as $(1+0.785)\times 2,590=4,623$ billion USD.
+Table E.2
+GMT reforms when MNEs’ pre-tax profits depend on taxes.
+(a) Gains from a 15% GMT with $\phi=0.9$ Non-haven Haven World
+Tax rate after GMT introduction $20.7\%$ for all $15.0\%$ for all —–
+Welfare change in bUSD (% change) $108.9$ ($8.3\%$) 2.5-2.5 ($2.8\%$) $106.3$ ($7.6\%$)
+Revenue change in bUSD (% change) $140.7$ ($20.8\%$) 2.5-2.5 ($2.8\%$) $138.2$ ($18.0\%$)
+Profit change in bUSD (% change) $386.9$ $(39.9\%)$ 386.9-386.9 $(-39.9\%)$ 0 $(0\%)$
+(b) Gains from a 16% GMT with $\phi=0.9$ Non-haven Haven World
+Tax rate after GMT introduction $20.8\%$ for all $\begin{cases}$10.4\%$&\text{for small}\\ $16.0\%$&\text{for large}\end{cases}$ —–
+Welfare change in bUSD (% change) $118.1$ ($9.0\%$) 7.4-7.4 ($8.3\%$) $110.7$ ($7.9\%$)
+Revenue change in bUSD (% change) $152.0$ ($22.4\%$) 7.4-7.4 ($8.3\%$) $144.6$ ($18.8\%$)
+Profit change in bUSD (% change) $416.5$ $(43.0\%)$ 416.5-416.5 $(-43.0\%)$ 0 $(0\%)$
+[…]
+Tax rate after GMT introduction $21.0\%$ for all $16.0\%$ for all —–
+Welfare change in bUSD (% change) $129.3$ ($9.9\%$) 7.4-7.4 ($8.3\%$) $121.9$ ($8.7\%$)
+Revenue change in bUSD (% change) $166.3$ ($24.5\%$) 7.4-7.4 ($8.3\%$) $158.9$ ($20.7\%$)
+Shifting-profits change in bUSD (% change) $453.6$ $(46.8\%)$ 453.6-453.6 $(-46.8\%)$ 0 $(0\%)$
+Notes: Panel (a) shows the effects of raising the GMT rate from $t_{M}=0$ to $0.15$ with partial coverage ($\phi=0.9$); panel (b), from $t_{M}=0$ to $0.16$ with $\phi=0.9$; and panel (c), from $t_{M}=0$ to $0.17$ with full coverage ($\phi=1.0$).
+The tax rates in the initial equilibrium (Regime 0) are $0.186$ in the non-haven and $0.093$ in the haven.
+The regime-switching rate between Regimes 1 and 2 is 15.9%.
+Results are based on the calibrated parameter values described in Table E.1.
+The welfare levels are normalized by the MVPF, $\lambda$.
+The ‘haven’ here refers to the sum of the $H=40$ haven countries.